### v2

# Positivstellensätze for polynomial matrices with universal quantifiers

Abstract.
This paper studies Positivstellensätze for a polynomial matrix subject to polynomial matrix inequality constraints with universal quantifiers. We first present a Scherer-Hol-type Positivstellensatz under the Archimedean condition. When the objective is a scalar polynomial, we further provide a sparse Scherer-Hol-type Positivstellensatz in the presence of correlative sparsity. Next, without assuming the Archimedean condition, we derive Putinar-Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze under the same setting. These results can be viewed as common generalizations of corresponding Positivstellensätze in the cases of polynomials, polynomials with universal quantifiers, and polynomial matrices. For the proofs, techniques from $\star$-algebra, real algebraic geometry, operator theory, and convex optimization are employed. Applications of the established Positivstellensätze to robust polynomial matrix optimization are also discussed.
Key words and phrases:
Positivstellensatz, sum of squares, polynomial matrix inequality, polynomial matrix optimization, universal quantifier, correlative sparsity
1991 Mathematics Subject Classification:
90C23, 15A54, 13J30, 14P10, 11E25, 12D15
1. Introduction
Positivstellensätze are fundamental results in real algebraic geometry, asserting under which conditions a polynomial is guaranteed to be positive on a given set [6, 31, 35, 44]. Most Positivstellensätze achieve this by expressing the polynomial using sums of squares (SOS). These powerful results offer constructive methods to certify whether a polynomial $f$ is positive over a basic semialgebraic set defined by
$\mathbf{K}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid g_{j}({\boldsymbol{x}})\geq 0,j=1,\ldots,s\},$
and have widespread applications in areas like optimization, algebraic geometry, control theory, and more where polynomial positivity is a key concern [5, 14, 29, 31, 36].
Depending on whether the semialgebraic set $\mathbf{K}$ is assumed to be compact, Positivstellensätze are divided into two categories. When $\mathbf{K}$ is a compact polyhedron with non-empty interior, Handelman’s Positivstellensatz [13] states that a polynomial $f$ that is positive on $\mathbf{K}$ can be expressed as a positive linear combination of cross-products of $g_{j}$’s. In the case of a more general compact basic semialgebraic set $\mathbf{K}$, Schmüdgen [47] demonstrated that if a polynomial $f$ is positive on $\mathbf{K}$, then it belongs to the preordering generated by $g_{j}$’s, that is, $f$ can be represented as an SOS-weighted combination of cross-products of $g_{j}$’s. Putinar’s Positivstellensatz [41] provides an alternative representation that avoids the need for cross-products of $g_{j}$’s under the Archimedean condition (slightly stronger than compactness). Building on Putinar’s Positivstellensatz, Lasserre [26, 28, 29] introduced a moment-SOS hierarchy of semidefinite relaxations for polynomial optimization. This framework generates a non-decreasing sequence of lower bounds converging to the optimal value of a given polynomial optimization problem. When $f$ is only nonnegative on $\mathbf{K}$, Lasserre and Netzer [30] provided an SOS approximation of $f$ via high-degree perturbations.
There have been several attempts to provide certificates of positivity of a polynomial over a non-compact semialgebraic set. Without requiring compactness of $\mathbf{K}$, Krivine-Stengle Positivstellensatz [25, 50] states that a polynomial $f$ is positive over $\mathbf{K}$ if and only if $\psi f=1+\phi$ for some $\psi,\phi$ from the preordering generated by $g_{j}$’s. Pólya [38] demonstrated that if $f$ is homogeneous and positive on +${\mathbb{R}}^{n}_{+}\setminus\{\mathbf{0}\}$, then multiplying $f$ by some power of $\sum_{i=1}^{n}x_{i}$, one obtains a polynomial with nonnegative coefficients. Dickinson and Povh [9] generalized Pólya’s Positivstellensatz for homogeneous polynomials being positive on the intersection +${\mathbb{R}}^{n}_{+}\cap\mathbf{K}\setminus\{\mathbf{0}\}$, where $g_{j}$’s are assumed to be also homogeneous. Reznick [43] proved that after be multiplying by some power of $\sum_{i=1}^{n}x_{i}^{2}$, any positive definite (PD) form is a sum of even powers of linear forms. Putinar and Vasilescu [42] extended Reznick’s result to the constrained case where $f$ and $g_{j}$’s are homogeneous polynomial of even degree. When $\mathbf{K}$ is non-compact, SOS-structured certificates of positivity of $f$ over $\mathbf{K}$ could be also established by investigating various geometric objects associated with the data, e.g., gradient varieties [37], principal gradient tentacles [49], truncated tangency varieties [52], and the polar varieties [11].
Recently, Hu, Klep, and Nie [18] studied Positivstellensätze concerning semialgebraic sets defined by universal quantifiers (UQ). Specifically, for a given tuple $g=(g_{1},\ldots,g_{s})$ of real polynomials in ${\boldsymbol{x}}$ and ${\boldsymbol{y}}=(y_{1},\ldots,y_{m})$, and a closed set $\mathcal{Y}\subset{\mathbb{R}}^{m}$, they consider representations of polynomials that are positive over the semialgebraic set
$\mathcal{U}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid g_{1}({\boldsymbol{x}},{\boldsymbol{y}})\geq 0,\ldots,g_{s}({\boldsymbol{x}},{\boldsymbol{y}})\geq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\}.$
For a fixed measure $\nu$ with support exactly on $\mathcal{Y}$, under the Carleman condition on $\nu$ and the Archimedean condition, they proved that if a polynomial $f$ is positive on $\mathcal{U}$, then $f$ belongs to the quadratic module associated to $(g,\nu)$. In other words, $f$ admits a representation
$\sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\sigma_{1}({\boldsymbol{x}},{\boldsymbol{y}})g_{1}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})+\cdots+\int_{\mathcal{Y}}\sigma_{s}({\boldsymbol{x}},{\boldsymbol{y}})g_{s}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}}),$
where each $\sigma_{j}$, $j=0,1,\ldots,s$, are SOS polynomials. They also investigated the corresponding moment problem on the semialgebraic set $\mathcal{U}$. As an important application, their results could be used to solve semi-infinite optimization problems which are highly challenge.
Most Positivstellensätze could be generalized to the matrix setting that both $f$ and $g_{j}$’s are polynomial matrices. Scherer and Hol [45] developed a matrix-version of Putinar’s Positivstellensatz. Cimprič [7] extended Krivine-Stengle Positivstellensatz to the case of polynomial matrices with polynomial constraints. A matrix version of Handelman’s Positivstellensatz was proposed in [32]. Building on Scherer and Hol’s Positivstellensatz, Dinh et al. [10] generalized the classical Schmüdgen, Putinar-Vasilescu, and Dickinson–Povh Positivstellensätze to the polynomial matrix setting. Given the projections of two semialgebraic sets defined by polynomial matrix inequalities (PMI), Klep and Nie [22] provided a matrix Positivstellensatz with lifting polynomials to determine whether one is contained in the other. These generalizations have a wide range of applications, particularly in areas such as optimal control, systems theory [15, 16, 19, 40, 51].
Computing SOS-structured representations involved in Positivstellensätze could be typically cast as semidefinite programs (SDP). However, the size of the SDPs grows rapidly with the problem dimension. Hence from the perspective of computation, it becomes appealing to develop sparse versions of Positivstellensätze for sparse data, e.g., correlative sparsity [21, 27, 53], term sparsity [21, 33, 54, 55, 56], and matrix (chordal) sparsity [57, 58].
Due to estimation errors or lack of information, the data of real-world problems often involve uncertainty. As a result, ensuring the robustness of PMIs over a prescribed set with uncertainty is crucial for some safety-critical applications with little tolerance for failure [3]. Consequently, Positivstellensätze for polynomial matrices with UQs will be a powerful mathematical tool for addressing this issue, which serves as the primary motivation for this work. Precisely, consider the semialgebraic set
$\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$ (1)
where $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ (the set of $q\times q$ symmetric polynomial matrices in ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$), and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is closed. The goal of this paper is to provide certificates for positive definiteness of a $p\times p$ symmetric polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ over $\mathcal{X}$.
Table 1. Summary of Positivstellensätze for polynomials, polynomials with UQs, polynomial matrices, and polynomial matrices with UQs
Positivstellensatz Polynomials Polynomials with UQs PMIs PMIs with UQs
Putinar Putinar [41] Hu, Klep, and Nie [18] Scherer and Hol [45] Theorem 3.1
Putinar–Vasilescu Putinar and Vasilescu [42] Theorem 4.1 Dinh et al. [10] Theorem 4.1
Pólya Pólya [38] Theorem 4.2 Theorem 4.2 Theorem 4.2
Lasserre-Netzer Lasserre and Netzer [30] Theorem 4.4 Theorem 4.4 Theorem 4.4
Table 2. Summary of sparse Positivstellensätze for polynomials, polynomial matrices, and polynomial matrices with UQs
Case Literature
Polynomials with polynomial constraints Lasserre [27]
Polynomials with PMI constraints Kojima and Muramatsu [23]
Polynomial matrices with polynomial constraints Counterexample [21]
Polynomials with PMI constraints and UQs Theorem 3.2
Contributions. We generalize several classical Positivstellensätze for scalar polynomials to the matrix setting with UQs (see Tables 1 and 2 for summaries). In the following, we highlight the main results of this paper. Throughout the paper, let $\nu$ be a fixed Borel measure on ${\mathbb{R}}^{m}$ with support $\mathrm{supp}(\nu)=\mathcal{Y}$ and satisfying $\int_{\mathcal{Y}}|h({\boldsymbol{y}})|\mathrm{d}\nu({\boldsymbol{y}})<\infty$ for all $h({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]$. For any $H({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{p}$, let us write $H({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\beta}}}$ as a polynomial matrix in ${\boldsymbol{y}}$ with coefficient matrices $H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and let
$\int_{\mathcal{Y}}H({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\coloneqq\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\mathrm{d}\nu({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}.$
The main foundational result of this paper is the following Scherer-Hol-type Positivestellensatz which holds under the Archimedean condition (Assumption 3.1) and the Carleman condition on $\nu$ (Assumption 3.2).
Theorem A. (Theorem 3.1) Suppose that Assumptions 3.1 and 3.2 hold. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then there exists SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ such that
$F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
See Section 2.1 for the definition of the product $\langle\cdot,\cdot\rangle_{p}$. We adopt a notably nontrivial operator-theoretic approach to prove Theorem A, utilizing techniques such as the Gelfand-Naimark-Segal (GNS) construction for a carefully constructed $\star$-algebra, and the spectral theory of self-adjoint operators in a Hilbert space.
Then we consider the particular case of $p=1$, namely, the objective is a scalar polynomial. Suppose that $\mathcal{X}$ is defined by multiple PMIs with UQs ($G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,j=1,\ldots,s$). We assume the presence of correlative sparsity in the problem data, which implies that the variables ${\boldsymbol{x}}$ decompose as a union of subsets ${\boldsymbol{x}}=\cup_{\ell=1}^{t}{\boldsymbol{x}}(\mathcal{I}_{\ell})$ such that $\{G_{j}\}_{j=1}^{s}=\cup_{\ell=1}^{t}\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, $\ell=1,\ldots,t$ (see Assumption 3.3). Under these conditions, we could give the following sparse version of the Scherer-Hol-type Positivstellensatz.
Theorem B. (Theorem 3.2) Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\mathcal{X}$, and Assumption 3.1 holds with respect to each ${\boldsymbol{x}}(\mathcal{I}_{\ell})$. If $f>0$ on $\mathcal{X}$, then there exist SOS polynomials $\sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}],\sigma_{\ell,j}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]$ such that
$f({\boldsymbol{x}})=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}})+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\langle\sigma_{\ell,j}({\boldsymbol{x}},{\boldsymbol{y}}),G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\rangle\mathrm{d}\nu({\boldsymbol{y}})\right).$
Next, building on Theorem A and existing techniques, we establish several Positivstellensätze for polynomial matrices with UQs and without assuming the Archimedean condition. The first such Positivstellensatz is of Putinar–Vasilescu-type.
Theorem C. (Theorem 4.1) Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$ of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and $\deg\Sigma_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}\Sigma=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The second such Positivstellensatz is of Pólya-type.
Theorem D. (Theorem 4.2) Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{{\boldsymbol{\alpha}}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy each $S_{0,{\boldsymbol{\alpha}}}\succ 0$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$\left(\sum_{i=1}^{n}x_{i}\right)^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The third such Positivstellensatz is of Lasserre-Netzer-type.
Theorem E. (Theorem 4.4) Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is positive semidefinite (PSD) on $\mathcal{X}\cap[-1,1]^{n}$. Then for any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1},k\geq\ell_{2}$, it holds
$F({\boldsymbol{x}})+\varepsilon\left(1+\sum_{i=1}^{n}x_{i}^{2d}\right)I_{p}=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg(\Sigma_{0})\leq 2d$, $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$, and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$.
To prove Theorem E, we studied the exponentially bounded matrix-valued moment problem, providing a characterization for an exponentially bounded matrix-valued sequence to admit a matrix-valued representing measure supported on the set $\mathcal{X}\cap[-C,C]^{n}$ for a given $C>0$ (Theorem 4.3). This result extends the work of Berg and Maserick [2] to the matrix case and offers novel insights on its own.
The rest of this paper is organized as follows. In Section 2, we review some preliminary concepts. Section 3 presents a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs under the Archimedean condition, along with a sparse version in the presence of correlative sparsity. In Section 4, without relying on the Archimedean condition, we derive Putinar–Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze for polynomial matrices with UQs. Section 5 briefly discusses applications of the established Positivstellensätze to robust polynomial matrix optimization. For the sake of readability, some lengthy and technical proofs are deferred to Section 6. Conclusions are given in Section 7.
2. Preliminaries
We collect some notation and basic concepts which will be used in this paper. We denote by ${\boldsymbol{x}}$ (resp., ${\boldsymbol{y}}$) the $n$-tuple (resp., $m$-tuple) of variables $(x_{1},\ldots,x_{n})$ (resp., $(y_{1},\ldots,y_{m})$). The symbol $\mathbb{N}$ (resp., ${\mathbb{R}}$, ${\mathbb{R}}_{+}$) denotes the set of nonnegative integers (resp., real numbers, nonnegative real numbers). For positive integer $n\in\mathbb{N}$, denote by $[n]$ the set $\{1,\ldots,n\}$. Denote by ${\mathbb{R}}^{p}$ (resp. ${\mathbb{R}}^{l_{1}\times l_{2}}$, ${\mathbb{S}}^{p}$, ${\mathbb{S}}_{+}^{p}$) the $p$-dimensional real vector (resp. $l_{1}\times l_{2}$ real matrix, $p\times p$ symmetric real matrix, $p\times p$ PSD matrix) space. Denote by +${\mathbb{R}}^{n}_{+}$ the nonnegative orthant of ${\mathbb{R}}^{n}$. For ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ (resp., $N\in{\mathbb{R}}^{l_{1}\times l_{2}}$), the symbol ${\boldsymbol{v}}^{\intercal}$ (resp., $N^{\intercal}$) denotes the transpose of ${\boldsymbol{v}}$ (resp., $N$). For a matrix $N\in{\mathbb{R}}^{p\times p}$, $\mbox{tr}\left(N\right)$ denotes its trace. For two matrices $N_{1}$ and $N_{2}$, $N_{1}\otimes N_{2}$ denotes the Kronecker product of $N_{1}$ and $N_{2}$. For two matrices $N_{1}$ and $N_{2}$ of the same size, $\langle N_{1},N_{2}\rangle$ denotes the inner product $\mbox{tr}\left(N_{1}^{\intercal}N_{2}\right)$ of $N_{1}$ and $N_{2}$. The notation $I_{p}$ denotes the $p\times p$ identity matrix. For any $t\in{\mathbb{R}}$, $\lceil t\rceil$ (resp., $\lfloor t\rfloor$) denotes the smallest (resp., largest) integer that is not smaller (resp., larger) than $t$. For ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$, $\|{\boldsymbol{u}}\|$ denotes the standard Euclidean norm of ${\boldsymbol{u}}$. For $N\in{\mathbb{R}}^{l_{1}\times l_{2}}$, $\|N\|$ denotes the spectral norm of $N$. For a vector ${\boldsymbol{\alpha}}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}$, let $|{\boldsymbol{\alpha}}|=\alpha_{1}+\cdots+\alpha_{n}$. For a set $A$, we use $|A|$ to denote its cardinality. For $k\in\mathbb{N}$, let $\mathbb{N}^{n}_{k}\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid|{\boldsymbol{\alpha}}|\leq k\}$ and $|\mathbb{N}^{n}_{k}|=\binom{n+k}{k}$ be its cardinality. For variables ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$, ${\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ denotes the monomial $x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. Let ${\mathbb{R}}[{\boldsymbol{x}}]$ (resp. ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$) denote the set of real polynomials (resp. $p\times p$ symmetric real polynomial matrices) in ${\boldsymbol{x}}$. For $h\in{\mathbb{R}}[{\boldsymbol{x}}]$ (resp., $h\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]$), we denote by $\deg(h)$ (resp., $\deg_{{\boldsymbol{x}}}(h)$) its total degree in ${\boldsymbol{x}}$. For a polynomial matrix $T({\boldsymbol{x}})=[T_{ij}({\boldsymbol{x}})]$ (resp., $T({\boldsymbol{x}},{\boldsymbol{y}})=[T_{ij}({\boldsymbol{x}},{\boldsymbol{y}})]$), denote $\deg(T)\coloneqq\max_{i,j}\deg(T_{ij})$ (resp., $\deg_{{\boldsymbol{x}}}(T)\coloneqq\max_{i,j}\deg_{{\boldsymbol{x}}}(T_{ij})$). For $k\in\mathbb{N}$, denote by ${\mathbb{R}}[{\boldsymbol{x}}]_{k}$ (resp., ${\mathbb{S}}[{\boldsymbol{x}}]^{p}_{k}$) the subset of ${\mathbb{R}}[{\boldsymbol{x}}]$ (resp., ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$) of degree up to $k$. For any $P({\boldsymbol{x}})=[P_{ij}({\boldsymbol{x}})]\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $Q({\boldsymbol{x}},{\boldsymbol{y}})=[Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})]\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$, denote
$\mathrm{supp}(P)$$\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\ \text{appears in some }P_{ij}({\boldsymbol{x}})\},$
$\mbox{supp}_{{\boldsymbol{x}}}(Q)$$\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\ \text{appears in some }Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})\text{ for some }{\boldsymbol{\beta}}\in\mathbb{N}^{m}\},$
$\mbox{supp}_{{\boldsymbol{y}}}(Q)$$\coloneqq\{{\boldsymbol{\beta}}\in\mathbb{N}^{m}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\ \text{appears in some }Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})\text{ for some }{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\}.$
2.1. SOS matrices and positivstellensätz for polynomial matrices
For a polynomial $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$, if there exist polynomials $f_{1}({\boldsymbol{x}}),\ldots,f_{t}({\boldsymbol{x}})$ such that $f({\boldsymbol{x}})=\sum_{i=1}^{t}f_{i}({\boldsymbol{x}})^{2}$, then we call $f({\boldsymbol{x}})$ an SOS. A polynomial matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is said to be an SOS matrix if there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. For $d\in\mathbb{N}$, denote by $[{\boldsymbol{x}}]_{d}$ the canonical basis of ${\mathbb{R}}[{\boldsymbol{x}}]_{d}$, i.e.,
$[{\boldsymbol{x}}]_{d}\coloneqq[1,\ x_{1},\ x_{2},\ \cdots,\ x_{n},\ x_{1}^{2},\ x_{1}x_{2},\ \cdots,\ x_{n}^{d}]^{\intercal},$ (2)
whose cardinality is $|\mathbb{N}^{n}_{d}|=\binom{n+d}{d}$. With $d=\deg(T)$, we can write $T({\boldsymbol{x}})$ as
$T({\boldsymbol{x}})=Q([{\boldsymbol{x}}]_{d}\otimes I_{p})\text{ with }Q=[Q_{1},\ldots,Q_{|\mathbb{N}^{n}_{d}|}],\quad Q_{i}\in{\mathbb{R}}^{l\times p},$
where $Q$ is the vector of coefficient matrices of $T({\boldsymbol{x}})$ with respect to $[{\boldsymbol{x}}]_{d}$. Hence, $\Sigma({\boldsymbol{x}})$ is an SOS matrix with respect to $[{\boldsymbol{x}}]_{d}$ if there exists some $Q\in{\mathbb{R}}^{l\times p|\mathbb{N}^{n}_{d}|}$ satisfying
$\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}(Q^{\intercal}Q)([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
We thus have the following results.
Proposition 2.1.
[45, Lemma 1] A polynomial matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is an SOS matrix with respect to the monomial basis $[{\boldsymbol{x}}]_{d}$ if and only if there exists $Z\in\mathbb{S}_{+}^{p|\mathbb{N}^{n}_{d}|}$ such that
$\Sigma({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}Z([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
Lemma 2.1.
Let $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be an SOS matrix and $\Sigma_{k}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{k\times k}$ be a principal submatrix of $\Sigma({\boldsymbol{x}})$ whose rows and columns are indexed by $(p_{1},\ldots,p_{k})$ with $1\leq p_{1}<\cdots<p_{k}\leq p$, then $\Sigma_{k}({\boldsymbol{x}})$ is an SOS matrix.
Proof.
As $\Sigma({\boldsymbol{x}})$ is an SOS matrix, there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$. Denote by $T_{k}({\boldsymbol{x}})$ the submatrix of $T({\boldsymbol{x}})$ consisting of the colums of $T(x)$indexed by $(p_{1},\ldots,p_{k})$. Then, $\Sigma_{k}({\boldsymbol{x}})=T_{k}({\boldsymbol{x}})^{\intercal}T_{k}({\boldsymbol{x}})$ and hence is an SOS matrix. ∎
We next recall Scherer-Hol’s Positivstellensatz for polynomial matrices obtained in [45]. Define the bilinear mapping
$\left\langle\cdot,\cdot\right\rangle_{p}\colon{\mathbb{R}}^{pq\times pq}\times{\mathbb{R}}^{q\times q}\to{\mathbb{R}}^{p\times p},\quad\left\langle A,B\right\rangle_{p}=\mbox{tr}_{p}\left(A^{\intercal}(I_{p}\otimes B)\right),$
with
$\mbox{tr}_{p}\left(C\right)\coloneqq\left[\begin{array}[]{ccc}\mbox{tr}\left(C_{11}\right)&\cdots&\mbox{tr}\left(C_{1p}\right)\\ \vdots&\ddots&\vdots\\ \mbox{tr}\left(C_{p1}\right)&\cdots&\mbox{tr}\left(C_{pp}\right)\end{array}\right]\quad\text{for }C=[C_{ij}]_{i,j\in[p]}\in{\mathbb{R}}^{pq\times pq},C_{ij}\in{\mathbb{R}}^{q\times q}.$
When $p=1$, the product $\left\langle A,B\right\rangle_{1}$ coicides with the matrix inner product $\langle A,B\rangle=\mbox{tr}\left(A^{\intercal}B\right)$.
Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$.
Lemma 2.2.
Let $H({\boldsymbol{x}})=\mbox{diag}(H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))$ be a block diagonal matrix, then for any SOS matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr}$ where $p\in\mathbb{N}$ and $r=r_{1}+\cdots+r_{t}$, there are SOS matrices $\Sigma_{j}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$, $j\in[t]$, such that
$\left\langle\Sigma({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
Proof.
For each $j\in[t]$, let $\Sigma_{j}({\boldsymbol{x}})$ be the $pr_{j}\times pr_{j}$ principal submatrix of $\Sigma({\boldsymbol{x}})$ whose rows and columns are indexed by
$\left(\sum_{\ell=1}^{j-1}r_{\ell}+1,\sum_{\ell=1}^{j}r_{\ell},r+\sum_{\ell=1}^{j-1}r_{\ell}+1,r+\sum_{\ell=1}^{j}r_{\ell},\ldots,(p-1)r+\sum_{\ell=1}^{j-1}r_{\ell}+1,(p-1)r+\sum_{\ell=1}^{j}r_{\ell}\right).$
As $H({\boldsymbol{x}})$ is block diagonal, by the definition of the mapping .,.$\left\langle.,.\right\rangle_{p}$, it is easy to see that
$\left\langle\Sigma({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
By Lemma 2.1, each $\Sigma_{j}({\boldsymbol{x}})$ is an SOS matrix. ∎
The matrix quadratic module $\mathcal{Q}^{p}(\mathbf{H})$ generated by $\mathbf{H}$ is defined as
$\mathcal{Q}^{p}(\mathbf{H})\coloneqq\left\{\Sigma_{0}({\boldsymbol{x}})+\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr_{j}},j\in[t],\ \text{are SOS}\right\}.$
Assumption 2.1.
$\mathcal{Q}^{p}(\mathbf{H})$ is Archimedean, i.e., there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(\mathbf{H})$.
Theorem 2.1.
(Scherer-Hol’s Positivestellensatz) Let Assumption 2.1 and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be PD on $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$. Then, $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{H})$.
Proof.
It was proved in [45, Corollary 1] for the case $t=1$. The case $t>1$ can be derived from [45, Corollary 1] and Lemma 2.2. ∎
2.2. Matrix-valued measures and moment problem
Now we recall some background on the concept of matrix-valued measures. Denote by $\mathfrak{B}(\mathcal{X})$ the smallest $\sigma$-algebra generated from the open subsets of $\mathcal{X}$ and by $\mathfrak{m}(\mathcal{X})$ the set of all finite Borel measures on $\mathcal{X}$. The support $\mathrm{supp}(\phi)$ of a Borel measure $\phi\in\mathfrak{m}(\mathcal{X})$ is the (unique) smallest closed set $\mathbf{A}\in\mathfrak{B}(\mathcal{X})$ such that $\phi(\mathcal{X}\setminus\mathbf{A})=0$. Let $\phi_{ij}\in\mathfrak{m}(\mathcal{X})$, $i,j=1,\ldots,p$. The $p\times p$ matrix-valued measure $\Phi$ on $\mathcal{X}$ is defined as the matrix-valued function $\Phi\colon\mathfrak{B}(\mathcal{X})\to{\mathbb{R}}^{p\times p}$ with
$\Phi(\mathbf{A})\coloneqq[\phi_{ij}(\mathbf{A})]\in{\mathbb{R}}^{p\times p},\quad\forall\mathbf{A}\in\mathfrak{B}(\mathcal{X}).$
If $\phi_{ij}=\phi_{ji}$ for all $i,j=1,\ldots,p$, we call $\Phi$ a symmetric matrix-valued measure. If ${\boldsymbol{v}}^{\intercal}\Phi(\mathbf{A}){\boldsymbol{v}}\geq 0$ holds for all $\mathbf{A}\in\mathfrak{B}(\mathcal{X})$ and for all column vectors ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$, we call $\Phi$ a PSD matrix-valued measure. The set $\mathrm{supp}(\Phi)\coloneqq\bigcup_{i,j=1}^{p}\mathrm{supp}(\phi_{ij})$ is called the support of the matrix-valued measure $\Phi$. We denote by $\mathfrak{M}^{p}_{+}(\mathcal{X})$ the set of all $p\times p$ PSD symmetric matrix-valued measures on $\mathcal{X}$.
For a polynomial matrix $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and a matrix-valued measure $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$, the integral of $H$ with respect to $\Phi$ is defined by
$\int_{\mathcal{X}}H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\coloneqq\int_{\mathcal{X}}\ \mbox{tr}\left(H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\right)=\sum_{i,j}\int_{\mathcal{X}}\ H_{ij}({\boldsymbol{x}})\mathrm{d}\phi_{ij}({\boldsymbol{x}}).$
Let $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ be a multi-indexed sequence of symmetric matrices in ${\mathbb{S}}^{p}$. We define a linear functional $\mathscr{L}_{\mathbf{S}}\colon{\mathbb{S}}[{\boldsymbol{x}}]^{p}\to{\mathbb{R}}$ in the following way:
$\mathscr{L}_{\mathbf{S}}(H)\coloneqq\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}\mbox{tr}\left(H_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\alpha}}}\right),\quad\forall H({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}H_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}.$
We call $\mathscr{L}_{\mathbf{S}}$ the Riesz functional associated to the sequence $\mathbf{S}$. We say the sequence $\mathbf{S}$ has a matrix-valued representing measure $\Phi=[\phi_{ij}]\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ if
$S_{{\boldsymbol{\alpha}}}=\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})\coloneqq\left[\int_{\mathbf{K}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi_{ij}({\boldsymbol{x}})\right]_{i,j\in[p]},\ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}.$ (3)
The following theorem is a matrix version of Haviland’s theorem ([8, 46]).
Theorem 2.2 (Haviland’s theorem for polynomial matrices).
[8, Theorem 3] A given sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ has a matrix-valued representing measure $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ if and only if $\mathscr{L}_{\mathbf{S}}(H)\geq 0$ for all $H({\boldsymbol{x}})$ which are PSD on $\mathcal{X}$.
[…]
By a $\star$-algebra we mean an algebra $\mathcal{A}$ over the field ${\mathbb{K}}={\mathbb{R}}$ or ${\mathbb{K}}={\mathbb{C}}$ equipped with an involution $a\to a^{\star}$ of $\mathcal{A}$ into itself such that $(\lambda a+\mu b)^{\star}=\bar{\lambda}a^{\star}+\bar{\mu}b^{\star}$, $(ab)^{\star}=b^{\star}a^{\star}$ and $(a^{\star})^{\star}$ for all $a,b\in\mathcal{A}$ and $\lambda,\mu\in{\mathbb{K}}$. Denote by $1$ the unit element of $\mathcal{A}$. If $a$ is an operator on a Hilbert space, denote by $\mathcal{D}(a)$ its domain.
[…]
A state of $\mathcal{A}$ is a $\star$-linear functional $\Pi$ on $\mathcal{A}$ such that $\Pi(1)=1$ and $\Pi(a^{\star}a)\geq 0$ for all $a\in\mathcal{A}$. If $\pi$ is a $\star$-representation of $\mathcal{A}$ and $\varphi\in\mathcal{D}(\pi)$ is a unit vector, then $\Pi(\cdot)\coloneqq\langle\pi(\cdot)\varphi,\varphi\rangle$ is a state of $\mathcal{A}$. Conversely, for each state $\Pi$ on $\mathcal{A}$ there exist a $\star$-representation of $\mathcal{A}$, called the GNS representation $\pi_{\Pi}$ of $\Pi$, and a vector $\varphi_{\Pi}\in\mathcal{D}(\pi_{\Pi})$ such that $\mathcal{D}(\pi_{\Pi})=\pi_{\Pi}(\mathcal{A})\varphi_{\Pi}$ and $\Pi(a)=\langle\pi_{\Pi}(a)\varphi_{\Pi},\varphi_{\Pi}\rangle$ for all $a\in\mathcal{A}$.
3. Positivstellensätze for polynomial matrices with UQs under the Archimedean condition
In this section, we shall present a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs, assuming the Archimedean condition, along with a sparse version in the presence of correlative sparsity.
3.1. Scherer-Hol-type Positivestellensatz
Recall the set $\mathcal{X}$ in (1). Throughout the paper, let $\nu$ be a fixed Borel measure on ${\mathbb{R}}^{m}$ with support $\mathrm{supp}(\nu)=\mathcal{Y}$ and satisfying $\int_{\mathcal{Y}}|h({\boldsymbol{y}})|\mathrm{d}\nu({\boldsymbol{y}})<\infty$ for all $h({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]$. Similarly to the scalar case in [18], let us define the matrix quadratic module associated with $(G,\nu)$ as follows.
Definition 3.1.
The matrix quadratic module $\mathcal{Q}^{p}(G,\nu)$ generated by $G$ and $\nu$ is defined as
$\mathcal{Q}^{p}(G,\nu)\coloneqq\left\{\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices}\right\}.$
Let us first derive the Scherer-Hol-type Positivestellensatz to provide SOS-structured characterizations for polynomial matrices that are PD over $\mathcal{X}$.
Assumption 3.1.
$\mathcal{Q}^{p}(G,\nu)$ is Archimedean, i.e., there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)$.
Lemma 3.1.
Under Assumption 3.1, the identity matrix $I_{p}$ belongs to the algebraic interior of the cone $\mathcal{Q}^{p}(G,\nu)$ in ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$, i.e., for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, there exists a real number $t_{0}\geq 0$ such that $H({\boldsymbol{x}})+tI_{p}\in\mathcal{Q}^{p}(G,\nu)$ for all $t\geq t_{0}$.
Proof.
Without loss of generality, we assume that $H({\boldsymbol{x}})=\pm{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ for some ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ and ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. We first show by induction that for any ${\boldsymbol{\beta}}\in\mathbb{N}^{n}$, $C^{|{\boldsymbol{\beta}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}\in\mathcal{Q}^{1}(G,\nu)$. Note that $C-x_{j}^{2}\in\mathcal{Q}^{1}(G,\nu)$ since
$C-x_{j}^{2}=(C-\|{\boldsymbol{x}}\|^{2})+\sum_{i\neq j}x_{j}^{2}\in\mathcal{Q}^{1}(G,\nu).$
Assume that the claim holds for ${\boldsymbol{\beta}}$ and consider the monomial ${\boldsymbol{x}}^{{\boldsymbol{\beta}}}x_{j_{0}}$ for some $j_{0}\in[n]$. We have
$C^{|{\boldsymbol{\beta}}|+1}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}x_{j_{0}}^{2}=C(C^{|{\boldsymbol{\beta}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}})+{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}(C-x_{j_{0}}^{2})\in\mathcal{Q}^{1}(G,\nu),$
by induction.
Now note that
±$\pm 2{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}={\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(({\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\pm 1)^{2}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}}-1)$
and
${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}}-1)={\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(C^{|{\boldsymbol{\alpha}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}})+(C^{|{\boldsymbol{\alpha}}|}+1)({\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal})-(C^{|{\boldsymbol{\alpha}}|}+1){\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}.$
By Cauchy’s inequality, ${\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}\succeq 0$. Hence, by the above claim,
${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(C^{|{\boldsymbol{\alpha}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}})+(C^{|{\boldsymbol{\alpha}}|}+1)({\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal})\in\mathcal{Q}^{p}(G,\nu).$
Thus, by letting $t_{0}=\frac{(C^{|{\boldsymbol{\alpha}}|}+1){\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}}{2}$, we have $H({\boldsymbol{x}})+tI_{p}\in\mathcal{Q}^{p}(G,\nu)$ for all $t\geq t_{0}$. ∎
Corollary 3.1.
If $F({\boldsymbol{x}})\not\in\mathcal{Q}^{p}(G,\nu)$, then there exists a linear functional $\mathscr{L}:{\mathbb{S}}[{\boldsymbol{x}}]^{p}\to{\mathbb{R}}$ such that $\mathscr{L}(\mathcal{Q}^{p}(G,\nu))\subseteq{\mathbb{R}}_{+}$, $\mathscr{L}(F)\leq 0$, and $\mathscr{L}(I_{p})=1$.
Proof.
As the cone $\mathcal{Q}^{p}(G,\nu)$ has non-empty algebraic interior by Lemma 3.1, this is due to the Hahn-Banach separation theorem [1, Corollary III.1.7]. ∎
Consider the Carleman condition imposed on $\nu$ which is automatically satisfied when $\mathcal{Y}$ is compact.
Assumption 3.2.
The Borel measure $\nu$ satisfies the multivariable Carleman condition
$\sum_{d=1}^{\infty}\left(\int_{\mathcal{Y}}y_{j}^{2d}\mathrm{d}\nu({\boldsymbol{y}})\right)^{-\frac{1}{2d}}=\infty,\quad\forall j\in[m].$
Proposition 3.1.
[18, Proposition 3.2] Suppose that $\nu$ satisfies Assumption 3.2. Then, SOS polynomials are dense in the cone of nonnegative functions in $L^{2}({\mathbb{R}}^{m},\nu)$.
Now, we present our main fundamental result concerning a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs.
Theorem 3.1.
Suppose that Assumptions 3.1 and 3.2 hold. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)$.
Proof.
The proof is by contradiction, using the functional $\mathscr{L}$ from Corollary 3.1, and proceeds as follows:
  1. 1.
    Construct a Hilbert space $\mathcal{H}$ and operators $\widehat{\mathbf{A}}\coloneqq(\widehat{A}_{i})_{i\in[n]}$ with $\widehat{A}_{i}\in\mathcal{B}(\mathcal{H})$, where $\mathcal{B}(\mathcal{H})$ denotes the set of bounded linear operators on $\mathcal{H}$. Then construct a state $\Pi$ on the the $p\times p$ matrix algebra $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ over $\mathcal{B}(\mathcal{H})$ such that $\Pi(F(\widehat{\mathbf{A}}))=\mathscr{L}(F)$;
  2. 2.
    By applying the GNS construction to $(\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})),\Pi)$, construct a Hilbert space $\mathcal{K}_{1}$ with inner product $\langle\cdot,\cdot\rangle_{\mathcal{K}_{1}}$ and operators $\mathbf{A}=(A_{i})_{i\in[n]}$ with $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ such that $\mathscr{L}(F)=\sum_{i,j=1}^{p}\left\langle F_{ij}(\mathbf{A})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}_{1}}$ for some $\eta_{i}\in\mathcal{K}_{1}$, $i\in[p]$;
  3. 3.
    By applying the spectral theorem to $\mathbf{A}=(A_{i})_{i\in[n]}$, construct a $p\times p$ PSD matrix-valued measure $\Phi$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ such that $\mathscr{L}(F)=\int_{{\mathbb{R}}^{n}}F({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})$;
  4. 4.
    Show that the support of $\Phi$ is contained in $\mathcal{X}$ and derive a contradiction by Corollary 3.1.
For the sake of readability, the details of the proof is postponed to Section 6.1. ∎
To conclude this part, let us derive a simple corollary of Theorem 3.1, which will be used in Section 4. We need the following lemma.
Lemma 3.2.
Suppose that $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ is an SOS matrix in ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$, then $\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in ${\boldsymbol{x}}$.
Proof.
As $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})$ is an SOS matrix, there exists an ×$\ell\times q$ polynomial matrix $T({\boldsymbol{x}},{\boldsymbol{y}})$ for some $\ell\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}T({\boldsymbol{x}},{\boldsymbol{y}})$. With $d=\deg_{{\boldsymbol{x}}}(T)$, we could write $T({\boldsymbol{x}},{\boldsymbol{y}})$ as
$T({\boldsymbol{x}},{\boldsymbol{y}})=Q({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})\text{ with }Q({\boldsymbol{y}})=[Q_{1}({\boldsymbol{y}}),\ldots,Q_{|\mathbb{N}^{n}_{d}|}({\boldsymbol{y}})],\quad Q_{i}({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]^{\ell\times q},$
where $Q({\boldsymbol{y}})$ is the vector of coefficient matrices of $T({\boldsymbol{x}},{\boldsymbol{y}})$ (considered as a polynomial matrix in ${\mathbb{R}}[{\boldsymbol{x}}]^{\ell\times q}$) with respect to $[{\boldsymbol{x}}]_{d}$. Hence,
$\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ $=\int_{\mathcal{Y}}([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}(Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}}))([{\boldsymbol{x}}]_{d}\otimes I_{q})\mathrm{d}\nu({\boldsymbol{y}})$
$=([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})$
As $\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is PSD, $\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in ${\boldsymbol{x}}$. ∎
Corollary 3.2.
Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$. Suppose that Assumption 3.2 holds and there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)+\mathcal{Q}^{1}(\mathbf{H})$. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$.
Proof.
It is clear that $F({\boldsymbol{x}})\succ 0$ on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succ 0,\ j\in[t]\}$ if and only if $F({\boldsymbol{x}})\succ 0$ on
$\left\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\,\middle|\,\widehat{G}({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\mbox{diag}(G({\boldsymbol{x}},{\boldsymbol{y}}),H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
By Lemmas 2.2 and 3.2, it is easy to see that $\mathcal{Q}^{p}(\widehat{G},\nu)=\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$. So the conclusion follows from Theorem 3.1. ∎
3.2. A sparse Positivstellensatz for polynomial matrices with UQs
There are sparse Positivstellensätze for scalar polynomials being positive on a basic semialgebraic set in the presence of correlative sparsity; see [12, 27] where the set is defined by polynomial inequalities and [23] where the set is defined PMIs. Let $p=1$. Building on the ideas from [12] and [23], we now prove a sparse Positivstellensatz for polynomial matrices with UQs. Let $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and
$\widehat{\mathcal{X}}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall j\in[s],\ {\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$
where each $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$, $q_{j}\in\mathbb{N}$.
Assumption 3.3 (correlative sparsity pattern).
Subsets {}$\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ of $[n]$ and subsets $\{\mathcal{J}_{\ell}\}_{\ell\in[t]}$ of $[s]$ satisfy the following conditions:
  1. (i)
    The running intersection property holds for {}$\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$, i.e.,
    for =$\text{for }\ell=2,\ldots,t,\ \exists k<\ell\quad\text{s.t.}\quad\mathcal{I}_{\ell}\cap\bigcup_{j<\ell}\mathcal{I}_{j}\subseteq\mathcal{I}_{k};$
  2. (ii)
    For every $\ell\in[t]$ and $j\in\mathcal{J}_{\ell}$, $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}}$, where ${\boldsymbol{x}}(\mathcal{I}_{\ell})\coloneqq\{x_{i}\}_{i\in\mathcal{I}_{\ell}}$;
  3. (iii)
    $f$ decomposes as $f=f_{1}+\cdots+f_{t}$ with each $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$.
For each $\ell\in[t]$, let $\mathbf{G}^{\ell}\coloneqq\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)$ be the quadratic module generated by $\mathbf{G}^{\ell}$ and $\nu$ in ${\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p}$, i.e.,
$\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)\coloneqq\left\{\Sigma_{0}+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle\Sigma_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p},\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{pq_{j}},j\in\mathcal{J}_{\ell},\text{ are SOS matrices}\right\}.$
Using the correlative sparsity pattern and the Archimedean condition, we are able to derive the following sparse Positivstellensatz.
Theorem 3.2.
Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f$ and $\widehat{\mathcal{X}}$, and Assumption 3.1 holds for each $\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$, $\ell\in[t]$. If $f>0$ on $\widehat{\mathcal{X}}$, then $f\in\sum_{\ell=1}^{t}\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$.
To prove Theorem 3.2, we need the following intermediary results.
Proposition 3.2.
[12, Lemma 3] Let {}$\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ satisfy the running intersection property. For any $C>0$, if $f=f_{1}+\cdots+f_{t}$ with $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ satisfies $f>0$ on $[-C,C]^{n}$, then $f=h_{1}+\ldots+h_{t}$ for some $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ with $h_{\ell}>0$ on $[-C,C]^{|\mathcal{I}_{\ell}|}$.
Proposition 3.3.
Suppose $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\widehat{\mathcal{X}}$. Then for any $C>0$, there exists $M>0$ and $\bar{k}\in\mathbb{N}$ such that
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succ 0$ (4)
on $[-C,C]^{n}$ for all $k\geq\bar{k}$.
The proof of Proposition 3.3 is postponed to Section 6.2. Note that using Proposition 3.3, we can provide a simpler proof of the Scherer-Hol-type Positivestellensatz (Theorem 3.1) assuming that $\mathcal{Y}$ is compact (see Section 6.2).
Proof of Theorem 3.2.
By assumption, let $C>0$ be such that $C-\|{\boldsymbol{x}}(\mathcal{I}_{\ell})\|^{2}\in\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$ holds for all $\ell\in[t]$. By Proposition 3.3, there exists $M>0$ and $k^{\prime}\in\mathbb{N}$ such that
$f({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})>0$
on $[-\sqrt{C},\sqrt{C}]^{n}$. Note that for each $j\in[s]$,
$\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$
for some $\ell\in[t]$. By Proposition 3.2, there exist $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$, $\ell\in[t]$, such that
$f({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})=h_{1}+\cdots+h_{t},$
and each $h_{\ell}>0$ on $[-\sqrt{C},\sqrt{C}]^{|\mathcal{I}_{\ell}|}$. By Putinar’s Positivstellensatz [41], for each $h_{\ell}$, there exist SOS polynomials $\sigma_{\ell,0},\sigma_{\ell,1}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ such that
$h_{\ell}=\sigma_{\ell,0}+\sigma_{\ell,1}(C-\|{\boldsymbol{x}}(\mathcal{I}_{\ell})\|^{2}).$
As each $C-\|{\boldsymbol{x}}(\mathcal{I}_{\ell})\|^{2}\in\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$, it holds $h_{\ell}\in\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu)$. Therefore,
$f({\boldsymbol{x}})=\sum_{\ell\in[t]}\left(\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})+h_{\ell}\right)\in\sum_{\ell=1}^{t}\mathcal{Q}^{1}(\mathbf{G}^{\ell},\nu).$
Remark 3.1.
One might wonder whether the result in Theorem 3.2 holds for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ with $p>1$. Indeed, this is not true even in the absence of UQs; see [21] for a counterexample.
4. Positivstellensätze for polynomial matrices with UQs and without the Archimedean condition
The results in Section 3 are derived under the Archimedean condition on the quadratic module generated by $(G,\nu)$, which requires $\mathcal{X}$ to be bounded. Lots of efforts have been made to provide SOS-structured representations for polynomials that are PD on a unbounded semialgebraic set. The goal of this section is to extend some of these well-known results to the setting of polynomial matrices with UQs.
4.1. Putinar–Vasilescu-type Positivstellensatz
Let $\theta\coloneqq 1+\|{\boldsymbol{x}}\|^{2}$. For every polynomial $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ nonnegative on a general basic semialgebraic set, Putinar and Vasilescu [42] proved that for a given $\varepsilon>0$ and $2d\geq\deg(f)$, there exists a nonnegative integer $k$ such that $\theta^{k}(f+\varepsilon\theta^{d})$ belongs to the quadratic module associated with the basic semialgebraic set. Using Jacobi’s technique [20], Mai et al. [34] provided an alternative proof of Putinar and Vasilescu’s result with an effective degree bound on polynomials involved in such certificates.
Building on Theorem 3.1 and similiar techniques from [34], we next derive a Putinar–Vasilescu-type Positivstellensatz for polynomial matrices with UQs.
In the homogeneous case, the result is stated as follows.
Theorem 4.1.
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$ of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and $\deg\Sigma_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}\Sigma=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
See Section 6.3 for its proof.
For any $H=[H_{ij}]\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{r}$, let $d_{H}\coloneqq\max\left\{\lfloor\deg_{{\boldsymbol{x}}}(H_{ij})/2\rfloor+1:i,j\in[r]\right\}$ and $\tilde{{\boldsymbol{x}}}\coloneqq({\boldsymbol{x}},x_{n+1})$. By applying Theorem 4.1, we obtain its inhomogeneous counterpart, as given below.
Corollary 4.1.
Suppose that Assumption 3.2 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}$. Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$, SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg\Sigma_{0}\leq 2(N_{\varepsilon}+d_{F})$ and $\deg_{{\boldsymbol{x}}}\Sigma\leq 2(N_{\varepsilon}+d_{F}-d_{G})$, such that
$\theta^{N_{\varepsilon}}(F({\boldsymbol{x}})+\varepsilon\theta^{d_{F}}I_{p})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Proof.
Let $\widetilde{F}\coloneqq[x_{n+1}^{2d_{F}}F_{ij}({\boldsymbol{x}}/x_{n+1})]_{i,j\in[p]}\in{\mathbb{S}}[\tilde{{\boldsymbol{x}}}]^{p}$, $\widetilde{G}\coloneqq[x_{n+1}^{2d_{G}}G_{ij}({\boldsymbol{x}}/x_{n+1},{\boldsymbol{y}})]_{i,j\in[q]}\in{\mathbb{S}}[\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}]^{q}$, and consider
$\widetilde{\mathcal{X}}\coloneqq\left\{\tilde{{\boldsymbol{x}}}\in{\mathbb{R}}^{n+1}\middle|\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
We first prove that $\widetilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}\succ 0$ on $\widetilde{\mathcal{X}}\setminus\{\mathbf{0}\}$. Fix a point $\tilde{{\boldsymbol{u}}}=({\boldsymbol{u}},u_{n+1})\in\widetilde{\mathcal{X}}\setminus\{\mathbf{0}\}$.
  1. -
    Case 1: $u_{n+1}\neq 0$. As $\widetilde{G}(\tilde{{\boldsymbol{u}}},{\boldsymbol{y}})=u_{n+1}^{2d_{G}}G({\boldsymbol{u}}/u_{n+1},{\boldsymbol{y}})\succeq 0$, we have $G({\boldsymbol{u}}/u_{n+1},{\boldsymbol{y}})\succeq 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, which implies ${\boldsymbol{u}}/u_{n+1}\in\mathcal{X}$. Hence, $\widetilde{F}(\tilde{{\boldsymbol{u}}})=u_{n+1}^{2d_{F}}F({\boldsymbol{u}}/x_{n+1})\succeq 0$. Since $\|\tilde{{\boldsymbol{u}}}\|^{2}\neq 0$, $\widetilde{F}(\tilde{{\boldsymbol{u}}})+\varepsilon\|\tilde{{\boldsymbol{u}}}\|^{2d_{F}}I_{p}\succ 0$.
  2. -
    Case 2: $u_{n+1}=0$. By the definition of $d_{F}$, $x_{n+1}$ divides $\widetilde{F}(\tilde{{\boldsymbol{x}}})$. Thus, $\widetilde{F}(\tilde{{\boldsymbol{u}}})=0$. Since $\|\tilde{{\boldsymbol{u}}}\|^{2}\neq 0$, $\widetilde{F}(\tilde{{\boldsymbol{u}}})+\varepsilon\|\tilde{{\boldsymbol{u}}}\|^{2d_{F}}I_{p}\succ 0$.
Applying Theorem 4.1 to the polynomial matrices $\tilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}$ and $\widetilde{G}$, we obtain $N_{\varepsilon}\in\mathbb{N}$ and SOS matrices $\widetilde{\Sigma}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\widetilde{\Sigma}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, which are homogeneous in ${\boldsymbol{x}}$ and $\deg\widetilde{\Sigma}_{0}=2N_{\varepsilon}+2d_{F}$, $\deg_{{\boldsymbol{x}}}\widetilde{\Sigma}=2N_{\varepsilon}+2d_{F}-2d_{G}$, such that
$\|\tilde{{\boldsymbol{x}}}\|^{2N_{\varepsilon}}(\widetilde{F}(\tilde{{\boldsymbol{x}}})+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p})=\widetilde{\Sigma}_{0}(\tilde{{\boldsymbol{x}}})+\int_{\mathcal{Y}}\left\langle\widetilde{\Sigma}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}),\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$ (5)
Then, letting $x_{n+1}=1$ in (5), we achieve the desired conclusion. ∎
4.2. Pólya-type Positivstellensatz
Pólya [38] proved that multiplying a positive homogeneous polynomial on the nonnegative orthant by some power of $\sum_{i=1}^{n}x_{i}$ yields a polynomial with nonnegative coefficients. A “positive” version of Pólya’s theorem was given by Powers and Reznick [39]. Dickinson and Povh [9] extended the result of Pólya [38] to provide a certificate for positive homogeneous polynomials on the intersection of the nonnegative orthant with a basic semialgebraic set.
We next employ Theorem 4.1 and similar techniques from [9] to present a Pólya-type Positivstellensatz for homogeneous polynomial matrices being PD on the intersection of $\mathcal{X}$ with the nonnegative orthant. Denote by ${\boldsymbol{e}}$ the column vector in ${\mathbb{R}}^{n}$ of all ones.
Theorem 4.2.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$ and ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The proof is postponed in Section 6.4.
In particular, when $\mathcal{Y}$ is a semialgebriac set defined by PMIs, the coefficient matrices $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})$ in Theorem 4.2 have SOS-structured representations.
Corollary 4.2.
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Moreover, suppose that
$\mathcal{Y}=\{{\boldsymbol{y}}\in{\mathbb{R}}^{m}\mid H_{1}({\boldsymbol{y}})\succeq 0,\ldots,H_{t}({\boldsymbol{y}})\succeq 0\},\ H_{i}({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{y}}]^{\ell_{i}},\ \ell_{i}\in\mathbb{N},\ i\in[t],$
and the Archimedean condition holds for the quadratic module $\mathcal{Q}(\mathbf{H})$ where $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}(\mathbf{H})$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$,
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Proof.
It follows from Theorem 4.2 and Scherer-Hol’s Positivestellensatz (Theorem 2.1). ∎
4.3. Lasserre-Netzer perturbative Positivstellensatz
Lasserre and Netzer [30] provided SOS approximations of polynomials that are nonnegative on $[-1,1]^{n}$ via simple high-degree perturbations. Motivated by their work, we next derive SOS-structured approximations for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ that is PSD on $\mathcal{X}\cap[-1,1]^{n}$ using high-degree perturbations.
4.3.1. Exponentially bounded matrix-valued moment problem
For a given $C>0$, Berg and Maserick [2] showed that a real sequence has representing measure supported on the hypercube $[-C,C]^{n}$ if and only if the sequence is exponentially bounded and the associated moment matrix is PSD. We next extend their results to characterize exponentially bounded matrix-valued sequence with a matrix-valued representing measure being supported on the set $\mathcal{X}\cap[-C,C]^{n}$. Beyond its intrinsic interest, we will use this result to establish a Lasserre-Netzer perturbative Positivstellensatz for polynomial matrices with UQs.
For a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$ and $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, let us recall the definition of the moment matrix $M(\mathbf{S})$ and localizing moment matrix $M(H\mathbf{S})$.
Definition 4.1.
Given a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, the associated moment matrix $M(\mathbf{S})$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}$ and the $({\boldsymbol{\alpha}},{\boldsymbol{\beta}})$-th block entry is $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$. For $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, the localizing matrix $M(H\mathbf{S})$ associated to $\mathbf{S}$ and $H$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}$ and the $({\boldsymbol{\alpha}},{\boldsymbol{\beta}})$-th block entry is $\sum_{{\boldsymbol{\gamma}}\in\mathrm{supp}(H)}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\gamma}}}\otimes H_{{\boldsymbol{\gamma}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$. For $d\in\mathbb{N}$, the $d$-th order moment matrix $M_{d}(\mathbf{S})$ (resp. localizing matrix $M_{d}(H\mathbf{S})$) is the submatrix of $M(\mathbf{S})$ (resp. $M(H\mathbf{S})$) whose block row and block column are both indexed by $\mathbb{N}^{n}_{d}$.
Similar to the scalar case in [18], we give the definition of localizing moment matrix associated with $(\mathbf{S},G,\nu)$, where we write $G({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}G_{{\boldsymbol{\gamma}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\gamma}}}$ with $G_{{\boldsymbol{\gamma}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
Definition 4.2.
Given a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ and a Borel measure $\nu$ on $\mathcal{Y}$, the localizing matrix $M^{\nu}(G{\mathbf{S}})$ associated to ${\mathbf{S}}$, $G$ and $\nu$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}\times\mathbb{N}^{m}$ and the $(({\boldsymbol{\alpha}},{\boldsymbol{\eta}}),({\boldsymbol{\beta}},{\boldsymbol{\xi}}))$-th block entry is
$\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)$
for all $({\boldsymbol{\alpha}},{\boldsymbol{\eta}}),({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}\times\mathbb{N}^{m}$. For $d,k\in\mathbb{N}$, the $(d,k)$-th order localizing matrix $M^{\nu}_{d,k}(G{\mathbf{S}})$ is the submatrix of $M^{\nu}(G{\mathbf{S}})$ whose block row and block column are both indexed by $\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}$.
Remark 4.1.
Throughout the paper, for a sequence indexed by $\mathbb{N}^{n}\times\mathbb{N}^{m}$ (resp., $\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}$), the indices are arranged according to the order of the exponents in the monomial basis $[{\boldsymbol{x}}]_{\infty}\otimes[{\boldsymbol{y}}]_{\infty}$ (resp., $[{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}$).
With the definition of the localizing matrix $M_{d,k}^{\nu}(G{\mathbf{S}})$, we immediately have the following result.
Proposition 4.1.
Given a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, for any SOS matrix $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d$ and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$,
$\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0$
if and only if $M^{\nu}_{d,k}(G{\mathbf{S}})\succeq 0$.
Proof.
For the polynomial matrix $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in\mathbf{S}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, there exists some $Q\in{\mathbb{S}}_{+}^{pq|\mathbb{N}^{n}_{d}||\mathbb{N}^{m}_{k}|}$ satisfying
$\Sigma({\boldsymbol{x}},{\boldsymbol{y}})$ $=(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq})^{\intercal}Q(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq}).$
Then, we have
$\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)=\langle M^{\nu}_{d,k}(G{\mathbf{S}}),Q\rangle,$
which implies the desired conclusion. ∎
We now provide the following result for the exponentially bounded matrix-valued moment problem, which generalizes the result of Berg and Maserick [2]. The proof of Theorem 4.3 is postponed to Section 6.5.
Theorem 4.3.
Suppose that Assumption 3.2 holds. Then for a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ and $C>0$, the following are equivalent:
  1. (i)
    $\mathbf{S}$ has a matrix-valued representing measure supported on $\mathcal{X}\cap[-C,C]^{n}$;
  2. (ii)
    $M(\mathbf{S})\succeq 0$, $M^{\nu}(G\mathbf{S})\succeq 0$ and there is a constant $C_{0}>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
In case that UQs are not present, we obtain the following corollary.
Corollary 4.3.
For a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ and $C>0$, the following are equivalent:
  1. (i)
    $\mathbf{S}$ admits a matrix-valued representing measure supported on
    $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\}\cap[-C,C]^{n},$
    where $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, $q\in\mathbb{N}$;
  2. (ii)
    $M(\mathbf{S})\succeq 0$, $M(H\mathbf{S})\succeq 0$ and there is a constant $C_{0}>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
4.3.2. Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs
Inspired by the result of Lasserre and Netzer [30], we next derive a Positivstellensatz to characterize a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ that is PSD on $\mathcal{X}\cap[-1,1]^{n}$ using high-degree perturbations. For $d\in\mathbb{N}$, let $\Theta_{d}\coloneqq 1+\sum_{i=1}^{n}x_{i}^{2d}$.
Let $d(G)\coloneqq\lceil\deg_{{\boldsymbol{x}}}G/2\rceil$. Given $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, let us consider the following optimization problem:
$r_{d,k}^{\star}\coloneqq\left\{\begin{aligned} \inf_{r,\Sigma_{0},\Sigma}&\ r\\ \text{s.t.}\,\,&\ F({\boldsymbol{x}})+r\Theta_{d}I_{p}=\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\ \text{are SOS},\\ &\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G),\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2k,\end{aligned}\right.$ (6)
and it dual problem reads as
$\sup_{\mathbf{S}}&\ -\mathscr{L}_{\mathbf{S}}(F)\\ \text{s.t.}&\ \mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}}\subset{\mathbb{S}}^{p},\ \mathscr{L}_{\mathbf{S}}(\Theta_{d}I_{p})\leq 1,\\ &\ M_{d}(\mathbf{S})\succeq 0,\ M_{d-d(G),k}^{\nu}(G\mathbf{S})\succeq 0.\end{aligned}\right.$ (7)
As the sequence of zero matrices is feasible to (7), it holds $r^{\star}_{d,k}\geq 0$ by the weak duality.
Assumption 4.1.
There are positive number $\lambda>0$, and open and bounded subsets $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ of $\mathcal{X}$ and $\mathcal{Y}$, respectively, such that $G({\boldsymbol{x}},{\boldsymbol{y}})\succeq\lambda I_{q}$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$.
The following proposition shows that if Assumption 4.1 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$, then the optimal value $r^{\star}_{d,k}\to 0$ as $d,k\to\infty$. Its proof is postponed to Section 6.6.
Proposition 4.2.
Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$. Then, for any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that $r_{d,k}^{\star}\leq\varepsilon$ for all $d\geq\ell_{1},k\geq\ell_{2}$.
We can now give the Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs.
Theorem 4.4.
Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$. For any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1},k\geq\ell_{2}$, it holds
$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$ and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$.
Proof.
By Proposition 4.2, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1}$ and $k\geq\ell_{2}$, there is a feasible point $(r,\Sigma_{0}^{\prime},\Sigma)$ of (6) with $r\leq\varepsilon$. We have
$F({\boldsymbol{x}})+r\Theta_{d}I_{p}=\Sigma_{0}^{\prime}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
with $\deg(\Sigma_{0}^{\prime})\leq 2d$, $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$, and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$. Then, for all $d\geq\ell_{1}$ and $k\geq\ell_{2}$, it holds
$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=F({\boldsymbol{x}})+r\Theta_{d}I_{p}+(\varepsilon-r)\Theta_{d}I_{p}=(\Sigma_{0}^{\prime}+(\varepsilon-r)\Theta_{d}I_{p})+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The conclusion then follows. ∎
In case that UQs are not present, we obtain the following corollary.
Corollary 4.4.
Let $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$ and $\mathcal{W}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\}.$ Suppose that there are positive number $\lambda>0$, and open and bounded subset $\mathcal{O}$ of $\mathcal{X}$ such that $H({\boldsymbol{x}})\succeq\lambda I_{q}$ on $\mathcal{O}$. If $F({\boldsymbol{x}})$ is PSD on $\mathcal{W}\cap[-1,1]^{n}$, then for any $\varepsilon>0$, there exists some $\ell\in\mathbb{N}$ such that for all $d\geq\ell$, it holds
$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=\Sigma_{0}+\left\langle\Sigma,G\right\rangle_{p}$
for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d-2d(G)}^{pq}$.
5. Applications to robust PMI constrained optimization
Verifying PMIs over a prescribed set has a wide range of applications in many fields. For instance, many control problems for systems of ordinary differential equations can be formulated as convex optimization problems with PMI constraints that must be satisfied over a specified portion of the state space [15, 16, 19, 40, 51]. These problems are typically formulated as follows:
$\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})\coloneqq P_{0}({\boldsymbol{x}})-\sum_{i=1}^{r}P_{i}({\boldsymbol{x}})\gamma_{i}\succeq 0,\ \forall{\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\},$
where ${\boldsymbol{c}}\in{\mathbb{R}}^{r}$, $P_{0},\ldots,P_{r}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
However, due to estimation errors or lack of information, the data of real-world problems often involve uncertainty. Therefore, ensuring the robustness of PMIs over a prescribed set under uncertainty is a critical issue, which could be formulated as the following robust optimization problem:
$\tau^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t. }\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\},$ (8)
where $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is closed.
Consequently, Positivstellensätze for polynomial matrices with UQs developed in this paper serve as a powerful mathematical tool for addressing this issue. Indeed, by leveraging the SOS-structured certificates provided by Positivstellensätze for positive definiteness of $P({\boldsymbol{x}},{\boldsymbol{\gamma}})$ over $\mathcal{X}$, we are able to establish converging hierarchies of SDP relaxations for the problem (8).
Concretely, if $\mathcal{X}$ is compact, then one could apply the Scherer-Hol-type Positivestellensatz (Theorem 3.1) to construct a hierarchy of SDP relaxations for (8):
$\tau_{k}\coloneqq\inf_{{\boldsymbol{\gamma}},\Sigma_{0},\Sigma}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg\Sigma_{0},\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2k,\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2k-\deg_{{\boldsymbol{x}}}G.\end{aligned}\right.$
It is clear that the sequence $(\tau_{k})_{k\in\mathbb{N}}$ is non-increasing upper bounds of $\tau^{\star}$. Assuming that the Slater condition holds for (8), i.e., there exists a point $\bar{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$ such that $P({\boldsymbol{x}},\bar{{\boldsymbol{\gamma}}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$, the convergence of $\tau_{k}\to\tau^{\star}$ as $k\to\infty$ is guaranteed under Assumptions 2.1 and 3.2 by Theorem 3.1.
Other Positivstellensätze presented in this paper can also be employed to derive the corresponding hierarchies of SDP relaxations for (8). The study of convergence of these hierarchies provides an intriguing avenue for future research. For instance, if $\mathcal{X}$ is non-compact, then one could apply the Putinar–Vasilescu-type Positivstellensatz for the inhomogeneous case (Corollary 4.1) with a fixed $\varepsilon>0$ to construct the following hierarchy of SDP relaxations for (8) ($d_{P}\coloneqq\max_{i,j\in[p]}\lfloor\deg_{{\boldsymbol{x}}}(P_{ij})/2\rfloor+1$):
$\tau_{k}(\varepsilon)\coloneqq\inf_{{\boldsymbol{\gamma}},\Sigma_{0},\Sigma}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \theta^{k}(P({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon\theta^{d_{P}}I_{p})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg\Sigma_{0},\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2(k+d_{P}),\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2(k+d_{P}-d_{G}).\end{aligned}\right.$
Assuming that $\tau^{\star}$ is attainable and Assumption 3.2 holds, by Corollary 4.1, there exists $\bar{k}\in\mathbb{N}$ such that $\tau_{k}(\varepsilon)\leq\tau^{\star}$ for all $k\geq\bar{k}$. How to bound the number $\bar{k}$ from above, and whether $\tau_{k}(\varepsilon)\to\tau^{\star}$ as $\varepsilon\to 0$ and $k\to\infty$ under certain conditions, could be interesting topics for further study.
6. Proofs
6.1. Proof of Theorem 3.1
Proof of Theorem 3.1.
As mentioned earlier, the proof proceeds in four steps: Step 1: Let $\mathscr{L}$ be the linear functional in Corollary 3.1. The linear functional $\mathscr{L}$ induces a symmetric bilinear form on the space ${\mathbb{R}}[{\boldsymbol{x}}]^{p}$ by
$(\psi,\phi)\mapsto\langle\psi,\phi\rangle_{\mathscr{L}}\coloneqq\frac{1}{2}\mathscr{L}(\psi\phi^{\intercal}+\phi\psi^{\intercal}),\quad\forall\psi,\phi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}.$
This form is PSD since $\psi\psi^{\intercal}\in\mathcal{Q}^{p}(G,\nu)$ for any $\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}$. Let
$\mathcal{N}\coloneqq\{\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}\mid\langle\psi,\psi\rangle_{\mathscr{L}}=0\}$
be the $\langle\cdot,\cdot\rangle_{\mathscr{L}}$-null vector space corresponding to $\mathscr{L}$. By the Cauchy-Schwarz inequality, $\mathcal{N}$ is a vector subspace of ${\mathbb{R}}[{\boldsymbol{x}}]^{p}$ and $\mathscr{L}$ induces an inner product on the quotient space ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$. Denote by $\mathcal{H}$ the Hilbert space completion of ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$ and $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ its inner product. Recall that by the proof of Lemma 3.1, there exists some $C>0$ such that $C-x_{i}^{2}\in\mathcal{Q}^{1}(G,\nu)$ for all $i\in[n]$. It implies that $(C-x_{i}^{2})\psi\psi^{\intercal}\in\mathcal{Q}^{p}(G,\nu)$ for any $\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}$. Hence,
$0\leq\langle x_{i}\psi+\mathcal{N},x_{i}\psi+\mathcal{N}\rangle_{\mathcal{H}}=\mathscr{L}(x_{i}^{2}\psi\psi^{\intercal})\leq C\mathscr{L}(\psi\psi^{\intercal}).$
Therefore, the multiplication $\hat{x}_{i}:\psi\mapsto x_{i}\psi$ defines a bounded operator on ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$ and can be extended to a bounded operator on $\mathcal{H}$. Fix an orthonormal basis of $\mathcal{H}$ and let $\widehat{A}_{i}$ be the corresponding representative of the multiplication by $x_{i}$ in $\mathcal{B}(\mathcal{H})$ with respect to this basis. Let $\widehat{\mathbf{A}}\coloneqq(\widehat{A}_{i})_{i\in[n]}$.
Denote by $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ the $p\times p$ matrix algebra over the $C^{\star}$-algebra $\mathcal{B}(\mathcal{H})$. That is, $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ consists of $p\times p$ matrices whose entries are in $\mathcal{B}(\mathcal{H})$. Then, $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ is a $C^{\star}$-algebra with involution $[B_{ij}]^{\star}=[B_{ji}^{\star}]$ where $B_{ij}\in\mathcal{B}(\mathcal{H})$. For $k\in[p]$, let ${\boldsymbol{v}}_{k}\in\mathcal{H}$ be the image of the constant polynomial vector $e_{k}\in{\mathbb{R}}^{p}$, where $\{e_{1},\ldots,e_{p}\}$ is the standard basis for ${\mathbb{R}}^{p}$. Define a linear functional $\Pi$ on $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ by
$\Pi(\mathbf{B})=\sum_{1\leq i,j\leq p}\langle B_{ij}{\boldsymbol{v}}_{j},{\boldsymbol{v}}_{i}\rangle_{\mathcal{H}},\ \ \forall\mathbf{B}=[B_{ij}]\in\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})).$
Then, $\Pi$ is a state on $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$. Note that for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, it holds that $\Pi(H(\widehat{\mathbf{A}}))=\mathscr{L}(H)$.
Step 2: By applying the GNS construction to $(\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})),\Pi)$, we obtain a Hilbert space $\mathcal{K}$ with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{K}}$, a representation $\pi:\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))\to\mathcal{B}(\mathcal{K})$ and a vector $\xi\in\mathcal{K}$ such that $\Pi(\mathbf{B})=\langle\pi(\mathbf{B})\xi,\xi\rangle_{\mathcal{K}}$ for all $\mathbf{B}\in\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$.
For each $i,j\in[p]$, denote by $E_{ij}$ the $p\times p$ matrix whose $(i,j)$-th entry is the identity operator and all other entries are the zero operator in $\mathcal{B}(\mathcal{H})$. For each $i\in[p]$, let $\mathcal{K}_{i}=\pi(E_{ii})(\mathcal{K})$. Then, the subspaces $\mathcal{K}_{i}$’s are all mutually orthogonal and isomorphic. In fact, $\pi(E_{ij})$ is a linear isomorphism between $\mathcal{K}_{j}$ and $\mathcal{K}_{i}$ with the inverse $\pi(E_{ji})$. Moreover,
$\mathcal{K}=1_{\mathcal{B}(\mathcal{K})}\mathcal{K}=\pi\left(\sum_{i=1}^{p}E_{ii}\right)\mathcal{K}=\bigoplus_{i=1}^{p}\mathcal{K}_{i}.$
Hence, each $\mathcal{K}_{i}$ is a Hilbert subspace of $\mathcal{K}$. Let us decompose $\xi$ as $\xi=\xi_{1}+\cdots+\xi_{p}$ where each $\xi_{i}\in\mathcal{K}_{i}$, i.e., $\xi_{i}=\pi(E_{ii})\xi$.
For $H({\boldsymbol{x}})=[H_{ij}({\boldsymbol{x}})]\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, it holds that
$\mathscr{L}(H)$ $=\Pi(H(\widehat{\mathbf{A}}))=\left\langle\pi\left(\sum_{i,j}H_{ij}(\widehat{\mathbf{A}})E_{ij}\right)\sum_{\ell=1}^{p}\xi_{\ell},\sum_{\ell=1}^{p}\xi_{\ell}\right\rangle_{\mathcal{K}}$ (9)
$=\left\langle\sum_{i,j}\left(\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{ij})\right)\sum_{\ell=1}^{p}\xi_{\ell},\sum_{\ell=1}^{p}\xi_{\ell}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{ij})\xi_{j},\xi_{i}\right\rangle_{\mathcal{K}}.$
For $i,j\in[p]$, let $\eta_{j}=\pi(E_{1i})\pi(E_{ij})\xi_{j}$ and $\eta_{i}=\pi(E_{1i})\xi_{i}$. Then, $\eta_{j},\eta_{i}\in\mathcal{K}_{1}$ and
$\mathscr{L}(H)$ $=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{i1})\eta_{j},\pi(E_{i1})\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\pi(E_{1i})\pi(E_{i1})\eta_{i}\right\rangle_{\mathcal{K}}$
$=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\pi(E_{11})\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}}.$
For each $i\in[p]$, let $\mathbf{A}=(A_{1},\ldots,A_{n})$ with $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ being the restriction of $\pi(\widehat{A}_{i}I_{p})$ to $\mathcal{K}_{1}$ and denote by $\langle\cdot,\cdot\rangle_{\mathcal{K}_{1}}$ the inner product on $\mathcal{K}_{1}$ induced from $\langle\cdot,\cdot\rangle_{\mathcal{K}}$. Then, we have
$\mathscr{L}(H)=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle H_{ij}(\mathbf{A})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}_{1}}.$ (10)
Step 3: For each $i\in[n]$, $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ is self-adjoint. Indeed, for every $\omega,\gamma\in\mathcal{K}_{1}$,
$\langle A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}=\langle\pi(\widehat{A}_{i}I_{p})\omega,\gamma\rangle_{\mathcal{K}}=\langle\omega,(\pi(\widehat{A}_{i}I_{p}))^{\star}\gamma\rangle_{\mathcal{K}}=\langle\omega,\pi((\widehat{A}_{i}I_{p})^{\star})\gamma\rangle_{\mathcal{K}}=\langle\omega,\pi(\widehat{A}_{i}I_{p})\gamma\rangle_{\mathcal{K}}=\langle\omega,A_{i}\gamma\rangle_{\mathcal{K}_{1}}.$
Moreover, $(A_{i})_{i\in[n]}$ are pairwise commute. In fact, for every $\omega,\gamma\in\mathcal{K}_{1}$,
$\langle A_{i}A_{j}\omega,\gamma\rangle_{\mathcal{K}_{1}}$ $=\langle\pi(\widehat{A}_{i}I_{p})\pi(\widehat{A}_{j}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle\pi(\widehat{A}_{i}\widehat{A}_{j}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}$
$=\langle\pi(\widehat{A}_{j}\widehat{A}_{i}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle\pi(\widehat{A}_{j}I_{p})\pi(\widehat{A}_{i}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle A_{j}A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}.$
For each $i\in[n]$, as $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ is self-adjoint, by the spectral theorem for bounded self-adjoint operators on a Hilbert space (see, e.g. [48, Theorem 5.1]), there exists a unique projection-valued measure $\mathcal{E}_{i}$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}})$ such that the support of $\mathcal{E}_{i}$ is contained in a compact interval, $A_{i}=\int_{{\mathbb{R}}}t\mathrm{d}\mathcal{E}_{i}(t)$ and
$\langle A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}=\int_{{\mathbb{R}}}t\mathrm{d}\langle\mathcal{E}_{i}(t)\omega,\gamma\rangle_{\mathcal{K}_{1}},\quad\forall\omega,\gamma\in\mathcal{K}_{1}.$
Moreover, as $(A_{i})_{i\in[n]}$ are pairwise commute, by [48, Theorem 4.10], we can define on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ the product measure $\mathcal{E}$ of $\mathcal{E}_{i},i\in[n]$, such that $\mathcal{E}\left(\prod_{i\in[n]}\mathcal{N}_{i}\right)=\prod_{i\in[n]}\mathcal{E}_{i}(\mathcal{N}_{i})$ for all $\mathcal{N}_{i}\in\mathfrak{B}({\mathbb{R}}),\ i\in[n].$ By the multidimensional spectral theorem (see, [48, Theorem 5.23] or [4, Sec. 6.5, Theorem 1]), $A_{i}=\int_{{\mathbb{R}}^{n}}x_{i}\mathrm{d}\mathcal{E}({\boldsymbol{x}})$, $i\in[n]$, and
$\left\langle\left(\prod_{i\in[n]}A_{i}^{\alpha_{i}}\right)\omega,\gamma\right\rangle_{\mathcal{K}_{1}}=\int_{{\mathbb{R}}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\langle\mathcal{E}({\boldsymbol{x}})\omega,\gamma\rangle_{\mathcal{K}_{1}},\quad\forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n},\ \omega,\gamma\in\mathcal{K}_{1}.$
Note that the support of the product measure $\mathcal{E}$ is compact since each $\mathcal{E}_{i}$ has compact support.
For $i,j\in[p]$, recall that the vectors $\eta_{i},\eta_{j}\in\mathcal{K}_{1}$. Define the real-valued measure $\phi_{ij}(\cdot)\coloneqq\langle\mathcal{E}(\cdot)\eta_{j},\eta_{i}\rangle_{\mathcal{K}_{1}}$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ and the symmetric matrix-valued measure $\Phi\coloneqq[\phi_{ij}]$. Then, each $\phi_{ij}$ has a compact support and for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, we have $\mathscr{L}(H)=\int_{{\mathbb{R}}^{n}}H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})$ by (10). Moreover, $\Phi$ is a PSD matrix-valued measure. In fact, for any $\mathcal{N}\in\mathfrak{B}({\mathbb{R}}^{n})$ and ${\boldsymbol{v}}=(v_{i})\in{\mathbb{R}}^{p}$, as $\mathcal{E}(\mathcal{N})$ is a projection operator on $\mathcal{K}_{1}$, we have
${\boldsymbol{v}}^{\intercal}\Phi(\mathcal{N}){\boldsymbol{v}}=\sum_{i,j}\phi_{ij}(\mathcal{N})v_{i}v_{j}=\sum_{i,j}\langle\mathcal{E}(\mathcal{N})v_{j}\eta_{j},v_{i}\eta_{i}\rangle_{\mathcal{K}_{1}}=\left\langle\mathcal{E}(\mathcal{N})\sum_{i}v_{i}\eta_{i},\mathcal{E}(\mathcal{N})\sum_{j}v_{j}\eta_{j}\right\rangle_{\mathcal{K}_{1}}\geq 0.$
Step 4: Now we show that for $i,j\in[p]$, the support of $\phi_{ij}$ is contained in $\mathcal{X}$. Suppose that ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$ and ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$. Then, there is a ball $\mathcal{O}_{1}\subset{\mathbb{R}}^{n}$ (resp., $\mathcal{O}_{2}\subset{\mathbb{R}}^{m}$) with a radius $\rho>0$ around ${\boldsymbol{u}}$ (resp., ${\boldsymbol{w}}$) such that $G({\boldsymbol{x}},{\boldsymbol{y}})\not\succeq 0$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$. Without loss of genericity, we may assume that there exists ${\boldsymbol{v}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}\leq-\delta$ on $2\mathcal{O}_{1}\times 2\mathcal{O}_{2}$. Define a continuous function $h^{(1)}({\boldsymbol{x}})$ (resp. $h^{(2)}({\boldsymbol{y}})$) on ${\mathbb{R}}^{n}$ (resp., ${\mathbb{R}}^{m}$) by $h^{(1)}({\boldsymbol{x}})=2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|$ (resp., $h^{(2)}({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$)for ${\boldsymbol{x}}\in 2\mathcal{O}_{1}$ (resp., ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$) and $h^{(1)}({\boldsymbol{x}})=0$ (resp., $h^{(2)}({\boldsymbol{y}})=0$) otherwise. Take a compact set $\mathcal{C}\subset{\mathbb{R}}^{n}$ containing $2\mathcal{O}_{1}$ and the support of $\phi_{ii}$, $i\in[p]$. By the Weierstrass theorem, there is a sequence of square polynomials $\{h^{2}_{\ell}\}_{\ell\in\mathbb{N}}\subset{\mathbb{R}}[{\boldsymbol{x}}]$ which converges to $h^{(1)}$ uniformly on $\mathcal{C}$. By Proposition 3.1, there exists a sequence of SOS polynomilas $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h^{(2)}$ in the $L^{2}$-norm. Hence,
$\lim_{\ell\to\infty}\lim_{k\to\infty}\sum_{i=1}^{p}\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$ (11)
$=$ $\lim_{\ell\to\infty}\lim_{k\to\infty}\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $\lim_{\ell\to\infty}\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})h^{(2)}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h^{(1)}({\boldsymbol{x}})h^{(2)}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}\left(\int_{\mathcal{Y}\cap 2\mathcal{O}_{2}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$\leq$ $\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}\left(\int_{\mathcal{Y}\cap\mathcal{O}_{2}}-\delta\rho(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$=$ $-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
$\leq$ $-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\int_{\mathcal{O}_{1}}\rho\mathrm{d}\phi_{ii}({\boldsymbol{x}})=-\delta\rho^{2}\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\phi_{ii}(\mathcal{O}_{1}).$
As $\mathcal{C}$ is compact, the second and third equalities in (11) can be verified by Dominated Convergence Theorem and the convergence of $\{\sigma_{k}\}_{k}$ to $h^{(2)}$ in the $L^{2}$-norm.
Note that for all $\ell,k\in\mathbb{N}$,
$\sum_{i=1}^{p}\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$ (12)
$=$ $\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}\left\langle h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\Phi({\boldsymbol{x}})$
$=$ $\mathscr{L}\left(\int_{\mathcal{Y}}\left\langle h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
where the last inequality is due to the fact that $h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}$ is an SOS and $\mathscr{L}(\mathcal{Q}^{p}(G,\nu))\subseteq{\mathbb{R}}_{+}$. Since $\mathrm{supp}(\nu)=\mathcal{Y}$, we have $\nu(\mathcal{Y}\cap\mathcal{O}_{2})>0$. Therefore, by combining (11) and (12), we have $\phi_{ii}(\mathcal{O}_{1})=0$ for all $i\in[p]$. As $\Phi$ is PSD, it implies that $\phi_{ij}(\mathcal{O}_{1})=0$ for all $i,j\in[p]$. Hence, the support of each $\phi_{ij}$ is contained in $\mathcal{X}$. Then, as $F$ is positive on $\mathcal{X}$ by assumption and $\Phi$ is PSD, we have
$0\geq\mathscr{L}(F)=\int_{\mathcal{X}}F({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})>0,$
a contradiction. ∎
6.2. Proof of Proposition 3.3
For a matrix $N\in{\mathbb{S}}^{q}$, denote by $\lambda_{\max}(N)$ and $\lambda_{\min}(N)$ the largest and smallest eigenvalues of $N$, respectively.
Proof of Proposition 3.3.
For each $j\in[s]$, let
$M\coloneqq\max\left\{|\lambda_{\max}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|\colon{\boldsymbol{x}}\in[-C,C]^{n},\ {\boldsymbol{y}}\in\mathcal{Y},\ j\in[s]\right\}.$
As $\mathcal{Y}$ is compact, the quantity $M$ is well-defined. For this $M$, we show that there exists some $\bar{k}\in\mathbb{N}$ such that (4) holds for all $k\geq\bar{k}$. Suppose on the contrary that for any $k\in\mathbb{N}$, there exists ${\boldsymbol{x}}^{(k)}\in[-C,C]^{n}$ such that
$F({\boldsymbol{x}}^{(k)})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}}^{(k)},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}}^{(k)},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\not\succ 0.$ (13)
As $[-C,C]^{n}$ is compact, without loss of generality, we may assume that $\lim_{k\to\infty}{\boldsymbol{x}}^{(k)}={\boldsymbol{x}}^{\star}$ for some ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}$. Next, we prove that there exists $k^{\prime}\in\mathbb{N}$ and a neighborhood $\mathcal{O}_{1}$ of ${\boldsymbol{x}}^{\star}$ such that (4) holds for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$, which yields a contradiction.
We first consider the case that ${\boldsymbol{x}}^{\star}\in\widehat{\mathcal{X}}$. As $F({\boldsymbol{x}})\succ 0$ on $\widehat{\mathcal{X}}$, there exists $\varepsilon>0$ and a neighborhood $\mathcal{O}_{1}$ of ${\boldsymbol{x}}^{\star}$ such that $F({\boldsymbol{x}})\succeq\varepsilon I_{p}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$. We now prove that there exsits $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$, it holds
$\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\leq\frac{\varepsilon}{2},$
for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Fix a pair of ${\boldsymbol{u}}\in[-C,C]^{n}$ and ${\boldsymbol{w}}\in\mathcal{Y}$. Let $\{\lambda^{(j)}_{i}\}_{i\in[q_{j}]}$ be the set of eigenvalues of $G_{j}({\boldsymbol{u}},{\boldsymbol{w}})$. Take the decomposition $G_{j}({\boldsymbol{u}},{\boldsymbol{w}})=Q_{j}D_{j}Q_{j}^{\intercal}$ where $D_{j}=\mbox{diag}(\lambda_{1}^{(j)},\ldots,\lambda^{(j)}_{q_{j}})$ and $Q_{j}\in{\mathbb{R}}^{q_{j}\times q_{j}}$ with $Q_{j}^{\intercal}Q_{j}=Q_{j}Q_{j}^{\intercal}=I_{q_{j}}$. Then,
[…]
Let $q^{\prime}\coloneqq\max_{j\in[s]}q_{j}$. Note that there exists $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$,
$\max\left\{\xi(1-\xi)^{2k}:\xi\in[0,1]\right\}\leq\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}.$
Hence, for any pair of ${\boldsymbol{x}}\in[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, from (14) we obtain that
$\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\leq M\sum_{\lambda_{i}^{(j)}>0}\frac{\lambda_{i}^{(j)}}{M}\left(1-\frac{\lambda_{i}^{(j)}}{M}\right)^{2k}\leq Mq_{j}\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}\leq\frac{\varepsilon}{2s\nu(\mathcal{Y})}$ (15)
for all $k\geq k^{\prime}$, where $k^{\prime}$ does not depends on the choice of ${\boldsymbol{x}}\in[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{Y}$. Therefore, for all ${\boldsymbol{x}}\in[-C,C]^{n}$ and $k\geq k^{\prime}$, we have
$\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\leq\sum_{j=1}^{s}\int_{\mathcal{Y}}\frac{\varepsilon}{2s\nu(\mathcal{Y})}\mathrm{d}\nu({\boldsymbol{y}})\leq\frac{\varepsilon}{2}.$
Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succeq\varepsilon I_{p}-\frac{\varepsilon}{2}I_{p}=\frac{\varepsilon}{2}I_{p}\succ 0.$
Next we consider the case that ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}\setminus\widehat{\mathcal{X}}$. There exists $j_{0}\in[s]$ and ${\boldsymbol{y}}^{(0)}\in\mathcal{Y}$ such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}}^{\star},{\boldsymbol{y}}^{(0)})<0$. By continuity, for some $\bar{\lambda}<0$, there exists a neighborhood $\mathcal{O}_{1}$ (resp., $\mathcal{O}_{2}$) of ${\boldsymbol{x}}^{\star}$ (resp., ${\boldsymbol{y}}^{(0)}$) such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))\leq\bar{\lambda}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}$. For any pair of ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}\cap\mathcal{Y}$, letting $\lambda_{\min}^{(j_{0})}\coloneqq\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))$, from (14) we obtain that
$\left\langle\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle$ $\leq\sum_{\lambda_{i}^{(j_{0})}>0}\lambda_{i}^{(j_{0})}+\sum_{\lambda_{i}^{(j_{0})}<0}\lambda_{i}^{(j_{0})}\left(1-\frac{\lambda_{i}^{(j_{0})}}{M}\right)^{2k}$
$\leq Mq^{\prime}+\lambda_{\min}^{(j_{0})}\left(1-\frac{\lambda_{\min}^{(j_{0})}}{M}\right)^{2k}\leq Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}.$
Since $\mathrm{supp}(\nu)=\mathcal{Y}$, we have $\nu(\mathcal{Y}\cap\mathcal{O}_{2})>0$. Note that for each $j\in[s]$, by (15), the maximal eigenvalue of
$\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}=\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle I_{p}$
could be uniformly bounded from above on $[-C,C]^{n}\times\mathcal{Y}$ for all $k\in\mathbb{N}$. Therefore, as $[-C,C]^{n}$ and $\mathcal{Y}$ are compact, there exists $R\in{\mathbb{R}}$ such that
$F({\boldsymbol{x}})-\sum_{j\neq j_{0}}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$-\int_{\mathcal{Y}\setminus\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})-(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1)I_{p}\succeq RI_{p}$
for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Since $11-\bar{\lambda}/M>$ 1, there exists $k^{\prime}\in\mathbb{N}$ such that $\bar{\lambda}\left(1-\bar{\lambda}/{M}\right)^{2k}<R/\nu(\mathcal{Y}\cap\mathcal{O}_{2})$ holds for all $k\geq k^{\prime}$. Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$\succeq$ $-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R)I_{p}$
$\succeq$ $\left(-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left(Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}\right)\mathrm{d}\nu({\boldsymbol{y}})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}$
$\succeq$ $\left(-\left(Mq^{\prime}+\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}\right)\nu(\mathcal{Y}\cap\mathcal{O}_{2})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}=I_{p},$
which completes the proof. ∎
Using Proposition 3.3, we can recover Theorem 2.1 in the case that $\mathcal{Y}$ is compact.
Lemma 6.1.
For any $h\in\mathcal{Q}^{1}(\mathbf{G},\nu)$ and SOS matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, we have $S({\boldsymbol{x}})h({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{G},\nu)$.
Proof.
Write $h=\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\langle\Sigma_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})$ where all $\sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$ are SOS. Then,
$Sh=S\sigma_{0}+\sum_{j=1}^{s}S\int_{\mathcal{Y}}\langle\Sigma_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})=S\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S\otimes\Sigma_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
As $S\sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $S\otimes\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq_{j}}$ are all SOS matrices, we have $Sh\in\mathcal{Q}^{p}(\mathbf{G},\nu)$. ∎
Corollary 6.1.
Suppose that $\mathcal{Y}$ is compact and Assumption 3.1 holds. If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\widehat{\mathcal{X}}$, then $F\in\mathcal{Q}^{p}(\mathbf{G},\nu)$.
Proof.
Let $C>0$ be given in Assumption 3.1. Then, by Proposition 3.3, there exists $M>0$ and $\bar{k}\in\mathbb{N}$ such that
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succ 0$
holds on the set $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid\|{\boldsymbol{x}}\|^{2}\leq C\}$ for all $k\geq\bar{k}$. By Scherer-Hol’s Positivestellensatz (Theorem 2.1), there exist SOS matrices $S_{0},S_{1}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ such that
$F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})=S_{0}({\boldsymbol{x}})+S_{1}({\boldsymbol{x}})(C-\|{\boldsymbol{x}}\|^{2}).$
By Lemma 6.1, $S_{1}({\boldsymbol{x}})(C-\|{\boldsymbol{x}}\|^{2})\in\mathcal{Q}^{p}(\mathbf{G},\nu)$ which implies that $F\in\mathcal{Q}^{p}(\mathbf{G},\nu)$. ∎
6.3. Proof of Theorem 4.1
Proof of Theorem 4.1.
Let
$\widetilde{\mathcal{X}}\coloneqq\left\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid 1-\|{\boldsymbol{x}}\|^{2}=0,\ G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
Then $F({\boldsymbol{u}})\succ 0$ for all ${\boldsymbol{u}}\in\widetilde{\mathcal{X}}$. By Corollary 3.2, it holds that
$F({\boldsymbol{x}})=\Sigma^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+H({\boldsymbol{x}})(1-\|{\boldsymbol{x}}\|^{2}),$
[…]
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{-\deg{F}}=\Sigma^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)+\int_{\mathcal{Y}}\left\langle\Sigma^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right),G\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Let
$k^{\prime}\coloneqq\max\left\{\deg{F},\ \deg(\Sigma^{\prime}_{0}),\ \deg_{{\boldsymbol{x}}}{G}+\deg_{{\boldsymbol{x}}}\Sigma^{\prime}\right\}.$
By assumption, $k^{\prime}$ is even. Multiplying the two sides of the last equality with $\|{\boldsymbol{x}}\|^{k^{\prime}}$ gives
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=\overline{\Sigma}_{0}\left({\boldsymbol{x}}\right)+\int_{\mathcal{Y}}\left\langle\overline{\Sigma}\left({\boldsymbol{x}},{\boldsymbol{y}}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (16)
where
$\overline{\Sigma}_{0}\coloneqq\Sigma^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)\|{\boldsymbol{x}}\|^{k^{\prime}}\ \ \text{and}\ \ \overline{\Sigma}\coloneqq\Sigma^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\|{\boldsymbol{x}}\|^{k^{\prime}-\deg_{{\boldsymbol{x}}}{G}}.$
Since $\Sigma^{\prime}$ is an SOS matrix and $k^{\prime}-\deg_{{\boldsymbol{x}}}{G}\geq\deg_{{\boldsymbol{x}}}\Sigma^{\prime}$, we have $\overline{\Sigma}=H^{\intercal}H$ with $H=H_{1}+H_{2}\|{\boldsymbol{x}}\|$ where $H_{1},H_{2}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times m}$ for some $\ell\in\mathbb{N}$, are homogeneous in ${\boldsymbol{x}}$ of degree $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2$ and $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2-1$, respectively. Thus,
$\overline{\Sigma}=H^{\intercal}H=(H_{1}+H_{2}\|{\boldsymbol{x}}\|)^{\intercal}(H_{1}+H_{2}\|{\boldsymbol{x}}\|)=\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right)\|{\boldsymbol{x}}\|.$
Similarly, there exist homogeneous polynomial matrices $H_{0,1}\in{\mathbb{R}}[x]^{\ell_{0}\times p}_{k^{\prime}/2}$ and $H_{0,2}\in{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}_{k^{\prime}/2-1}$ for some $\ell_{0}\in\mathbb{N}$ such that
$\overline{\Sigma}_{0}=\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)\|{\boldsymbol{x}}\|.$
Then, by (16), it holds that
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=$ $\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$+\left(\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\|{\boldsymbol{x}}\|.$
Since $\|{\boldsymbol{x}}\|$ is not a polynomial and the left hand side of the above equation is a polynomial matrix, we must have
$\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})=0.$
Then, letting $N\coloneqq(k^{\prime}-\deg{F})/2\in\mathbb{N}$, $\Sigma_{0}\coloneqq H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}$ and $\Sigma\coloneqq H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}$, we obtain the desired conclusion. ∎
6.4. Proof of Theorem 4.2
Let ${\boldsymbol{z}}=(z_{1},\ldots,z_{n})$ and ${\boldsymbol{z}}\circ{\boldsymbol{z}}=(z_{1}^{2},\ldots,z_{n}^{2})$. We first derive an intermediary result.
Theorem 6.1.
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ ($\deg(S_{0})=N+\deg F$ and $\deg_{{\boldsymbol{x}}}S=N+\deg F-\deg_{{\boldsymbol{x}}}G$), and satisfy that $S_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$ (resp., $S({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$) are SOS matrices in ${\boldsymbol{z}}$ (resp., ${\boldsymbol{z}}$ and ${\boldsymbol{y}}$), such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Proof.
Consider the polynomial matrices $F({\boldsymbol{z}}\circ{\boldsymbol{z}})\in{\mathbb{S}}[{\boldsymbol{z}}]^{p}$ and $G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{q}$. Then, by assumption and Theorem 4.1, there exists $N$ and polynomial matrices $H_{0}\in{\mathbb{R}}[{\boldsymbol{z}}]^{\ell_{0}\times p}$, $H\in{\mathbb{R}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell_{0},\ell\in\mathbb{N}$ which are homogeneous in ${\boldsymbol{z}}$, such that
$\|{\boldsymbol{z}}\|^{2N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})=(H_{0}^{\intercal}H_{0})({\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle(H^{\intercal}H)({\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Moreover, by Theorem 4.1, $\deg(H_{0})=N+\deg F$ and $\deg_{{\boldsymbol{x}}}(H)=N+\deg F-\deg_{{\boldsymbol{x}}}G$. Note that there are sets of polynomial matrices
$\{H_{0,{\boldsymbol{\alpha}}}:{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}\quad\text{and}\quad\{H_{{\boldsymbol{\alpha}}}:{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times p}$
which are homogeneous in ${\boldsymbol{x}}$, such that
$H_{0}({\boldsymbol{z}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{z}}\circ{\boldsymbol{z}}),\quad H({\boldsymbol{z}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}).$
Then,
$({\boldsymbol{e}}^{\intercal}({\boldsymbol{z}}\circ{\boldsymbol{z}}))^{N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})$
$=$ $\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
$+$ $\sum_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\{0,1\}^{n},{\boldsymbol{\alpha}}\neq{\boldsymbol{\beta}}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\beta}}})({\boldsymbol{z}}\circ{\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\{0,1\}^{n},{\boldsymbol{\alpha}}\neq{\boldsymbol{\beta}}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\beta}}})({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Comparing the even and odd terms in ${\boldsymbol{z}}$ in the above equation, we get
$({\boldsymbol{e}}^{\intercal}({\boldsymbol{z}}\circ{\boldsymbol{z}}))^{N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})$
$=$ $\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}({\boldsymbol{z}}\circ{\boldsymbol{z}})^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
It follows that the equality
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
holds for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}$. As +${\mathbb{R}}^{n}_{+}$ has interior points, the polynomial matrices on the left and right side of the above equality are identical. Letting $S_{0}\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})$ and $S\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})$, the conclusion follows. ∎
As an extension of Pólya’s result [38], Scherer and Hol [45] provided the following certificate for homogeneous polynomial matrices being positive on the nonnegative orthant.
Theorem 6.2.
[45, Theorem 3] Suppose that the polynomial matrix $P({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is homogeneous and $P({\boldsymbol{x}})\succeq\lambda I_{p}$ for some $\lambda>0$ on $\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$. Write $P({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(P)}P_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ with each $P_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ and let
$L(P)\coloneqq\max_{{\boldsymbol{\alpha}}\in\mathrm{supp}(P)}\frac{{\boldsymbol{\alpha}}!}{\deg(P)!}\|P_{{\boldsymbol{\alpha}}}\|,$
[…]
$N\geq\frac{\deg(P)(\deg(P)-1)L(P)}{2\lambda}-\deg(P),$
all coefficients of $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}P({\boldsymbol{x}})$ are PD.
We will use this theorem as another intermediary result. Let $D\coloneqq\max\left\{\deg(F),\deg_{{\boldsymbol{x}}}(G)\right\}$.
Lemma 6.2.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$ and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $\varepsilon>0$ such that the homogeneous polynomial matrix
$F_{\varepsilon}({\boldsymbol{x}})\coloneqq({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg(F)}F({\boldsymbol{x}})-\varepsilon\left(({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D}I_{p}+\int_{\mathcal{Y}}\left\langle({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G}I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$
for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$.
Proof.
By homogeneity, we only need to prove $F_{\varepsilon}\succ 0$ on $\overline{\mathcal{X}}\coloneqq{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}.$ That is, there exists $\varepsilon>0$ such that $F({\boldsymbol{x}})-\varepsilon\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$ for all ${\boldsymbol{x}}\in\overline{\mathcal{X}}$. Since $\overline{\mathcal{X}}$ is compact and $F({\boldsymbol{x}})\succ 0$ on +${\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, letting
$\varepsilon\coloneqq\frac{\min_{{\boldsymbol{x}}\in\overline{\mathcal{X}}}\lambda_{\min}(F({\boldsymbol{x}}))}{2\max_{{\boldsymbol{x}}\in\overline{\mathcal{X}}}|\lambda_{\max}\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)|},$ (17)
the conclusion follows. ∎
We now prove the following Pólya-type Positivstellensatz for polynomial matrices with UQs.
Proof of Theorem 4.2.
Let $F_{\varepsilon}({\boldsymbol{x}})$ be the polynomial matrix in Lemma 6.2 where $\varepsilon>0$ is defined in (17). Then, applying Theorem 6.1 to $F_{\varepsilon}({\boldsymbol{x}})$ yields $N_{1}\in\mathbb{N}$ and polynomial matrices $S^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$’s are SOS matrices, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{1}}F_{\varepsilon}({\boldsymbol{x}})=S^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
Then, by the definition of $F_{\varepsilon}({\boldsymbol{x}})$,
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg F+N_{1}}F({\boldsymbol{x}})=\Gamma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Gamma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (18)
where
$\Gamma_{0}({\boldsymbol{x}})\coloneqq\varepsilon({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D+N_{1}}I_{p}+S^{\prime}_{0}({\boldsymbol{x}}),\ \ \Gamma({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\varepsilon({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G+N_{1}}I_{pq}+S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}).$
Note that $\Gamma_{0}({\boldsymbol{x}})$ and $\Gamma({\boldsymbol{x}},{\boldsymbol{y}})$ are homogeneous in ${\boldsymbol{x}}$ because by Theorem 6.1,
$\deg S^{\prime}_{0}=N_{1}+\deg F_{\varepsilon}=D+N_{1}\ \ \text{and}\ \ \deg_{{\boldsymbol{x}}}S^{\prime}=N_{1}+\deg F_{\varepsilon}-\deg_{{\boldsymbol{x}}}G=D-\deg_{{\boldsymbol{x}}}G+N_{1}.$
Since $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$ are SOS matrices, for all ${\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$, it holds
$\Gamma_{0}({\boldsymbol{x}})\succeq\varepsilon I_{p}\ \text{ and }\ \Gamma({\boldsymbol{x}},{\boldsymbol{y}})\succeq\varepsilon I_{pq},\ \forall{\boldsymbol{y}}\in\mathcal{Y}.$
Write $\Gamma({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ and define
$L(\Gamma)\coloneqq\max_{{\boldsymbol{y}}\in\mathcal{Y}}\max_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\frac{{\boldsymbol{\alpha}}!}{\deg_{{\boldsymbol{x}}}(\Gamma)!}\|\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\|.$
As $\mathcal{Y}$ is compact, $L(\Gamma)$ is well-defined. Let
$N_{2}\coloneqq\max\left\{\frac{\deg(\Gamma_{0})(\deg(\Gamma_{0})-1)L(\Gamma_{0})}{2\varepsilon}-\deg(\Gamma_{0}),\ \frac{\deg_{{\boldsymbol{x}}}(\Gamma)(\deg_{{\boldsymbol{x}}}(\Gamma)-1)L(\Gamma)}{2\varepsilon}-\deg_{{\boldsymbol{x}}}(\Gamma)\right\}.$
Let $S_{0}({\boldsymbol{x}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma_{0}$ and $S({\boldsymbol{x}},{\boldsymbol{y}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma$. Write
$S_{0}({\boldsymbol{x}})=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\quad\text{ and }\quad S({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}.$
By Theorem 6.2, $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$ and ${\boldsymbol{y}}\in\mathcal{Y}$. Letting $N=D-\deg F+N_{1}+N_{2}$, the conclusion follows from (18). ∎
6.5. Proof of Theorem 4.3
To proof Theorem 4.3, we need some intermediary lemmas.
Lemma 6.3.
Assume that $M(\mathbf{S})\succeq 0$ and there are constants $C_{0},C>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. Then, $\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{\mathbf{0}}\|C^{|{\boldsymbol{\alpha}}|}\leq\mbox{tr}\left(S_{\mathbf{0}}\right)C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
Proof.
As $S_{\mathbf{0}}\succeq 0$, we only need to prove the first inequality. If $\|S_{\mathbf{0}}\|=0$, then $S_{\mathbf{0}}=0$ and hence $S_{{\boldsymbol{\alpha}}}=0$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$ since the null space of $S_{\mathbf{0}}$ contains that of $S_{{\boldsymbol{\alpha}}}$ by [31, Lemma 1.2 (i)]. As $M(\mathbf{S})\succeq 0$, by [17, Theorem 7.7.11], there is a contraction $N_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$ such that $S_{{\boldsymbol{\alpha}}}=S_{\mathbf{0}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\alpha}}}^{1/2}$. Then, it holds
$\|S_{{\boldsymbol{\alpha}}}\|=\|S_{\mathbf{0}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\alpha}}}^{1/2}\|\leq\|S_{\mathbf{0}}^{1/2}\|\|N_{{\boldsymbol{\alpha}}}\|\|S_{2{\boldsymbol{\alpha}}}^{1/2}\|\leq\|S_{\mathbf{0}}\|^{1/2}\|S_{2{\boldsymbol{\alpha}}}\|^{1/2}.$
By induction, we have
$\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{\mathbf{0}}\|^{\sum_{i=1}^{k}1/2^{i}}\|S_{2^{k}{\boldsymbol{\alpha}}}\|^{1/2^{k}}\leq\|S_{\mathbf{0}}\|^{1-1/2^{k}}(C_{0}C^{2^{k}|{\boldsymbol{\alpha}}|})^{1/2^{k}}=\|S_{\mathbf{0}}\|^{1-1/2^{k}}C_{0}^{1/2^{k}}C^{|{\boldsymbol{\alpha}}|}.$
for any $k\geq 1$. We obtain $\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{\mathbf{0}}\|C^{|{\boldsymbol{\alpha}}|}$ by letting $k\to\infty$. ∎
Lemma 6.4.
Suppose that Assumption 3.2 holds. For $C>0$, the set
$\mathscr{M}\coloneqq\{\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}\mid\mbox{tr}\left(S_{\mathbf{0}}\right)=1,\ M(\mathbf{S})\succeq 0,\ M^{\nu}(G\mathbf{S})\succeq 0,\ \|S_{{\boldsymbol{\alpha}}}\|\leq C^{|{\boldsymbol{\alpha}}|},\ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\}$
is a convex set whose extreme points are ${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}\zeta_{{\boldsymbol{u}}}$ where ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ with $\|{\boldsymbol{v}}\|=1$ and $\zeta_{{\boldsymbol{u}}}=({\boldsymbol{u}}^{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ is the Zeta vector at some ${\boldsymbol{u}}\in\mathcal{X}\cap[-C,C]^{n}$.
Proof.
It is clear that $\mathscr{M}$ is a convex set. Let $\mathbf{S}$ be an extreme point of $\mathscr{M}$. We next show that there exists a sequence $(\xi_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{R}}^{\mathbb{N}^{n}}$ such that $S_{{\boldsymbol{\alpha}}}=\xi_{{\boldsymbol{\alpha}}}S_{\mathbf{0}}$ and $\xi_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}\xi_{{\boldsymbol{\beta}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$.
Now we fix an arbitrary ${\boldsymbol{\alpha}}_{0}\in\mathbb{N}^{n}$ and prove that $S_{{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{\mathbf{0}}$ for some $\xi_{{\boldsymbol{\alpha}}_{0}}\in{\mathbb{R}}$. For $\varepsilon\in\{\pm 1\}$, define two sequences $\mathbf{S}^{(\varepsilon)}\coloneqq(S^{(\varepsilon)}_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ by letting $S^{(\varepsilon)}_{{\boldsymbol{\alpha}}}\coloneqq C^{|{\boldsymbol{\alpha}}_{0}|}S_{{\boldsymbol{\alpha}}}+\varepsilon S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
We claim that $M(\mathbf{S}^{(\varepsilon)})\succeq 0$ and $M^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$, that is, $M_{d}(\mathbf{S}^{(\varepsilon)})\succeq 0$ and $M_{d,k}^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$ for each $d,k\in\mathbb{N}$. We only prove that $M^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$ and that $M(\mathbf{S}^{(\varepsilon)})\succeq 0$ could be proven in a similar way. Fix an arbitrary vector ${\boldsymbol{w}}=(w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}}})_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}}\in{\mathbb{R}}^{pq|\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}|}$ with each $w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\in{\mathbb{R}}^{pq}$, we need to prove that
${\boldsymbol{w}}^{\intercal}M^{\nu}_{d,k}(G\mathbf{S}^{(\varepsilon)}){\boldsymbol{w}}$
$=$ $\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S^{(\varepsilon)}_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
$=$ $C^{|{\boldsymbol{\alpha}}_{0}|}\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
$+\varepsilon\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\alpha}}_{0}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}\geq 0.$
Define a new sequence ${\boldsymbol{z}}\coloneqq(z_{{\boldsymbol{\kappa}}})_{\kappa\in\mathbb{N}^{n}}\in{\mathbb{R}}^{\mathbb{N}^{n}}$ by letting
$z_{{\boldsymbol{\kappa}}}\coloneqq\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\kappa}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
for each ${\boldsymbol{\kappa}}\in\mathbb{N}^{n}$. Then to prove ${\boldsymbol{w}}^{\intercal}M^{\nu}_{d,k}(G\mathbf{S}^{(\varepsilon)}){\boldsymbol{w}}\geq 0$, it suffices to show that $-z_{0}C^{|{\boldsymbol{\kappa}}|}\leq z_{{\boldsymbol{\kappa}}}\leq z_{0}C^{|{\boldsymbol{\kappa}}|}$ for all ${\boldsymbol{\kappa}}\in\mathbb{N}^{n}$ and then let ${\boldsymbol{\kappa}}={\boldsymbol{\alpha}}_{0}$. To this end, we next show that $M({\boldsymbol{z}})\succeq 0$ and ${\boldsymbol{z}}$ is exponentially bounded, and then apply Lemma 6.3. For $\ell\in\mathbb{N}$ and ${\boldsymbol{v}}\coloneqq(v_{{\boldsymbol{\kappa}}})_{{\boldsymbol{\kappa}}\in\mathbb{N}^{n}_{\ell}}\in{\mathbb{R}}^{|\mathbb{N}^{n}_{\ell}|}$, we have
${\boldsymbol{v}}^{\intercal}M_{k}({\boldsymbol{z}}){\boldsymbol{v}}$
$=$ $\sum_{{\boldsymbol{\kappa}},{\boldsymbol{\delta}}\in\mathbb{N}^{n}_{\ell}}\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})v_{{\boldsymbol{\kappa}}}w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\kappa}}+{\boldsymbol{\delta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)v_{{\boldsymbol{\delta}}}w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
$=$ $\,({\boldsymbol{v}}\otimes{\boldsymbol{w}})^{\intercal}M_{d+\ell,k}^{\nu}(G\mathbf{S})({\boldsymbol{v}}\otimes{\boldsymbol{w}})\geq 0.$
Moreover,
$|z_{{\boldsymbol{\kappa}}}|$ $\leq\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\left|\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\right|\|w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\|\|w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}\|\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}\|S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\kappa}}}\|\|G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\|$
$\leq\left(\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\left|\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\right|\|w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\|\|w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}\|\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}\|G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\|C^{|{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}|}\right)C^{|{\boldsymbol{\kappa}}|}.$
Then, applying Lemma 6.3 to ${\boldsymbol{z}}$ yielding $-z_{0}C^{|{\boldsymbol{\kappa}}|}\leq z_{{\boldsymbol{\kappa}}}\leq z_{0}C^{|{\boldsymbol{\kappa}}|}$. Letting ${\boldsymbol{\gamma}}={\boldsymbol{\alpha}}_{0}$, we get ${\boldsymbol{w}}^{\intercal}M_{d,k}^{\nu}(G\mathbf{S}^{(\varepsilon)}){\boldsymbol{w}}\geq 0$, i.e., $M^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$. Similarly, we can prove that $M(\mathbf{S}^{(\varepsilon)})\succeq 0$. As $\|S^{(\varepsilon)}_{{\boldsymbol{\alpha}}}\|\leq(2C^{|{\boldsymbol{\alpha}}_{0}|})C^{|{\boldsymbol{\alpha}}|}$, applying Lemma 6.3 again yields that
$\|S^{(\varepsilon)}_{{\boldsymbol{\alpha}}}\|\leq\mbox{tr}\left(S^{(\varepsilon)}_{0}\right)C^{|{\boldsymbol{\alpha}}|}=(C^{|{\boldsymbol{\alpha}}_{0}|}+\varepsilon\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right))C^{|{\boldsymbol{\alpha}}|}$
for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. Additionally, as $M(\mathbf{S}^{(\varepsilon)})\succeq 0$, we have $\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)\geq 0$, i.e., $|\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right)|\leq C^{|{\boldsymbol{\alpha}}_{0}|}$.
Case 1: $\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)=0$ for some $\varepsilon\in\{\pm 1\}$. Without loss of generality, we may assume that $\mbox{tr}\left(S^{(-1)}_{\mathbf{0}}\right)=0$ which implies $S^{(-1)}_{{\boldsymbol{\alpha}}}=0$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$ since $M(\mathbf{S}^{(-1)})\succeq 0$. Letting ${\boldsymbol{\alpha}}=\mathbf{0}$, we get $S_{{\boldsymbol{\alpha}}_{0}}=C^{|{\boldsymbol{\alpha}}_{0}|}S_{\mathbf{0}}$. So, we can set $\xi_{{\boldsymbol{\alpha}}_{0}}=C^{|{\boldsymbol{\alpha}}_{0}|}$ in this case and it holds that $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{{\boldsymbol{\alpha}}}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
Case 2: Now we assume $\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)\neq 0$ for both $\varepsilon\in\{\pm 1\}$ which implies $\mathbf{S}^{(\varepsilon)}/\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)\in\mathscr{M}$. Then, we have a convex combination of two points in $\mathscr{M}$:
[…]
which implies $\mathbf{S}=\mathbf{S}^{(1)}/\mbox{tr}\left(S^{(1)}_{\mathbf{0}}\right)$ or $\mathbf{S}^{(-1)}/\mbox{tr}\left(S^{(-1)}_{\mathbf{0}}\right)$. Without loss of generality, we may assume that $\mathbf{S}=\mathbf{S}^{(1)}/\mbox{tr}\left(S^{(1)}_{\mathbf{0}}\right)$. Then, for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$,
$C^{|{\boldsymbol{\alpha}}_{0}|}S_{{\boldsymbol{\alpha}}}+S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}=S^{(1)}_{{\boldsymbol{\alpha}}}=\mbox{tr}\left(S^{(1)}_{\mathbf{0}}\right)S_{{\boldsymbol{\alpha}}}=C^{|{\boldsymbol{\alpha}}_{0}|}S_{{\boldsymbol{\alpha}}}+\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right)S_{{\boldsymbol{\alpha}}}.$
Letting ${\boldsymbol{\alpha}}={\mathbf{0}}$ and $\xi_{{\boldsymbol{\alpha}}_{0}}=\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right)$, we get $S_{{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{\mathbf{0}}$. Also, in this case, it holds that $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{{\boldsymbol{\alpha}}}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
Now we have shown that there exists a sequence $(\xi_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{R}}^{\mathbb{N}^{n}}$ such that $S_{{\boldsymbol{\alpha}}}=\xi_{{\boldsymbol{\alpha}}}S_{\mathbf{0}}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. Moreover, it is clear from the above arguments that $\xi_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}S_{\mathbf{0}}=S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}\xi_{{\boldsymbol{\beta}}}S_{\mathbf{0}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$ which implies $\xi_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}\xi_{{\boldsymbol{\beta}}}$. Since $S_{\mathbf{0}}\succeq 0$, decompose $S_{\mathbf{0}}$ as $S_{\mathbf{0}}=\sum_{i=1}^{t}\lambda_{i}{\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}$ for some $\lambda_{i}>0$ and ${\boldsymbol{v}}^{(i)}\in{\mathbb{R}}^{p}$, $i\in[t]$, with $\sum_{i=1}^{t}\lambda_{i}=1$ and $\|{\boldsymbol{v}}^{(i)}\|=1$. Letting ${\boldsymbol{u}}=(u_{1},\ldots,u_{n})$ with $u_{i}=\xi_{{\boldsymbol{e}}_{i}}$, we get $\mathbf{S}=\sum_{i=1}^{t}\lambda_{i}{\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}\zeta_{{\boldsymbol{u}}}$. We have ${\boldsymbol{u}}\in[-C,C]^{n}$ because $|\xi_{{\boldsymbol{e}}_{i}}|=|\mbox{tr}\left(S_{{\boldsymbol{e}}_{i}}\right)|\leq C^{|{\boldsymbol{e}}_{i}|}=C$.
It remains to show that ${\boldsymbol{u}}\in\mathcal{X}$ and $t=1$. First, suppose on the contrary that ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$. Then, there is a ball $\mathcal{O}_{2}\subset{\mathbb{R}}^{m}$ with a radius $\rho>0$ around ${\boldsymbol{w}}$ such that $G({\boldsymbol{u}},{\boldsymbol{y}})\not\succeq 0$ for all ${\boldsymbol{y}}\in\mathcal{O}_{2}$. Without loss of genericity, we may assume that there exists ${\boldsymbol{\xi}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\leq-\delta$ for all ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$. Define a continuous function $h^{(2)}({\boldsymbol{y}})$ on ${\mathbb{R}}^{m}$ by $h^{(2)}({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$ for ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$ and $h^{(2)}({\boldsymbol{y}})=0$ otherwise. By Proposition 3.1, there exists a sequence of SOS polynomilas $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h^{(2)}$ in the $L^{2}$-norm. Hence,
$\lim_{k\to\infty}\int_{\mathcal{Y}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}\cap 2\mathcal{O}_{2}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{\xi}}(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})$ (19)
$\leq$ $\int_{\mathcal{Y}\cap\mathcal{O}_{2}}-\delta\rho\mathrm{d}\nu({\boldsymbol{y}})=-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})<0.$
The last inequality is due to the fact that $\mathrm{supp}(\nu)=\mathcal{Y}$. Then for $k\in\mathbb{N}$ large enough, we have the following contradiction,
$0>\sum_{i=1}^{t}\lambda_{i}\|{\boldsymbol{v}}^{(i)}\|^{2}\int_{\mathcal{Y}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=$ $\sum_{i=1}^{t}\int_{\mathcal{Y}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\langle I_{p},\lambda_{i}{\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}\rangle$ (20)
$=$ $\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\left\langle\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{\xi}}{\boldsymbol{\xi}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
where the last inequality is due to the fact that $\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{\xi}}{\boldsymbol{\xi}}^{\intercal}$ is an SOS matrix and Proposition 4.1. As ${\boldsymbol{u}}\in\mathcal{X}\cap[-1,1]^{n}$, it is easy to see that each ${\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}\zeta_{{\boldsymbol{u}}}\in\mathscr{M}$. Then since $\mathbf{S}$ is an extreme point of $\mathscr{M}$, we must have $t=1$. ∎
Proof of Theorem 4.3.
(i) $\Longrightarrow$ (ii). Assume that $\mathbf{S}$ has a $p\times p$ matrix-valued representing measure $\Phi=[\phi_{ij}]_{i,j}$ supported on $\mathcal{X}\cap[-C,C]^{n}$. Then each measure $\phi_{ij}$ is supported on $\mathcal{X}\cap[-C,C]^{n}$. Fix $d\in\mathbb{N}$ and an arbitrary vector ${\boldsymbol{w}}\in{\mathbb{R}}^{p|\mathbb{N}^{n}_{d}|}$. Let $\Sigma_{0}({\boldsymbol{x}})\coloneqq([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}{\boldsymbol{w}}{\boldsymbol{w}}^{\intercal}([{\boldsymbol{x}}]_{d}\otimes I_{p})$. We have
${\boldsymbol{w}}^{\intercal}M_{d}(\mathbf{S}){\boldsymbol{w}}=\mathscr{L}_{\mathbf{S}}(\Sigma_{0})=\int_{\mathcal{X}}\mbox{tr}\left(\Sigma_{0}({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\right)\geq 0,$
which implies $M_{d}(\mathbf{S})\succeq 0$. Fix $d,k\in\mathbb{N}$ and an arbitrary PSD matrix $Z\in\mathbb{S}_{+}^{pq|\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}|}$. Let
$\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k})\otimes I_{pq})^{\intercal}Z(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k})\otimes I_{pq}).$
We have $\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succeq 0$ for any ${\boldsymbol{x}}\in\mathcal{X}\cap[-C,C]^{n}$. Hence,
$\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
which implies that $M_{d,k}^{\nu}(G\mathbf{S})\succeq 0$ by Proposition 4.1. Moreover, for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$,
$\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{{\boldsymbol{\alpha}}}\|_{F}=\left\|\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})\right\|_{F}=\left\|\left[\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi_{ij}\right]_{i,j}\right\|_{F}\leq\sqrt{\sum_{i,j=1}^{p}(\phi_{ij}(\mathcal{X}\cap[-C,C]^{n}))^{2}}C^{|{\boldsymbol{\alpha}}|}.$
(ii) $\Longrightarrow$ (i). Assume that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. By Lemma 6.3, $\|S_{{\boldsymbol{\alpha}}}\|\leq\mbox{tr}\left(S_{\mathbf{0}}\right)C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$. If $\mbox{tr}\left(S_{\mathbf{0}}\right)=0$, then $S_{\mathbf{0}}=0$ which implies that all $\mathbf{S}=0$ and we are done. Now rescale $\mathbf{S}$ if necessary and assume that $\mbox{tr}\left(S_{\mathbf{0}}\right)=1$. Then, by Lemma 6.3, $\mathbf{S}$ belongs the set $\mathscr{M}$ in Lemma 6.4. As the set $\mathscr{M}$ is compact by Tychonoff’s theorem, applying the Krein-Milman theorem [24] to the convex compact set $\mathscr{M}$ yields that $\mathbf{S}$ belongs to the closure of the convex hull of the extreme set of $\mathscr{M}$. Therefore, by Lemma 6.4, we have a coordinate-wise convergence $\mathbf{S}=\lim_{i\to\infty}\mathbf{S}^{(i)}$ where each $\mathbf{S}^{(i)}$ admits a finitely atomic matrix-valued measure supported by $\mathcal{X}\cap[-C,C]^{n}$. Then, by the coordinate-wise convergence and Haviland’s theorem (Theorem 2.2), we conclude that $\mathbf{S}$ has a matrix-valued representing measure supported by $\mathcal{X}\cap[-C,C]^{n}$. ∎
6.6. Proof of Proposition 4.2
We first show that (7) is solvable, and strong duality holds between (6) and (7) if Assumption 4.1 holds.
Lemma 6.5.
[…]
Proof.
Define a linear form $L:{\mathbb{R}}[{\boldsymbol{x}}]_{2d}\to{\mathbb{R}}$ by
$L(f)=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(f)}\mbox{tr}\left(S_{{\boldsymbol{\alpha}}}\right)f_{{\boldsymbol{\alpha}}},\ \ \forall f=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(f)}f_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{R}}[{\boldsymbol{x}}]_{2d}.$
Then it is clear that $L(f^{2})=\mathscr{L}_{\mathbf{S}}(f^{2}I_{p})$ for all $f\in{\mathbb{R}}[{\boldsymbol{x}}]_{d}$. As $M_{d}(\mathbf{S})\succeq 0$, we have $L(f^{2})\geq 0$ for all $f\in{\mathbb{R}}[{\boldsymbol{x}}]_{d}$. Moreover, since $\mathscr{L}_{\mathbf{S}}(\Theta_{d}I_{p})\leq 1$, it holds that $L(1)\leq 1$ and $L(x_{i}^{2d})\leq 1$ for all $i\in[n]$. Then, by [30, Lemmas 4.1 and 4.3], $|L({\boldsymbol{x}}^{{\boldsymbol{\alpha}}})|\leq 1$ and hence $|\mbox{tr}\left(S_{{\boldsymbol{\alpha}}}\right)|\leq 1$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}$. In particular, for any ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{d}$, as $S_{2{\boldsymbol{\alpha}}}\succeq 0$, we have $\|S_{2{\boldsymbol{\alpha}}}\|\leq\mbox{tr}\left(S_{2{\boldsymbol{\alpha}}}\right)\leq 1$. Fix an ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}$. We may write ${\boldsymbol{\alpha}}={\boldsymbol{\beta}}+{\boldsymbol{\gamma}}$ for some ${\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in\mathbb{N}^{n}_{d}$. As $M_{d}(\mathbf{S})\succeq 0$, by [17, Theorem 7.7.11], there is a contraction $N_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ such that $S_{{\boldsymbol{\alpha}}}=S_{2{\boldsymbol{\beta}}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\gamma}}}^{1/2}$. Therefore,
$\|S_{{\boldsymbol{\alpha}}}\|=\|S_{2{\boldsymbol{\beta}}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\gamma}}}^{1/2}\|\leq\|S_{2{\boldsymbol{\beta}}}^{1/2}\|\|N_{{\boldsymbol{\alpha}}}\|\|S_{2{\boldsymbol{\gamma}}}^{1/2}\|\leq\|S_{2{\boldsymbol{\beta}}}^{1/2}\|\|S_{2{\boldsymbol{\gamma}}}^{1/2}\|\leq\mbox{tr}\left(S_{2{\boldsymbol{\beta}}}\right)\mbox{tr}\left(S_{2{\boldsymbol{\gamma}}}\right)\leq 1.$
holds true. ∎
Proposition 6.1.
The problem (7) is solvable. If Assumption 4.1 holds, then there is no dual gap between (6) and (7).
Proof.
As the sequence of zero matrices is feasible to (7), by Lemma 6.5, the feasible set of (7) is compact and hence (7) is solvable.
Now let $\lambda$, $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ be as in Assumption 4.1. To prove strong duality, it suffices to show that (7) is strictly feasible. Let $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ be such that $\Phi=\mbox{diag}(\phi,\ldots,\phi)$ where $\phi$ is the probability measure with uniform distribution on $\mathcal{O}_{1}$, and ${\mathbf{S}}^{\circ}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}}$ where each
$S_{{\boldsymbol{\alpha}}}=\frac{1}{1+p\int_{\mathcal{X}}\Theta_{d}\mathrm{d}\phi({\boldsymbol{x}})}\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})=\left(\frac{1}{1+p\int_{\mathcal{X}}\Theta_{d}\mathrm{d}\phi({\boldsymbol{x}})}\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi({\boldsymbol{x}})\right)I_{p},\ \ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}.$
Clearly, $\mathscr{L}_{\mathbf{S}^{\circ}}(\Theta_{d}I_{p})<1$. We next prove that $M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ})\succ 0$. Since $\mathbf{S}^{\circ}$ has a PSD matrix-valued representing measure supported on $\mathcal{X}$, by Proposition 4.1, we obtain $M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ})\succeq 0$. Suppose on the contrary that $M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ})\not\succ 0$. Now fix a nonzero vector ${\boldsymbol{w}}=(w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}}})_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d-d(G)}\times\mathbb{N}^{m}_{k}}\in{\mathbb{R}}^{pq|\mathbb{N}^{n}_{d-d(G)}\times\mathbb{N}^{m}_{k}|}$ with each $w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\in{\mathbb{R}}^{pq}$ such that ${{\boldsymbol{w}}^{\intercal}}M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ}){\boldsymbol{w}}=0$. Let
$\Sigma({\boldsymbol{y}})=([{\boldsymbol{x}}]_{d-d(G)}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq})^{\intercal}{\boldsymbol{w}}{\boldsymbol{w}}^{\intercal}([{\boldsymbol{x}}]_{d-d(G)}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq}).$
Then it holds
$\mathscr{L}_{{\mathbf{S}}^{\circ}}\left(\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)=\langle M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ}),{\boldsymbol{w}}{\boldsymbol{w}}^{\intercal}\rangle={{\boldsymbol{w}}^{\intercal}}M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ}){\boldsymbol{w}}=0.$
Let
$T({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{d-d(G)}}\sum_{{\boldsymbol{\eta}}\in\mathbb{N}^{m}_{k}}w_{{\boldsymbol{\alpha}},{\boldsymbol{\eta}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\eta}}}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\quad\text{and}\quad\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}.$
For each $j\in[p]$, let
$H_{j}({\boldsymbol{x}},{\boldsymbol{y}})=[T_{(j-1)q+1}({\boldsymbol{x}},{\boldsymbol{y}}),\ldots,T_{jq}({\boldsymbol{x}},{\boldsymbol{y}})]\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}.$
Then,
$\left\langle\Sigma,G\right\rangle_{p}=[H_{i}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})]_{i,j\in[p]},$
and
$0=\mathscr{L}_{{\mathbf{S}}^{\circ}}\left(\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)=\frac{1}{1+p\int_{\mathcal{X}}\Theta_{d}\mathrm{d}\phi({\boldsymbol{x}})}\int_{\mathcal{X}}\sum_{j=1}^{p}\left(\int_{\mathcal{Y}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi({\boldsymbol{x}}).$
As $H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\geq 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, we have
$\int_{\mathcal{X}}\int_{\mathcal{Y}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})=0,\quad\forall j\in[p].$
As $G({\boldsymbol{x}},{\boldsymbol{y}})\succeq\lambda I_{p}$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$, for each $j\in[p]$, we have
00 0 $=\int_{\mathcal{X}}\int_{\mathcal{Y}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})\geq\int_{\mathcal{O}_{1}}\int_{\mathcal{O}_{2}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})$
$\geq\lambda\int_{\mathcal{O}_{1}}\int_{\mathcal{O}_{2}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})=\lambda\int_{\mathcal{O}_{1}}\int_{\mathcal{O}_{2}}\sum_{i=1}^{q}T_{(j-1)q+i}({\boldsymbol{x}},{\boldsymbol{y}})^{2}\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}}).$
Since $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are open, we have $T_{i}({\boldsymbol{x}},{\boldsymbol{y}})\equiv 0$ for each $i=1,\ldots,pq$. We then conclude ${\boldsymbol{w}}=0$, a contradiction. Similar arguments can be used to prove $M_{d}({\mathbf{S}}^{\circ})\succ 0$. Thus, $\mathbf{S}^{\circ}$ is strictly feasible to (7). ∎
Proof of Proposition 4.2.
Suppose that the conclusion is false. Then, we can find $d_{i},k_{i}\in\mathbb{N}$, $i\in\mathbb{N}$, such that $d_{i},k_{i}\to\infty$ as $i\to\infty$, and $r_{d_{i},k_{i}}^{\star}>\varepsilon$ for all $i\in\mathbb{N}$. By Proposition 6.1, let $\mathbf{S}^{(i)}=(S^{(i)}_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d_{i}}}$ be an optimal solution of (7) with $d=d_{i}$ and $k=k_{i}$. Hence, by Proposition 6.1, $\mathscr{L}_{\mathbf{S}^{(i)}}(F)=-r_{d_{i},k_{i}}^{\star}$. By Lemma 6.5, $\|S^{(i)}_{{\boldsymbol{\alpha}}}\|\leq 1$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d_{i}}$. Complete the sequence $\mathbf{S}^{(i)}$ with zero matrices to obtain $\widetilde{\mathbf{S}}^{(i)}=(S^{(i)}_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$. By Tychonoff’s Theorem, we could find a subsequence of $\{\mathbf{S}^{(i)}\}_{i\in\mathbb{N}}$ which pointwisely converges to some $\mathbf{S}^{\star}=(S^{\star}_{\boldsymbol{\alpha}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ in the product topology with each $\|S^{\star}_{\boldsymbol{\alpha}}\|\leq 1$. For the sake of simplicity and without loss of generality, we may assume that the whole sequence $\{\mathbf{S}^{(i)}\}_{i\in\mathbb{N}}$ converges to $\mathbf{S}^{\star}$. It holds that $M(\mathbf{S}^{\star})\succeq 0$ and $M^{\nu}(G\mathbf{S}^{\star})\succeq 0$ from the pointwise convergence. Then, by Theorem 4.3, $\mathbf{S}^{\star}$ admits a matrix-valued representing measure $\Phi^{\star}$ supported on $\mathcal{X}\cap[-1,1]^{n}$. Therefore, as $F({\boldsymbol{x}})$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$, we have
$-\varepsilon>-r_{d_{i},k_{i}}^{\star}=\mathscr{L}_{\mathbf{S}^{(i)}}(F)\to\mathscr{L}_{\mathbf{S}^{\star}}(F)=\int_{\mathcal{X}\cap[-1,1]^{n}}F\mathrm{d}\Phi^{\star}\geq 0,$
a contradiction. ∎
7. Conclusions
We extend various classical Positivstellensätze to provide SOS-structured characterizations for polynomial matrices that are positive (semi)definite over a semialgebraic set defined by a PMI with UQs. Under the Archimedean condition, we first present a Scherer-Hol-type Positivstellensatz, along with a sparse version for scalar polynomials being positive over the semialgebraic set in the presence of correlative sparsity. We also derive Putinar–Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze without assuming the Archimedean condition. These results significantly extend existing work on Positivstellensätze, and some of them remain novel and intriguing even in the absence of UQs. Our results are valuable for ensuring robustness of PMIs over a prescribed set with uncertainty, allowing potential applications in areas such as optimal control, systems theory, and certifying stability.
acknowledgements
Feng Guo was supported by the Chinese National Natural Science Foundation under grant 12471478. Jie Wang was supported by National Key R&D Program of China under grant No. 2023YFA1009401, the Strategic Priority Research Program of the Chinese Academy of Sciences XDB0640000 & XDB0640200, and the National Natural Science Foundation of China under grant No. 12201618 & 12171324.

### v4

@@
 Abstract.
-This paper studies Positivstellensätze for a polynomial matrix subject to polynomial matrix inequality constraints with universal quantifiers.
-We first present a Scherer-Hol-type Positivstellensatz under the Archimedean condition.
-When the objective is a scalar polynomial, we further provide a sparse Scherer-Hol-type Positivstellensatz in the presence of correlative sparsity.
-Next, without assuming the Archimedean condition, we derive Putinar-Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze under the same setting.
-These results can be viewed as common generalizations of corresponding Positivstellensätze in the cases of polynomials, polynomials with universal quantifiers, and polynomial matrices.
-For the proofs, techniques from $\star$-algebra, real algebraic geometry, operator theory, and convex optimization are employed.
-Applications of the established Positivstellensätze to robust polynomial matrix optimization are also discussed.
+This paper investigates Positivstellensätze for polynomial matrices subject to universally quantified polynomial matrix inequality constraints.
+We first establish a matrix-valued Positivstellensatz under the Archimedean condition, incorporating universal quantifiers.
+For scalar-valued polynomial objectives, we further develop a sparse Positivstellensatz that leverages correlative sparsity patterns within these quantified constraints.
+Moving beyond the Archimedean framework, we then derive two generalized Positivstellensätze under analogous settings.
+These results collectively unify and extend foundational theorems in three distinct contexts: classical polynomial Positivstellensätze, their universally quantified counterparts, and matrix polynomial formulations.
+Applications of the established Positivstellensätze to robust polynomial matrix inequality constrained optimization are also investigated.
 Key words and phrases:
@@
 Most Positivstellensätze achieve this by expressing the polynomial using sums of squares (SOS).
-These powerful results offer constructive methods to certify whether a polynomial $f$ is positive over a basic semialgebraic set defined by
+These powerful results offer constructive methods for certifying whether a polynomial $f$ is positive over a basic semialgebraic set defined by polynomial inequalities
 $\mathbf{K}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid g_{j}({\boldsymbol{x}})\geq 0,j=1,\ldots,s\},$
-and have widespread applications in areas like optimization, algebraic geometry, control theory, and more where polynomial positivity is a key concern [5, 14, 29, 31, 36].
+and have widespread applications in areas such as optimization, algebraic geometry, control theory, and more where polynomial positivity is a key concern [4, 14, 25, 26, 34].
 Depending on whether the semialgebraic set $\mathbf{K}$ is assumed to be compact, Positivstellensätze are divided into two categories.
 When $\mathbf{K}$ is a compact polyhedron with non-empty interior, Handelman’s Positivstellensatz [12] states that a polynomial $f$ that is positive on $\mathbf{K}$ can be expressed as a positive linear combination of cross-products of $g_{j}$’s.
-In the case of a more general compact basic semialgebraic set $\mathbf{K}$, Schmüdgen [47] demonstrated that if a polynomial $f$ is positive on $\mathbf{K}$, then it belongs to the preordering generated by $g_{j}$’s, that is, $f$ can be represented as an SOS-weighted combination of cross-products of $g_{j}$’s.
+In the case of a more general compact basic semialgebraic set $\mathbf{K}$, Schmüdgen [43] demonstrated that if a polynomial $f$ is positive on $\mathbf{K}$, then it belongs to the preorder generated by $g_{j}$’s, that is, $f$ can be represented as an SOS-weighted combination of cross-products of $g_{j}$’s.
 Putinar’s Positivstellensatz [39] provides an alternative representation that avoids the need for cross-products of $g_{j}$’s under the Archimedean condition (slightly stronger than compactness).
@@
 This framework generates a non-decreasing sequence of lower bounds converging to the optimal value of a given polynomial optimization problem.
-When $f$ is only nonnegative on $\mathbf{K}$, Lasserre and Netzer [30] provided an SOS approximation of $f$ via high-degree perturbations.
 There have been several attempts to provide certificates of positivity of a polynomial over a non-compact semialgebraic set.
-Without requiring compactness of $\mathbf{K}$, Krivine-Stengle Positivstellensatz [25, 50] states that a polynomial $f$ is positive over $\mathbf{K}$ if and only if $\psi f=1+\phi$ for some $\psi,\phi$ from the preordering generated by $g_{j}$’s.
-Pólya [38] demonstrated that if $f$ is homogeneous and positive on +${\mathbb{R}}^{n}_{+}\setminus\{\mathbf{0}\}$, then multiplying $f$ by some power of $\sum_{i=1}^{n}x_{i}$, one obtains a polynomial with nonnegative coefficients.
-Dickinson and Povh [9] generalized Pólya’s Positivstellensatz for homogeneous polynomials being positive on the intersection +${\mathbb{R}}^{n}_{+}\cap\mathbf{K}\setminus\{\mathbf{0}\}$, where $g_{j}$’s are assumed to be also homogeneous.
-Reznick [43] proved that after be multiplying by some power of $\sum_{i=1}^{n}x_{i}^{2}$, any positive definite (PD) form is a sum of even powers of linear forms.
-Putinar and Vasilescu [42] extended Reznick’s result to the constrained case where $f$ and $g_{j}$’s are homogeneous polynomial of even degree.
-When $\mathbf{K}$ is non-compact, SOS-structured certificates of positivity of $f$ over $\mathbf{K}$ could be also established by investigating various geometric objects associated with the data, e.g., gradient varieties [37], principal gradient tentacles [49], truncated tangency varieties [52], and the polar varieties [11].
+Without requiring compactness of $\mathbf{K}$, Krivine-Stengle Positivstellensatz [21, 46] states that a polynomial $f$ is positive over $\mathbf{K}$ if and only if $\psi f=1+\phi$ for some $\psi,\phi$ from the preorder generated by $g_{j}$’s.
+Pólya [35] demonstrated that if $f$ is homogeneous and positive on ${\mathbb{R}}^{n}_{+}\setminus\{\mathbf{0}\}$, then multiplying $f$ by some power of $\sum_{i=1}^{n}x_{i}$, one obtains a polynomial with nonnegative coefficients.
+Dickinson and Povh [7] generalized Pólya’s Positivstellensatz for homogeneous polynomials being positive on the intersection ${\mathbb{R}}^{n}_{+}\cap\mathbf{K}\setminus\{\mathbf{0}\}$, where $g_{j}$’s are assumed to be also homogeneous.
+Reznick [40] proved that, after multiplying by some power of $\sum_{i=1}^{n}x_{i}^{2}$, any positive definite (PD) form is a sum of even powers of linear forms.
+Putinar and Vasilescu [38] extended Reznick’s result to the constrained case where $f$ and $g_{j}$’s are homogeneous polynomials of even degree.
+When $\mathbf{K}$ is non-compact, SOS-structured positivity certificates of $f$ over $\mathbf{K}$ could be also established by investigating various geometric objects associated with the data, e.g., gradient varieties [33], principal gradient tentacles [45], truncated tangency varieties [48], and the polar varieties [10].
 Recently, Hu, Klep, and Nie [16] studied Positivstellensätze concerning semialgebraic sets defined by universal quantifiers (UQ).
@@
 They also investigated the corresponding moment problem on the semialgebraic set $\mathcal{U}$.
-As an important application, their results could be used to solve semi-infinite optimization problems which are highly challenge.
+As an important application, their results could be used to solve semi-infinite optimization problems which are highly challenging.
 Most Positivstellensätze could be generalized to the matrix setting that both $f$ and $g_{j}$’s are polynomial matrices.
 Scherer and Hol [42] developed a matrix-version of Putinar’s Positivstellensatz.
-Cimprič [7] extended Krivine-Stengle Positivstellensatz to the case of polynomial matrices with polynomial constraints.
+Cimprič [6] extended the Krivine-Stengle Positivstellensatz to the case of polynomial matrices with polynomial constraints.
 A matrix version of Handelman’s Positivstellensatz was proposed in [27].
@@
 Computing SOS-structured representations involved in Positivstellensätze could be typically cast as semidefinite programs (SDP).
-However, the size of the SDPs grows rapidly with the problem dimension.
-Hence from the perspective of computation, it becomes appealing to develop sparse versions of Positivstellensätze for sparse data, e.g., correlative sparsity [21, 27, 53], term sparsity [21, 33, 54, 55, 56], and matrix (chordal) sparsity [57, 58].
-Due to estimation errors or lack of information, the data of real-world problems often involve uncertainty.
+However, the size of the SDPs grows rapidly with the dimension of the problem.
+Hence, from the perspective of computation, it becomes appealing to develop sparse versions of Positivstellensätze for sparse data, e.g., correlative sparsity [32, 23, 49], term sparsity [32, 29, 51, 52, 50], and matrix (chordal) sparsity [54, 53].
+Due to estimation errors or lack of information, the data from real-world problems often involve uncertainty.
 As a result, ensuring the robustness of PMIs over a prescribed set with uncertainty is crucial for some safety-critical applications with little tolerance for failure [3].
@@
 Pólya Pólya [35] Theorem 4.2 Theorem 4.2 Theorem 4.2
-Lasserre-Netzer Lasserre and Netzer [30] Theorem 4.4 Theorem 4.4 Theorem 4.4
 Table 2. Summary of sparse Positivstellensätze for polynomials, polynomial matrices, and polynomial matrices with UQs
@@
 Contributions.
-We generalize several classical Positivstellensätze for scalar polynomials to the matrix setting with UQs (see Tables 1 and 2 for summaries).
+We generalize several classical Positivstellensätze from scalar polynomials to the matrix setting with UQs (see Tables 1 and 2 for summaries).
 In the following, we highlight the main results of this paper.
@@
 $\int_{\mathcal{Y}}H({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\coloneqq\sum_{{\boldsymbol{\beta}}\in\mbox{supp}_{{\boldsymbol{y}}}(H)}H_{{\boldsymbol{\beta}}}({\boldsymbol{x}})\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\mathrm{d}\nu({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}.$
-The main foundational result of this paper is the following Scherer-Hol-type Positivestellensatz which holds under the Archimedean condition (Assumption 3.1) and the Carleman condition on $\nu$ (Assumption 3.2).
+The main result of this paper is the following matrix-valued Positivestellensatz incorporating universal quantifiers, which holds under the Archimedean condition (Assumption 3.1) and the Carleman condition on $\nu$ (Assumption 3.2).
+See Section 2 for the definition of the product $\langle\cdot,\cdot\rangle_{p}$.
 Theorem A.
 (Theorem 3.1) Suppose that Assumptions 3.1 and 3.2 hold.
-If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then there exists SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ such that
-$F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-See Section 2.1 for the definition of the product $\langle\cdot,\cdot\rangle_{p}$.
-We adopt a notably nontrivial operator-theoretic approach to prove Theorem A, utilizing techniques such as the Gelfand-Naimark-Segal (GNS) construction for a carefully constructed $\star$-algebra, and the spectral theory of self-adjoint operators in a Hilbert space.
+If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}$, then there exist SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ such that
+$F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
 Then we consider the particular case of $p=1$, namely, the objective is a scalar polynomial.
@@
 We assume the presence of correlative sparsity in the problem data, which implies that the variables $\boldsymbol{x}$} decompose as a union of subsets ${\boldsymbol{x}}=\cup_{\ell=1}^{t}{\boldsymbol{x}}(\mathcal{I}_{\ell})$ such that $\{G_{j}\}_{j=1}^{s}=\cup_{\ell=1}^{t}\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, $\ell=1,\ldots,t$ (see Assumption 3.3).
-Under these conditions, we could give the following sparse version of the Scherer-Hol-type Positivstellensatz.
+Under these conditions, we could give the following sparse Positivstellensatz that leverages correlative sparsity patterns within these quantified constraints.
 Theorem B.
 (Theorem 3.2) Suppose that $\mathcal{Y}$ is compact, Assumption 3.3 holds for $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\mathcal{X}$, and Assumption 3.1 holds with respect to each ${\boldsymbol{x}}(\mathcal{I}_{\ell})$.
-If $f>0$ on $\mathcal{X}$, then there exist SOS polynomials $\sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}],\sigma_{\ell,j}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]$ such that
-$f({\boldsymbol{x}})=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}})+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\langle\sigma_{\ell,j}({\boldsymbol{x}},{\boldsymbol{y}}),G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\rangle\mathrm{d}\nu({\boldsymbol{y}})\right).$
-Next, building on Theorem A and existing techniques, we establish several Positivstellensätze for polynomial matrices with UQs and without assuming the Archimedean condition.
-The first such Positivstellensatz is of Putinar–Vasilescu-type.
+If $f>0$ on $\mathcal{X}$, then there exist SOS polynomials $\sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$, and SOS polynomial matrices $S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}}$, $j\in\mathcal{J}_{\ell}$, such that
+$f({\boldsymbol{x}})=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}}(\mathcal{I}_{\ell}))+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\langle S_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}),G_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}})\rangle\mathrm{d}\nu({\boldsymbol{y}})\right).$
+Next, building on Theorem A and existing techniques, we establish two generalized Positivstellensätze for polynomial matrices with UQs and without assuming the Archimedean condition.
+The first is a Positivstellensatz for non-compact case.
 Theorem C.
 (Theorem 4.1) Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$.
-Then, there exists $N\in\mathbb{N}$ and SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and $\deg\Sigma_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}\Sigma=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
-$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-The second such Positivstellensatz is of Pólya-type.
+Then, there exists $N\in\mathbb{N}$ and SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and $\deg S_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}S=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
+$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+The second is a Positivstellensatz on the nonnegative orthant.
 Theorem D.
 (Theorem 4.2) Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact.
-Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{{\boldsymbol{\alpha}}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy each $S_{0,{\boldsymbol{\alpha}}}\succ 0$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, such that
-$\left(\sum_{i=1}^{n}x_{i}\right)^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-The third such Positivstellensatz is of Lasserre-Netzer-type.
-Theorem E.
-(Theorem 4.4) Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is positive semidefinite (PSD) on $\mathcal{X}\cap[-1,1]^{n}$.
-Then for any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1},k\geq\ell_{2}$, it holds
-$F({\boldsymbol{x}})+\varepsilon\left(1+\sum_{i=1}^{n}x_{i}^{2d}\right)I_{p}=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
-for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg(\Sigma_{0})\leq 2d$, $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$, and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$.
-To prove Theorem E, we studied the exponentially bounded matrix-valued moment problem, providing a characterization for an exponentially bounded matrix-valued sequence to admit a matrix-valued representing measure supported on the set $\mathcal{X}\cap[-C,C]^{n}$ for a given $C>0$ (Theorem 4.3).
-This result extends the work of Berg and Maserick [2] to the matrix case and offers novel insights on its own.
+Then, there exists $N\in\mathbb{N}$ and polynomial matrices $P_{0}=\sum_{{\boldsymbol{\alpha}}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $P=\sum_{{\boldsymbol{\alpha}}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and satisfy each $P_{0,{\boldsymbol{\alpha}}}\succ 0$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all ${\boldsymbol{y}}\in\mathcal{Y}$, such that
+$\left(\sum_{i=1}^{n}x_{i}\right)^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+The inhomogeneous counterparts of Theorems C and D appear in Corollaries 4.1 and 4.3, respectively.
+As applications, leveraging the SOS-structured certificates provided by the established Positivstellensätze, we propose hierarchies of SDP relaxations for the robust PMI constrained optimization problem and analyze their convergence.
 The rest of this paper is organized as follows.
 In Section 2, we review some preliminary concepts.
-Section 3 presents a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs under the Archimedean condition, along with a sparse version in the presence of correlative sparsity.
-In Section 4, without relying on the Archimedean condition, we derive Putinar–Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze for polynomial matrices with UQs.
-Section 5 briefly discusses applications of the established Positivstellensätze to robust polynomial matrix optimization.
+Section 3 presents a matrix-valued Positivstellensatz for polynomial matrices with UQs under the Archimedean condition, along with a sparse version in the presence of correlative sparsity.
+In Section 4, without relying on the Archimedean condition, we derive two generalized Positivstellensätze for polynomial matrices with UQs.
+In Section 5, we apply the established Positivstellensätze to robust PMI constrained optimization.
 For the sake of readability, some lengthy and technical proofs are deferred to Section 6.
@@
 2. Preliminaries
-We collect some notation and basic concepts which will be used in this paper.
-We denote by ${\boldsymbol{x}}$ (resp., ${\boldsymbol{y}}$) the $n$-tuple (resp., $m$-tuple) of variables $(x_{1},\ldots,x_{n})$ (resp., $(y_{1},\ldots,y_{m})$).
-The symbol $\mathbb{N}$ (resp., ${\mathbb{R}}$, ${\mathbb{R}}_{+}$) denotes the set of nonnegative integers (resp., real numbers, nonnegative real numbers).
-For positive integer $n\in\mathbb{N}$, denote by $[n]$ the set $\{1,\ldots,n\}$.
-Denote by ${\mathbb{R}}^{p}$ (resp.
-${\mathbb{R}}^{l_{1}\times l_{2}}$, ${\mathbb{S}}^{p}$, ${\mathbb{S}}_{+}^{p}$) the $p$-dimensional real vector (resp.
-$l_{1}\times l_{2}$ real matrix, $p\times p$ symmetric real matrix, $p\times p$ PSD matrix) space.
-Denote by +${\mathbb{R}}^{n}_{+}$ the nonnegative orthant of ${\mathbb{R}}^{n}$.
-For ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ (resp., $N\in{\mathbb{R}}^{l_{1}\times l_{2}}$), the symbol ${\boldsymbol{v}}^{\intercal}$ (resp., $N^{\intercal}$) denotes the transpose of ${\boldsymbol{v}}$ (resp., $N$).
-For a matrix $N\in{\mathbb{R}}^{p\times p}$, $\mbox{tr}\left(N\right)$ denotes its trace.
-For two matrices $N_{1}$ and $N_{2}$, $N_{1}\otimes N_{2}$ denotes the Kronecker product of $N_{1}$ and $N_{2}$.
-For two matrices $N_{1}$ and $N_{2}$ of the same size, $\langle N_{1},N_{2}\rangle$ denotes the inner product $\mbox{tr}\left(N_{1}^{\intercal}N_{2}\right)$ of $N_{1}$ and $N_{2}$.
-The notation $I_{p}$ denotes the $p\times p$ identity matrix.
-For any $t\in{\mathbb{R}}$, $\lceil t\rceil$ (resp., $\lfloor t\rfloor$) denotes the smallest (resp., largest) integer that is not smaller (resp., larger) than $t$.
-For ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$, $\|{\boldsymbol{u}}\|$ denotes the standard Euclidean norm of ${\boldsymbol{u}}$.
-For $N\in{\mathbb{R}}^{l_{1}\times l_{2}}$, $\|N\|$ denotes the spectral norm of $N$.
-For a vector ${\boldsymbol{\alpha}}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}^{n}$, let $|{\boldsymbol{\alpha}}|=\alpha_{1}+\cdots+\alpha_{n}$.
-For a set $A$, we use $|A|$ to denote its cardinality.
-For $k\in\mathbb{N}$, let $\mathbb{N}^{n}_{k}\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid|{\boldsymbol{\alpha}}|\leq k\}$ and $|\mathbb{N}^{n}_{k}|=\binom{n+k}{k}$ be its cardinality.
-For variables ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$, ${\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ denotes the monomial $x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$.
-Let ${\mathbb{R}}[{\boldsymbol{x}}]$ (resp.
-${\mathbb{S}}[{\boldsymbol{x}}]^{p}$) denote the set of real polynomials (resp.
-$p\times p$ symmetric real polynomial matrices) in ${\boldsymbol{x}}$.
-For $h\in{\mathbb{R}}[{\boldsymbol{x}}]$ (resp., $h\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]$), we denote by $\deg(h)$ (resp., $\deg_{{\boldsymbol{x}}}(h)$) its total degree in ${\boldsymbol{x}}$.
-For a polynomial matrix $T({\boldsymbol{x}})=[T_{ij}({\boldsymbol{x}})]$ (resp., $T({\boldsymbol{x}},{\boldsymbol{y}})=[T_{ij}({\boldsymbol{x}},{\boldsymbol{y}})]$), denote $\deg(T)\coloneqq\max_{i,j}\deg(T_{ij})$ (resp., $\deg_{{\boldsymbol{x}}}(T)\coloneqq\max_{i,j}\deg_{{\boldsymbol{x}}}(T_{ij})$).
-For $k\in\mathbb{N}$, denote by ${\mathbb{R}}[{\boldsymbol{x}}]_{k}$ (resp., ${\mathbb{S}}[{\boldsymbol{x}}]^{p}_{k}$) the subset of ${\mathbb{R}}[{\boldsymbol{x}}]$ (resp., ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$) of degree up to $k$.
-For any $P({\boldsymbol{x}})=[P_{ij}({\boldsymbol{x}})]\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $Q({\boldsymbol{x}},{\boldsymbol{y}})=[Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})]\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$, denote
+[…]
 $\mathrm{supp}(P)$ $\coloneqq\{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\ \text{appears in some }P_{ij}({\boldsymbol{x}})\},$
@@
 $\mbox{supp}_{{\boldsymbol{y}}}(Q)$ $\coloneqq\{{\boldsymbol{\beta}}\in\mathbb{N}^{m}\mid{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\ \text{appears in some }Q_{ij}({\boldsymbol{x}},{\boldsymbol{y}})\text{ for some }{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\}.$
-2.1. SOS matrices and positivstellensätz for polynomial matrices
+Let us recall some background on SOS matrices and positivstellensätz for polynomial matrices.
 For a polynomial $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$, if there exist polynomials $f_{1}({\boldsymbol{x}}),\ldots,f_{t}({\boldsymbol{x}})$ such that $f({\boldsymbol{x}})=\sum_{i=1}^{t}f_{i}({\boldsymbol{x}})^{2}$, then we call $f({\boldsymbol{x}})$ an SOS.
-A polynomial matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is said to be an SOS matrix if there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$.
-For $d\in\mathbb{N}$, denote by $[{\boldsymbol{x}}]_{d}$ the canonical basis of ${\mathbb{R}}[{\boldsymbol{x}}]_{d}$, i.e.,
-$[{\boldsymbol{x}}]_{d}\coloneqq[1,\ x_{1},\ x_{2},\ \cdots,\ x_{n},\ x_{1}^{2},\ x_{1}x_{2},\ \cdots,\ x_{n}^{d}]^{\intercal},$ (2)
-whose cardinality is $|\mathbb{N}^{n}_{d}|=\binom{n+d}{d}$. With $d=\deg(T)$, we can write $T({\boldsymbol{x}})$ as
+A polynomial matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is said to be an SOS matrix if there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$.
+For $d\in\mathbb{N}$, denote by ${\boldsymbol{x}}]_{d}$ the canonical basis of ${\mathbb{R}}[{\boldsymbol{x}}]_{d}$, i.e.,
+${\boldsymbol{x}}]_{d}\coloneqq[1,\ x_{1},\ x_{2},\ \cdots,\ x_{n},\ x_{1}^{2},\ x_{1}x_{2},\ \cdots,\ x_{n}^{d}]^{\intercal},$ (2)
+whose cardinality is $|\mathbb{N}^{n}_{d}|=\binom{n+d}{d}$. With $d=\deg T$, we can write $T({\boldsymbol{x}})$ as
 $T({\boldsymbol{x}})=Q([{\boldsymbol{x}}]_{d}\otimes I_{p})\text{ with }Q=[Q_{1},\ldots,Q_{|\mathbb{N}^{n}_{d}|}],\quad Q_{i}\in{\mathbb{R}}^{l\times p},$
-where $Q$ is the vector of coefficient matrices of $T({\boldsymbol{x}})$ with respect to $[{\boldsymbol{x}}]_{d}$.
-Hence, $\Sigma({\boldsymbol{x}})$ is an SOS matrix with respect to $[{\boldsymbol{x}}]_{d}$ if there exists some $Q\in{\mathbb{R}}^{l\times p|\mathbb{N}^{n}_{d}|}$ satisfying
-$\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}(Q^{\intercal}Q)([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
-We thus have the following results.
+where $Q$ is the vector of the coefficient matrices of $T({\boldsymbol{x}})$ with respect to ${\boldsymbol{x}}]_{d}$.
+Hence, $S({\boldsymbol{x}})$ is an SOS matrix with respect to ${\boldsymbol{x}}]_{d}$ if there exists some $Q\in{\mathbb{R}}^{l\times p|\mathbb{N}^{n}_{d}|}$ satisfying
+$S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}(Q^{\intercal}Q)([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
+Thus, we have the following results.
 Proposition 2.1.
-[45, Lemma 1] A polynomial matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is an SOS matrix with respect to the monomial basis $[{\boldsymbol{x}}]_{d}$ if and only if there exists $Z\in\mathbb{S}_{+}^{p|\mathbb{N}^{n}_{d}|}$ such that
-$\Sigma({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}Z([{\boldsymbol{x}}]_{d}\otimes I_{p}).$
+[42, Lemma 1] A polynomial matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is an SOS matrix with respect to the monomial basis ${\boldsymbol{x}}]_{d}$ if and only if there exists $Z\in\mathbb{S}_{+}^{p|\mathbb{N}^{n}_{d}|}$ such that $S({\boldsymbol{x}})=([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}Z([{\boldsymbol{x}}]_{d}\otimes I_{p})$.
 Lemma 2.1.
-Let $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be an SOS matrix and $\Sigma_{k}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{k\times k}$ be a principal submatrix of $\Sigma({\boldsymbol{x}})$ whose rows and columns are indexed by $(p_{1},\ldots,p_{k})$ with $1\leq p_{1}<\cdots<p_{k}\leq p$, then $\Sigma_{k}({\boldsymbol{x}})$ is an SOS matrix.
+Let $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be an SOS matrix and $S_{k}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{k\times k}$ be a principal submatrix of $S({\boldsymbol{x}})$ whose rows and columns are indexed by $(p_{1},\ldots,p_{k})$ with $1\leq p_{1}<\cdots<p_{k}\leq p$, then $S_{k}({\boldsymbol{x}})$ is an SOS matrix.
 Proof.
-As $\Sigma({\boldsymbol{x}})$ is an SOS matrix, there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$.
-Denote by $T_{k}({\boldsymbol{x}})$ the submatrix of $T({\boldsymbol{x}})$ consisting of the colums of $T(x)$indexed by $(p_{1},\ldots,p_{k})$.
-Then, $\Sigma_{k}({\boldsymbol{x}})=T_{k}({\boldsymbol{x}})^{\intercal}T_{k}({\boldsymbol{x}})$ and hence is an SOS matrix.
+As $S({\boldsymbol{x}})$ is an SOS matrix, there exists an $l\times p$ polynomial matrix $T({\boldsymbol{x}})$ for some $l\in\mathbb{N}$ such that $S({\boldsymbol{x}})=T({\boldsymbol{x}})^{\intercal}T({\boldsymbol{x}})$.
+Denote by $T_{k}({\boldsymbol{x}})$ the submatrix of $T({\boldsymbol{x}})$ consisting of the columns of T(x)T(x)indexed by $(p_{1},\ldots,p_{k})$.
+Then $S_{k}({\boldsymbol{x}})=T_{k}({\boldsymbol{x}})^{\intercal}T_{k}({\boldsymbol{x}})$ and hence is an SOS matrix.
 ∎
@@
 $\mbox{tr}_{p}\left(C\right)\coloneqq\left[\begin{array}[]{ccc}\mbox{tr}\left(C_{11}\right)&\cdots&\mbox{tr}\left(C_{1p}\right)\\ \vdots&\ddots&\vdots\\ \mbox{tr}\left(C_{p1}\right)&\cdots&\mbox{tr}\left(C_{pp}\right)\end{array}\right]\quad\text{for }C=[C_{ij}]_{i,j\in[p]}\in{\mathbb{R}}^{pq\times pq},C_{ij}\in{\mathbb{R}}^{q\times q}.$
-When $p=1$, the product $\left\langle A,B\right\rangle_{1}$ coicides with the matrix inner product $\langle A,B\rangle=\mbox{tr}\left(A^{\intercal}B\right)$.
+When $p=1$, the product $\left\langle A,B\right\rangle_{1}$ coincides with the matrix inner product $\langle A,B\rangle=\mbox{tr}\left(A^{\intercal}B\right)$.
 Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$.
 Lemma 2.2.
-Let $H({\boldsymbol{x}})=\mbox{diag}(H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))$ be a block diagonal matrix, then for any SOS matrix $\Sigma({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr}$ where $p\in\mathbb{N}$ and $r=r_{1}+\cdots+r_{t}$, there are SOS matrices $\Sigma_{j}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$, $j\in[t]$, such that
-$\left\langle\Sigma({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
+Let $H({\boldsymbol{x}})=\mbox{diag}(H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))$ be a block diagonal matrix, then for any SOS matrix $S({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr}$ where $p\in\mathbb{N}$ and $r=r_{1}+\cdots+r_{t}$, there are SOS matrices $S_{j}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$, $j\in[t]$, such that
+$\left\langle S({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
 Proof.
-For each $j\in[t]$, let $\Sigma_{j}({\boldsymbol{x}})$ be the $pr_{j}\times pr_{j}$ principal submatrix of $\Sigma({\boldsymbol{x}})$ whose rows and columns are indexed by
+For each $j\in[t]$, let $S_{j}({\boldsymbol{x}})$ be the $pr_{j}\times pr_{j}$ principal submatrix of $S({\boldsymbol{x}})$ whose rows and columns are indexed by
 $\left(\sum_{\ell=1}^{j-1}r_{\ell}+1,\sum_{\ell=1}^{j}r_{\ell},r+\sum_{\ell=1}^{j-1}r_{\ell}+1,r+\sum_{\ell=1}^{j}r_{\ell},\ldots,(p-1)r+\sum_{\ell=1}^{j-1}r_{\ell}+1,(p-1)r+\sum_{\ell=1}^{j}r_{\ell}\right).$
-As $H({\boldsymbol{x}})$ is block diagonal, by the definition of the mapping .,.$\left\langle.,.\right\rangle_{p}$, it is easy to see that
-$\left\langle\Sigma({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
-By Lemma 2.1, each $\Sigma_{j}({\boldsymbol{x}})$ is an SOS matrix. ∎
-The matrix quadratic module $\mathcal{Q}^{p}(\mathbf{H})$ generated by $\mathbf{H}$ is defined as
-$\mathcal{Q}^{p}(\mathbf{H})\coloneqq\left\{\Sigma_{0}({\boldsymbol{x}})+\sum_{j=1}^{t}\left\langle\Sigma_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr_{j}},j\in[t],\ \text{are SOS}\right\}.$
+As $H({\boldsymbol{x}})$ is block diagonal, by the definition of the mapping $\left\langle.,.\right\rangle_{p}$, it is easy to see that
+$\left\langle S({\boldsymbol{x}}),H({\boldsymbol{x}})\right\rangle_{p}=\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}.$
+By Lemma 2.1, each $S_{j}({\boldsymbol{x}})$ is an SOS matrix. ∎
+The matrix quadratic module $\mathcal{Q}^{p}(\mathbf{H})$ generated by $\mathbf{H}$ is defined as the cone
+$\mathcal{Q}^{p}(\mathbf{H})\coloneqq\left\{S_{0}({\boldsymbol{x}})+\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{pr_{j}},j\in[t],\ \text{are SOS}\right\}.$
+The $k$-th truncated matrix quadratic module $\mathcal{Q}^{p}_{k}(\mathbf{H})$ is a subcone of $\mathcal{Q}^{p}(\mathbf{H})$ in which the degrees of the polynomial matrices $S_{0}({\boldsymbol{x}})$ and $\sum_{j=1}^{t}\left\langle S_{j}({\boldsymbol{x}}),H_{j}({\boldsymbol{x}})\right\rangle_{p}$ are bounded by $2k$.
 Assumption 2.1.
@@
 (Scherer-Hol’s Positivestellensatz) Let Assumption 2.1 and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ be PD on $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$.
-Then, $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{H})$.
+Then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(\mathbf{H})$.
 Proof.
-It was proved in [45, Corollary 1] for the case $t=1$. The case $t>1$ can be derived from [45, Corollary 1] and Lemma 2.2. ∎
-2.2. Matrix-valued measures and moment problem
-Now we recall some background on the concept of matrix-valued measures.
-Denote by $\mathfrak{B}(\mathcal{X})$ the smallest $\sigma$-algebra generated from the open subsets of $\mathcal{X}$ and by $\mathfrak{m}(\mathcal{X})$ the set of all finite Borel measures on $\mathcal{X}$.
-The support $\mathrm{supp}(\phi)$ of a Borel measure $\phi\in\mathfrak{m}(\mathcal{X})$ is the (unique) smallest closed set $\mathbf{A}\in\mathfrak{B}(\mathcal{X})$ such that $\phi(\mathcal{X}\setminus\mathbf{A})=0$.
-Let $\phi_{ij}\in\mathfrak{m}(\mathcal{X})$, $i,j=1,\ldots,p$.
-The $p\times p$ matrix-valued measure $\Phi$ on $\mathcal{X}$ is defined as the matrix-valued function $\Phi\colon\mathfrak{B}(\mathcal{X})\to{\mathbb{R}}^{p\times p}$ with
-$\Phi(\mathbf{A})\coloneqq[\phi_{ij}(\mathbf{A})]\in{\mathbb{R}}^{p\times p},\quad\forall\mathbf{A}\in\mathfrak{B}(\mathcal{X}).$
-If $\phi_{ij}=\phi_{ji}$ for all $i,j=1,\ldots,p$, we call $\Phi$ a symmetric matrix-valued measure.
-If ${\boldsymbol{v}}^{\intercal}\Phi(\mathbf{A}){\boldsymbol{v}}\geq 0$ holds for all $\mathbf{A}\in\mathfrak{B}(\mathcal{X})$ and for all column vectors ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$, we call $\Phi$ a PSD matrix-valued measure.
-The set $\mathrm{supp}(\Phi)\coloneqq\bigcup_{i,j=1}^{p}\mathrm{supp}(\phi_{ij})$ is called the support of the matrix-valued measure $\Phi$.
-We denote by $\mathfrak{M}^{p}_{+}(\mathcal{X})$ the set of all $p\times p$ PSD symmetric matrix-valued measures on $\mathcal{X}$.
-For a polynomial matrix $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and a matrix-valued measure $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$, the integral of $H$ with respect to $\Phi$ is defined by
-$\int_{\mathcal{X}}H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\coloneqq\int_{\mathcal{X}}\ \mbox{tr}\left(H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\right)=\sum_{i,j}\int_{\mathcal{X}}\ H_{ij}({\boldsymbol{x}})\mathrm{d}\phi_{ij}({\boldsymbol{x}}).$
-Let $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ be a multi-indexed sequence of symmetric matrices in ${\mathbb{S}}^{p}$.
-We define a linear functional $\mathscr{L}_{\mathbf{S}}\colon{\mathbb{S}}[{\boldsymbol{x}}]^{p}\to{\mathbb{R}}$ in the following way:
-$\mathscr{L}_{\mathbf{S}}(H)\coloneqq\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}\mbox{tr}\left(H_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\alpha}}}\right),\quad\forall H({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}H_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}.$
-We call $\mathscr{L}_{\mathbf{S}}$ the Riesz functional associated to the sequence $\mathbf{S}$.
-We say the sequence $\mathbf{S}$ has a matrix-valued representing measure $\Phi=[\phi_{ij}]\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ if
-$S_{{\boldsymbol{\alpha}}}=\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})\coloneqq\left[\int_{\mathbf{K}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi_{ij}({\boldsymbol{x}})\right]_{i,j\in[p]},\ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}.$ (3)
-The following theorem is a matrix version of Haviland’s theorem ([8, 46]).
-Theorem 2.2 (Haviland’s theorem for polynomial matrices).
-[8, Theorem 3] A given sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ has a matrix-valued representing measure $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ if and only if $\mathscr{L}_{\mathbf{S}}(H)\geq 0$ for all $H({\boldsymbol{x}})$ which are PSD on $\mathcal{X}$.
-[…]
-By a $\star$-algebra we mean an algebra $\mathcal{A}$ over the field ${\mathbb{K}}={\mathbb{R}}$ or ${\mathbb{K}}={\mathbb{C}}$ equipped with an involution $a\to a^{\star}$ of $\mathcal{A}$ into itself such that $(\lambda a+\mu b)^{\star}=\bar{\lambda}a^{\star}+\bar{\mu}b^{\star}$, $(ab)^{\star}=b^{\star}a^{\star}$ and $(a^{\star})^{\star}$ for all $a,b\in\mathcal{A}$ and $\lambda,\mu\in{\mathbb{K}}$.
-Denote by $1$ the unit element of $\mathcal{A}$.
-If $a$ is an operator on a Hilbert space, denote by $\mathcal{D}(a)$ its domain.
-[…]
-A state of $\mathcal{A}$ is a $\star$-linear functional $\Pi$ on $\mathcal{A}$ such that $\Pi(1)=1$ and $\Pi(a^{\star}a)\geq 0$ for all $a\in\mathcal{A}$.
-If $\pi$ is a $\star$-representation of $\mathcal{A}$ and $\varphi\in\mathcal{D}(\pi)$ is a unit vector, then $\Pi(\cdot)\coloneqq\langle\pi(\cdot)\varphi,\varphi\rangle$ is a state of $\mathcal{A}$.
-Conversely, for each state $\Pi$ on $\mathcal{A}$ there exist a $\star$-representation of $\mathcal{A}$, called the GNS representation $\pi_{\Pi}$ of $\Pi$, and a vector $\varphi_{\Pi}\in\mathcal{D}(\pi_{\Pi})$ such that $\mathcal{D}(\pi_{\Pi})=\pi_{\Pi}(\mathcal{A})\varphi_{\Pi}$ and $\Pi(a)=\langle\pi_{\Pi}(a)\varphi_{\Pi},\varphi_{\Pi}\rangle$ for all $a\in\mathcal{A}$.
+It was proved in [42, Corollary 1] for the case $t=1$. The case $t>1$ can be derived from [42, Corollary 1] and Lemma 2.2. ∎
 3. Positivstellensätze for polynomial matrices with UQs under the Archimedean condition
-In this section, we shall present a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs, assuming the Archimedean condition, along with a sparse version in the presence of correlative sparsity.
-3.1. Scherer-Hol-type Positivestellensatz
+In this section, assuming the Archimedean condition, we shall present a matrix-valued Positivstellensatz incorporating universal quantifiers, along with a sparse version in the presence of correlative sparsity.
+3.1. A Positivstellensatze for polynomial matrices with UQs
 Recall the set $\mathcal{X}$ in (1).
@@
 The matrix quadratic module $\mathcal{Q}^{p}(G,\nu)$ generated by $G$ and $\nu$ is defined as
-$\mathcal{Q}^{p}(G,\nu)\coloneqq\left\{\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices}\right\}.$
-Let us first derive the Scherer-Hol-type Positivestellensatz to provide SOS-structured characterizations for polynomial matrices that are PD over $\mathcal{X}$.
+$\mathcal{Q}^{p}(G,\nu)\coloneqq\left\{S_{0}+\int_{\mathcal{Y}}\left\langle S,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices}\right\}.$
+Next, we derive a Positivstellensatze for polynomial matrices with UQs, providing SOS-structured characterizations for polynomial matrices that are PD over $\mathcal{X}$.
 Assumption 3.1.
 $\mathcal{Q}^{p}(G,\nu)$ is Archimedean, i.e., there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)$.
-Lemma 3.1.
-Under Assumption 3.1, the identity matrix $I_{p}$ belongs to the algebraic interior of the cone $\mathcal{Q}^{p}(G,\nu)$ in ${\mathbb{S}}[{\boldsymbol{x}}]^{p}$, i.e., for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, there exists a real number $t_{0}\geq 0$ such that $H({\boldsymbol{x}})+tI_{p}\in\mathcal{Q}^{p}(G,\nu)$ for all $t\geq t_{0}$.
-Proof.
-Without loss of generality, we assume that $H({\boldsymbol{x}})=\pm{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ for some ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ and ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-We first show by induction that for any ${\boldsymbol{\beta}}\in\mathbb{N}^{n}$, $C^{|{\boldsymbol{\beta}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}\in\mathcal{Q}^{1}(G,\nu)$.
-Note that $C-x_{j}^{2}\in\mathcal{Q}^{1}(G,\nu)$ since
-$C-x_{j}^{2}=(C-\|{\boldsymbol{x}}\|^{2})+\sum_{i\neq j}x_{j}^{2}\in\mathcal{Q}^{1}(G,\nu).$
-Assume that the claim holds for ${\boldsymbol{\beta}}$ and consider the monomial ${\boldsymbol{x}}^{{\boldsymbol{\beta}}}x_{j_{0}}$ for some $j_{0}\in[n]$. We have
-$C^{|{\boldsymbol{\beta}}|+1}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}x_{j_{0}}^{2}=C(C^{|{\boldsymbol{\beta}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\beta}}})+{\boldsymbol{x}}^{2{\boldsymbol{\beta}}}(C-x_{j_{0}}^{2})\in\mathcal{Q}^{1}(G,\nu),$
-by induction.
-Now note that
-±$\pm 2{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}={\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(({\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\pm 1)^{2}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}}-1)$
-and
-${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}}-1)={\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(C^{|{\boldsymbol{\alpha}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}})+(C^{|{\boldsymbol{\alpha}}|}+1)({\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal})-(C^{|{\boldsymbol{\alpha}}|}+1){\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}.$
-By Cauchy’s inequality, ${\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}\succeq 0$. Hence, by the above claim,
-${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}(C^{|{\boldsymbol{\alpha}}|}-{\boldsymbol{x}}^{2{\boldsymbol{\alpha}}})+(C^{|{\boldsymbol{\alpha}}|}+1)({\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}I_{p}-{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal})\in\mathcal{Q}^{p}(G,\nu).$
-Thus, by letting $t_{0}=\frac{(C^{|{\boldsymbol{\alpha}}|}+1){\boldsymbol{v}}^{\intercal}{\boldsymbol{v}}}{2}$, we have $H({\boldsymbol{x}})+tI_{p}\in\mathcal{Q}^{p}(G,\nu)$ for all $t\geq t_{0}$. ∎
-Corollary 3.1.
-If $F({\boldsymbol{x}})\not\in\mathcal{Q}^{p}(G,\nu)$, then there exists a linear functional $\mathscr{L}:{\mathbb{S}}[{\boldsymbol{x}}]^{p}\to{\mathbb{R}}$ such that $\mathscr{L}(\mathcal{Q}^{p}(G,\nu))\subseteq{\mathbb{R}}_{+}$, $\mathscr{L}(F)\leq 0$, and $\mathscr{L}(I_{p})=1$.
-Proof.
-As the cone $\mathcal{Q}^{p}(G,\nu)$ has non-empty algebraic interior by Lemma 3.1, this is due to the Hahn-Banach separation theorem [1, Corollary III.1.7]. ∎
 Consider the Carleman condition imposed on $\nu$ which is automatically satisfied when $\mathcal{Y}$ is compact.
 Assumption 3.2.
-The Borel measure $\nu$ satisfies the multivariable Carleman condition
+The Borel measure $\nu$ satisfies the multivariate Carleman condition
 $\sum_{d=1}^{\infty}\left(\int_{\mathcal{Y}}y_{j}^{2d}\mathrm{d}\nu({\boldsymbol{y}})\right)^{-\frac{1}{2d}}=\infty,\quad\forall j\in[m].$
 Proposition 3.1.
-[18, Proposition 3.2] Suppose that $\nu$ satisfies Assumption 3.2. Then, SOS polynomials are dense in the cone of nonnegative functions in $L^{2}({\mathbb{R}}^{m},\nu)$.
-Now, we present our main fundamental result concerning a Scherer-Hol-type Positivstellensatz for polynomial matrices with UQs.
+[16, Proposition 3.2] Suppose that $\nu$ satisfies Assumption 3.2. Then, SOS polynomials are dense in the cone of nonnegative functions in $L^{2}({\mathbb{R}}^{m},\nu)$.
+Proposition 3.2.
+Suppose that $\nu$ satisfies Assumption 3.2. Then
+$\mathcal{X}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0,\ \forall H\in\mathcal{Q}^{p}(G,\nu)\}.$ (3)
+Proof.
+By the definition of $\mathcal{Q}^{p}(G,\nu)$, we only need to prove that $\mathcal{X}$ contains the set on the right-hand side of the equation in (3).
+Fix ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$ with $H({\boldsymbol{u}})\succeq 0$ for all $H\in\mathcal{Q}^{p}(G,\nu)$.
+Suppose on the contrary that ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$.
+Then, there is a ball $\mathcal{O}\subset{\mathbb{R}}^{m}$ with a radius $\rho>0$ around $\boldsymbol{w}$} such that $G({\boldsymbol{u}},{\boldsymbol{y}})\not\succeq 0$ on $2\mathcal{O}$.
+We may assume that there exists ${\boldsymbol{v}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\leq-\delta$ on $2\mathcal{O}$.
+Define a continuous function $h({\boldsymbol{y}})$ on ${\mathbb{R}}^{m}$ by $h({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$ for ${\boldsymbol{y}}\in 2\mathcal{O}$ and $h({\boldsymbol{y}})=0$ otherwise.
+By Proposition 3.1, there exists a sequence of SOS polynomials $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h$ in the $L^{2}$-norm.
+Hence,
+$\lim_{k\to\infty}\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}h({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$
+$=$ $\int_{\mathcal{Y}\cap 2\mathcal{O}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})\leq\int_{\mathcal{Y}\cap\mathcal{O}}-\delta\rho\mathrm{d}\nu({\boldsymbol{y}})=-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O})<0,$
+where the last inequality is due to the fact that $\mathrm{supp}(\nu)=\mathcal{Y}$ and thus $\nu(\mathcal{Y}\cap\mathcal{O})>0$. Note that for all $k\in\mathbb{N}$,
+$\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{v}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\frac{1}{p}\mbox{tr}\left(\int_{\mathcal{Y}}\left\langle\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{u}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
+since $\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}$ is an SOS and $H({\boldsymbol{u}})\succeq 0$ for all $H\in\mathcal{Q}^{p}(G,\nu)$.
+A contradiction follows.
+∎
+Now, we present our main result concerning a Positivstellensatz for polynomial matrices with UQs.
 Theorem 3.1.
@@
 Proof.
-The proof is by contradiction, using the functional $\mathscr{L}$ from Corollary 3.1, and proceeds as follows:
-  1. 1.
-    Construct a Hilbert space $\mathcal{H}$ and operators $\widehat{\mathbf{A}}\coloneqq(\widehat{A}_{i})_{i\in[n]}$ with $\widehat{A}_{i}\in\mathcal{B}(\mathcal{H})$, where $\mathcal{B}(\mathcal{H})$ denotes the set of bounded linear operators on $\mathcal{H}$.
-Then construct a state $\Pi$ on the the $p\times p$ matrix algebra $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ over $\mathcal{B}(\mathcal{H})$ such that $\Pi(F(\widehat{\mathbf{A}}))=\mathscr{L}(F)$;
-  2. 2.
-    By applying the GNS construction to $(\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})),\Pi)$, construct a Hilbert space $\mathcal{K}_{1}$ with inner product $\langle\cdot,\cdot\rangle_{\mathcal{K}_{1}}$ and operators $\mathbf{A}=(A_{i})_{i\in[n]}$ with $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ such that $\mathscr{L}(F)=\sum_{i,j=1}^{p}\left\langle F_{ij}(\mathbf{A})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}_{1}}$ for some $\eta_{i}\in\mathcal{K}_{1}$, $i\in[p]$;
-  3. 3.
-    By applying the spectral theorem to $\mathbf{A}=(A_{i})_{i\in[n]}$, construct a $p\times p$ PSD matrix-valued measure $\Phi$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ such that $\mathscr{L}(F)=\int_{{\mathbb{R}}^{n}}F({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})$;
-  4. 4.
-    Show that the support of $\Phi$ is contained in $\mathcal{X}$ and derive a contradiction by Corollary 3.1.
-For the sake of readability, the details of the proof is postponed to Section 6.1. ∎
-To conclude this part, let us derive a simple corollary of Theorem 3.1, which will be used in Section 4. We need the following lemma.
+Since the quadratic module $\mathcal{Q}^{p}(G,\nu)$ is Archimedean and the equality in (3) holds, the conclusion follows from the fundamental Positivstellensatz for matrix algebras of polynomials [44, Theorem 10.25].
+∎
+Next, we derive a corollary of Theorem 3.1, which will be used in Section 4.
 Lemma 3.1.
-Suppose that $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ is an SOS matrix in ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$, then $\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in ${\boldsymbol{x}}$.
+Suppose that $S({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ is an SOS matrix in $\boldsymbol{x}$} and $\boldsymbol{y}$}.
+Then $\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in $\boldsymbol{x}$}.
 Proof.
-As $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})$ is an SOS matrix, there exists an ×$\ell\times q$ polynomial matrix $T({\boldsymbol{x}},{\boldsymbol{y}})$ for some $\ell\in\mathbb{N}$ such that $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}T({\boldsymbol{x}},{\boldsymbol{y}})$.
-With $d=\deg_{{\boldsymbol{x}}}(T)$, we could write $T({\boldsymbol{x}},{\boldsymbol{y}})$ as
+As $S({\boldsymbol{x}},{\boldsymbol{y}})$ is an SOS matrix, there exists an $\ell\times q$ polynomial matrix $T({\boldsymbol{x}},{\boldsymbol{y}})$ for some $\ell\in\mathbb{N}$ such that $S({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}T({\boldsymbol{x}},{\boldsymbol{y}})$.
+With $d=\deg_{{\boldsymbol{x}}}T$, we could write $T({\boldsymbol{x}},{\boldsymbol{y}})$ as
 $T({\boldsymbol{x}},{\boldsymbol{y}})=Q({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})\text{ with }Q({\boldsymbol{y}})=[Q_{1}({\boldsymbol{y}}),\ldots,Q_{|\mathbb{N}^{n}_{d}|}({\boldsymbol{y}})],\quad Q_{i}({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{y}}]^{\ell\times q},$
-where $Q({\boldsymbol{y}})$ is the vector of coefficient matrices of $T({\boldsymbol{x}},{\boldsymbol{y}})$ (considered as a polynomial matrix in ${\mathbb{R}}[{\boldsymbol{x}}]^{\ell\times q}$) with respect to $[{\boldsymbol{x}}]_{d}$.
+where $Q({\boldsymbol{y}})$ is the vector of coefficient matrices of $T({\boldsymbol{x}},{\boldsymbol{y}})$ (considered as a polynomial matrix in ${\mathbb{R}}[{\boldsymbol{x}}]^{\ell\times q}$) with respect to ${\boldsymbol{x}}]_{d}$.
 Hence,
-$\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ $=\int_{\mathcal{Y}}([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}(Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}}))([{\boldsymbol{x}}]_{d}\otimes I_{q})\mathrm{d}\nu({\boldsymbol{y}})$
-$=([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})$
-As $\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is PSD, $\int_{\mathcal{Y}}\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in ${\boldsymbol{x}}$.
+$\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}}([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}(Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}}))([{\boldsymbol{x}}]_{d}\otimes I_{q})\mathrm{d}\nu({\boldsymbol{y}})=([{\boldsymbol{x}}]_{d}\otimes I_{q})^{\intercal}\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})([{\boldsymbol{x}}]_{d}\otimes I_{q})$
+As $\int_{\mathcal{Y}}Q({\boldsymbol{y}})^{\intercal}Q({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is PSD, $\int_{\mathcal{Y}}S({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})$ is an SOS matrix in $\boldsymbol{x}$}.
 ∎
@@
 Let $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$ where each $H_{j}\in{\mathbb{S}}[{\boldsymbol{x}}]^{r_{j}}$ for some $r_{j}\in\mathbb{N}$.
-Suppose that Assumption 3.2 holds and there is $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)+\mathcal{Q}^{1}(\mathbf{H})$.
+Suppose that Assumption 3.2 holds and there exists $C>0$ such that $C-\|{\boldsymbol{x}}\|^{2}\in\mathcal{Q}^{1}(G,\nu)+\mathcal{Q}^{1}(\mathbf{H})$.
 If $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PD on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$, then $F({\boldsymbol{x}})\in\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$.
 Proof.
-It is clear that $F({\boldsymbol{x}})\succ 0$ on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succ 0,\ j\in[t]\}$ if and only if $F({\boldsymbol{x}})\succ 0$ on
+It is clear that $F({\boldsymbol{x}})\succ 0$ on $\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H_{j}({\boldsymbol{x}})\succeq 0,\ j\in[t]\}$ if and only if $F({\boldsymbol{x}})\succ 0$ on
 $\left\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\,\middle|\,\widehat{G}({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\mbox{diag}(G({\boldsymbol{x}},{\boldsymbol{y}}),H_{1}({\boldsymbol{x}}),\ldots,H_{t}({\boldsymbol{x}}))\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\right\}.$
-By Lemmas 2.2 and 3.2, it is easy to see that $\mathcal{Q}^{p}(\widehat{G},\nu)=\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$. So the conclusion follows from Theorem 3.1. ∎
+By Lemmas 2.2 and 3.1, it is easy to see that $\mathcal{Q}^{p}(\widehat{G},\nu)=\mathcal{Q}^{p}(G,\nu)+\mathcal{Q}^{p}(\mathbf{H})$. So, the conclusion follows from Theorem 3.1. ∎
 3.2. A sparse Positivstellensatz for polynomial matrices with UQs
@@
 Let $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ and
-$\widehat{\mathcal{X}}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall j\in[s],\ {\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$
+$\widehat{\mathcal{X}}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall j\in[s],\ {\boldsymbol{y}}\in\mathcal{Y}\subset{\mathbb{R}}^{m}\},$ (4)
 where each $G_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$, $q_{j}\in\mathbb{N}$.
@@
     The running intersection property holds for $\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$, i.e.,
-    for =$\text{for }\ell=2,\ldots,t,\ \exists k<\ell\quad\text{s.t.}\quad\mathcal{I}_{\ell}\cap\bigcup_{j<\ell}\mathcal{I}_{j}\subseteq\mathcal{I}_{k};$
+    $\text{for }\ell=2,\ldots,t,\ \exists k<\ell\quad\text{s.t.}\quad\mathcal{I}_{\ell}\cap\bigcup_{j<\ell}\mathcal{I}_{j}\subseteq\mathcal{I}_{k};$
   2. (ii)
@@
 For each $\ell\in[t]$, let $\mathbf{G}^{\ell}\coloneqq\{G_{j}\}_{j\in\mathcal{J}_{\ell}}$ and $\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)$ be the quadratic module generated by $\mathbf{G}^{\ell}$ and $\nu$ in ${\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p}$, i.e.,
-$\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)\coloneqq\left\{\Sigma_{0}+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle\Sigma_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p},\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{pq_{j}},j\in\mathcal{J}_{\ell},\text{ are SOS matrices}\right\}.$
+$\mathcal{Q}^{p}(\mathbf{G}^{\ell},\nu)\coloneqq\left\{S_{0}+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle S_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\ \middle|\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]^{p},S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{pq_{j}},j\in\mathcal{J}_{\ell},\text{ are SOS matrices}\right\}.$
 Using the correlative sparsity pattern and the Archimedean condition, we are able to derive the following sparse Positivstellensatz.
@@
 Proposition 3.3.
-[12, Lemma 3] Let {}$\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ satisfy the running intersection property.
-For any $C>0$, if $f=f_{1}+\cdots+f_{t}$ with $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ satisfies $f>0$ on $[-C,C]^{n}$, then $f=h_{1}+\ldots+h_{t}$ for some $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ with $h_{\ell}>0$ on $[-C,C]^{|\mathcal{I}_{\ell}|}$.
+[11, Lemma 3] Let $\{\mathcal{I}_{\ell}\}_{\ell\in[t]}$ satisfy the running intersection property.
+For any $C>0$, if $f=f_{1}+\cdots+f_{t}$ with $f_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ satisfies $f>0$ on $C,C]^{n}$, then $f=h_{1}+\ldots+h_{t}$ for some $h_{\ell}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ with $h_{\ell}>0$ on $C,C]^{|\mathcal{I}_{\ell}|}$.
 Proposition 3.4.
@@
 $F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succ 0$ (5)
-on $[-C,C]^{n}$ for all $k\geq\bar{k}$.
-The proof of Proposition 3.3 is postponed to Section 6.2.
-Note that using Proposition 3.3, we can provide a simpler proof of the Scherer-Hol-type Positivestellensatz (Theorem 3.1) assuming that $\mathcal{Y}$ is compact (see Section 6.2).
+on $C,C]^{n}$ for all $k\geq\bar{k}$.
+The proof of Proposition 3.4 is postponed to Section 6.1. Note that using Proposition 3.4, we can provide a constructive proof of Theorem 3.1 assuming that $\mathcal{Y}$ is compact (see Section 6.1).
 Proof of Theorem 3.2.
@@
 $f({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})>0$
-on $[-\sqrt{C},\sqrt{C}]^{n}$. Note that for each $j\in[s]$,
+on $\sqrt{C},\sqrt{C}]^{n}$. Note that for each $j\in[s]$,
 $\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$
@@
 $f({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k^{\prime}},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})=h_{1}+\cdots+h_{t},$
-and each $h_{\ell}>0$ on $[-\sqrt{C},\sqrt{C}]^{|\mathcal{I}_{\ell}|}$.
+and each $h_{\ell}>0$ on $\sqrt{C},\sqrt{C}]^{|\mathcal{I}_{\ell}|}$.
 By Putinar’s Positivstellensatz [39], for each $h_{\ell}$, there exist SOS polynomials $\sigma_{\ell,0},\sigma_{\ell,1}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})]$ such that
@@
 Remark 3.1.
-One might wonder whether the result in Theorem 3.2 holds for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ with $p>1$.
+One might wonder whether the result of Theorem 3.2 holds for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ with $p>1$.
 Indeed, this is not true even in the absence of UQs; see [32] for a counterexample.
@@
 The results in Section 3 are derived under the Archimedean condition on the quadratic module generated by $(G,\nu)$, which requires $\mathcal{X}$ to be bounded.
-Lots of efforts have been made to provide SOS-structured representations for polynomials that are PD on a unbounded semialgebraic set.
+Many efforts have been made to provide SOS-structured representations for polynomials that are PD on a semialgebraic set without the Archimedean condition.
 The goal of this section is to extend some of these well-known results to the setting of polynomial matrices with UQs.
-4.1. Putinar–Vasilescu-type Positivstellensatz
+4.1. A Positivstellensatz for the non-compact case
 Let $\theta\coloneqq 1+\|{\boldsymbol{x}}\|^{2}$.
-For every polynomial $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ nonnegative on a general basic semialgebraic set, Putinar and Vasilescu [42] proved that for a given $\varepsilon>0$ and $2d\geq\deg(f)$, there exists a nonnegative integer $k$ such that $\theta^{k}(f+\varepsilon\theta^{d})$ belongs to the quadratic module associated with the basic semialgebraic set.
+For every nonnegative polynomial $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ on a general basic semialgebraic set, Putinar and Vasilescu [38] proved that for a given $\varepsilon>0$ and $2d\geq\deg f$, there exists a nonnegative integer $k$ such that $\theta^{k}(f+\varepsilon\theta^{d})$ belongs to the quadratic module associated with the basic semialgebraic set.
 Using Jacobi’s technique [18], Mai et al.
 [30] provided an alternative proof of Putinar and Vasilescu’s result with an effective degree bound on polynomials involved in such certificates.
-Building on Theorem 3.1 and similiar techniques from [34], we next derive a Putinar–Vasilescu-type Positivstellensatz for polynomial matrices with UQs.
+Building on Theorem 3.1 and similar techniques from [30], we next derive matrix-valued Positivstellensätz for the non-compact case, incorporating universal quantifiers.
 In the homogeneous case, the result is stated as follows.
+See Section 6.2 for its proof.
 Theorem 4.1.
 Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$.
-Then, there exists $N\in\mathbb{N}$ and SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and $\deg\Sigma_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}\Sigma=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
-$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-See Section 6.3 for its proof.
+Then, there exists $N\in\mathbb{N}$ and SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ that are homogeneous in $\boldsymbol{x}$} and $\deg S_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}S=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
+$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
 For any $H=[H_{ij}]\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{r}$, let $d_{H}\coloneqq\max\left\{\lfloor\deg_{{\boldsymbol{x}}}(H_{ij})/2\rfloor+1:i,j\in[r]\right\}$ and $\tilde{{\boldsymbol{x}}}\coloneqq({\boldsymbol{x}},x_{n+1})$.
@@
 Suppose that Assumption 3.2 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}$.
-Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$, SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg\Sigma_{0}\leq 2(N_{\varepsilon}+d_{F})$ and $\deg_{{\boldsymbol{x}}}\Sigma\leq 2(N_{\varepsilon}+d_{F}-d_{G})$, such that
-$\theta^{N_{\varepsilon}}(F({\boldsymbol{x}})+\varepsilon\theta^{d_{F}}I_{p})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$, SOS matrices $S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, and $S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg S_{0}\leq 2(N_{\varepsilon}+d_{F})$ and $\deg_{{\boldsymbol{x}}}S\leq 2(N_{\varepsilon}+d_{F}-d_{G})$, such that
+$\theta^{N_{\varepsilon}}(F({\boldsymbol{x}})+\varepsilon\theta^{d_{F}}I_{p})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
 Proof.
@@
   2. -
-    Case 2: $u_{n+1}=0$.
-By the definition of $d_{F}$, $x_{n+1}$ divides $\widetilde{F}(\tilde{{\boldsymbol{x}}})$.
-Thus, $\widetilde{F}(\tilde{{\boldsymbol{u}}})=0$.
-Since $\|\tilde{{\boldsymbol{u}}}\|^{2}\neq 0$, $\widetilde{F}(\tilde{{\boldsymbol{u}}})+\varepsilon\|\tilde{{\boldsymbol{u}}}\|^{2d_{F}}I_{p}\succ 0$.
-Applying Theorem 4.1 to the polynomial matrices $\tilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}$ and $\widetilde{G}$, we obtain $N_{\varepsilon}\in\mathbb{N}$ and SOS matrices $\widetilde{\Sigma}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\widetilde{\Sigma}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, which are homogeneous in ${\boldsymbol{x}}$ and $\deg\widetilde{\Sigma}_{0}=2N_{\varepsilon}+2d_{F}$, $\deg_{{\boldsymbol{x}}}\widetilde{\Sigma}=2N_{\varepsilon}+2d_{F}-2d_{G}$, such that
-$\|\tilde{{\boldsymbol{x}}}\|^{2N_{\varepsilon}}(\widetilde{F}(\tilde{{\boldsymbol{x}}})+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p})=\widetilde{\Sigma}_{0}(\tilde{{\boldsymbol{x}}})+\int_{\mathcal{Y}}\left\langle\widetilde{\Sigma}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}),\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$ (5)
+[…]
+Applying Theorem 4.1 to the polynomial matrices $\tilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}$ and $\widetilde{G}$, we obtain $N_{\varepsilon}\in\mathbb{N}$ and SOS matrices $\widetilde{S}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\widetilde{S}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, which are homogeneous in $\boldsymbol{x}$} and $\deg\widetilde{S}_{0}=2N_{\varepsilon}+2d_{F}$, $\deg_{{\boldsymbol{x}}}\widetilde{S}=2N_{\varepsilon}+2d_{F}-2d_{G}$, such that
+$\|\tilde{{\boldsymbol{x}}}\|^{2N_{\varepsilon}}(\widetilde{F}(\tilde{{\boldsymbol{x}}})+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p})=\widetilde{S}_{0}(\tilde{{\boldsymbol{x}}})+\int_{\mathcal{Y}}\left\langle\widetilde{S}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}),\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$ (6)
 Then, letting $x_{n+1}=1$ in (6), we achieve the desired conclusion. ∎
-4.2. Pólya-type Positivstellensatz
+4.2. A Positivstellensatz on the nonnegative orthant
 Pólya [35] proved that multiplying a positive homogeneous polynomial on the nonnegative orthant by some power of $\sum_{i=1}^{n}x_{i}$ yields a polynomial with nonnegative coefficients.
-A “positive” version of Pólya’s theorem was given by Powers and Reznick [39].
+Powers and Reznick [36] gave a ”positive” version of Pólya’s theorem.
 Dickinson and Povh [7] extended the result of Pólya [35] to provide a certificate for positive homogeneous polynomials on the intersection of the nonnegative orthant with a basic semialgebraic set.
-We next employ Theorem 4.1 and similar techniques from [9] to present a Pólya-type Positivstellensatz for homogeneous polynomial matrices being PD on the intersection of $\mathcal{X}$ with the nonnegative orthant.
+We next employ Theorem 4.1 and similar techniques from [7] to present a Positivstellensatz for homogeneous polynomial matrices being PD on the intersection of $\mathcal{X}$ with the nonnegative orthant.
 Denote by $\boldsymbol{e}$} the column vector in ${\mathbb{R}}^{n}$ of all ones.
@@
 Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact.
-Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$ and ${\boldsymbol{y}}\in\mathcal{Y}$, such that
-$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-The proof is postponed in Section 6.4.
-In particular, when $\mathcal{Y}$ is a semialgebriac set defined by PMIs, the coefficient matrices $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})$ in Theorem 4.2 have SOS-structured representations.
+Then, there exists $N\in\mathbb{N}$ and polynomial matrices
+$P_{0}=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
+which are homogeneous in $\boldsymbol{x}$} and satisfy $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G$ and ${\boldsymbol{y}}\in\mathcal{Y}$, such that
+$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+The proof is postponed to Section 6.3.
+In particular, when $\mathcal{Y}$ is a semialgebriac set defined by PMIs, the coefficient matrices $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})$ in Theorem 4.2 have SOS-structured representations.
 Corollary 4.2.
-Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$.
+Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$.
 Moreover, suppose that
-$\mathcal{Y}=\{{\boldsymbol{y}}\in{\mathbb{R}}^{m}\mid H_{1}({\boldsymbol{y}})\succeq 0,\ldots,H_{t}({\boldsymbol{y}})\succeq 0\},\ H_{i}({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{y}}]^{\ell_{i}},\ \ell_{i}\in\mathbb{N},\ i\in[t],$
-and the Archimedean condition holds for the quadratic module $\mathcal{Q}(\mathbf{H})$ where $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$.
-Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}(\mathbf{H})$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$,
-$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+$\mathcal{Y}=\{{\boldsymbol{y}}\in{\mathbb{R}}^{m}\mid H_{1}({\boldsymbol{y}})\succeq 0,\ldots,H_{t}({\boldsymbol{y}})\succeq 0\},\ H_{i}({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{y}}]^{\ell_{i}},\ \ell_{i}\in\mathbb{N},\ i\in[t],$ (7)
+and the Archimedean condition holds for the quadratic module $\mathcal{Q}^{pq}(\mathbf{H})$ where $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$.
+Then, there exists $N\in\mathbb{N}$ and polynomial matrices
+$P_{0}=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
+which are homogeneous in $\boldsymbol{x}$} and satisfy $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}^{pq}(\mathbf{H})$ for all $|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
+$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
 Proof.
 It follows from Theorem 4.2 and Scherer-Hol’s Positivestellensatz (Theorem 2.1). ∎
-4.3. Lasserre-Netzer perturbative Positivstellensatz
-Lasserre and Netzer [30] provided SOS approximations of polynomials that are nonnegative on $[-1,1]^{n}$ via simple high-degree perturbations.
-Motivated by their work, we next derive SOS-structured approximations for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ that is PSD on $\mathcal{X}\cap[-1,1]^{n}$ using high-degree perturbations.
-4.3.1. Exponentially bounded matrix-valued moment problem
-For a given $C>0$, Berg and Maserick [2] showed that a real sequence has representing measure supported on the hypercube $[-C,C]^{n}$ if and only if the sequence is exponentially bounded and the associated moment matrix is PSD.
-We next extend their results to characterize exponentially bounded matrix-valued sequence with a matrix-valued representing measure being supported on the set $\mathcal{X}\cap[-C,C]^{n}$.
-Beyond its intrinsic interest, we will use this result to establish a Lasserre-Netzer perturbative Positivstellensatz for polynomial matrices with UQs.
-For a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$ and $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, let us recall the definition of the moment matrix $M(\mathbf{S})$ and localizing moment matrix $M(H\mathbf{S})$.
-Definition 4.1.
-Given a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, the associated moment matrix $M(\mathbf{S})$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}$ and the $({\boldsymbol{\alpha}},{\boldsymbol{\beta}})$-th block entry is $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$.
-For $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, the localizing matrix $M(H\mathbf{S})$ associated to $\mathbf{S}$ and $H$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}$ and the $({\boldsymbol{\alpha}},{\boldsymbol{\beta}})$-th block entry is $\sum_{{\boldsymbol{\gamma}}\in\mathrm{supp}(H)}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\gamma}}}\otimes H_{{\boldsymbol{\gamma}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$.
-For $d\in\mathbb{N}$, the $d$-th order moment matrix $M_{d}(\mathbf{S})$ (resp.
-localizing matrix $M_{d}(H\mathbf{S})$) is the submatrix of $M(\mathbf{S})$ (resp.
-$M(H\mathbf{S})$) whose block row and block column are both indexed by $\mathbb{N}^{n}_{d}$.
-Similar to the scalar case in [18], we give the definition of localizing moment matrix associated with $(\mathbf{S},G,\nu)$, where we write $G({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}G_{{\boldsymbol{\gamma}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\gamma}}}$ with $G_{{\boldsymbol{\gamma}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
-Definition 4.2.
-Given a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ and a Borel measure $\nu$ on $\mathcal{Y}$, the localizing matrix $M^{\nu}(G{\mathbf{S}})$ associated to ${\mathbf{S}}$, $G$ and $\nu$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}\times\mathbb{N}^{m}$ and the $(({\boldsymbol{\alpha}},{\boldsymbol{\eta}}),({\boldsymbol{\beta}},{\boldsymbol{\xi}}))$-th block entry is
-$\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)$
-for all $({\boldsymbol{\alpha}},{\boldsymbol{\eta}}),({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}\times\mathbb{N}^{m}$.
-For $d,k\in\mathbb{N}$, the $(d,k)$-th order localizing matrix $M^{\nu}_{d,k}(G{\mathbf{S}})$ is the submatrix of $M^{\nu}(G{\mathbf{S}})$ whose block row and block column are both indexed by $\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}$.
-Remark 4.1.
-Throughout the paper, for a sequence indexed by $\mathbb{N}^{n}\times\mathbb{N}^{m}$ (resp., $\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}$), the indices are arranged according to the order of the exponents in the monomial basis $[{\boldsymbol{x}}]_{\infty}\otimes[{\boldsymbol{y}}]_{\infty}$ (resp., $[{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}$).
-With the definition of the localizing matrix $M_{d,k}^{\nu}(G{\mathbf{S}})$, we immediately have the following result.
-Proposition 4.1.
-Given a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, for any SOS matrix $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d$ and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$,
-$\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0$
-if and only if $M^{\nu}_{d,k}(G{\mathbf{S}})\succeq 0$.
-Proof.
-For the polynomial matrix $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in\mathbf{S}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, there exists some $Q\in{\mathbb{S}}_{+}^{pq|\mathbb{N}^{n}_{d}||\mathbb{N}^{m}_{k}|}$ satisfying
-$\Sigma({\boldsymbol{x}},{\boldsymbol{y}})$ $=(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq})^{\intercal}Q(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq}).$
-Then, we have
-$\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)=\langle M^{\nu}_{d,k}(G{\mathbf{S}}),Q\rangle,$
-which implies the desired conclusion. ∎
-We now provide the following result for the exponentially bounded matrix-valued moment problem, which generalizes the result of Berg and Maserick [2].
-The proof of Theorem 4.3 is postponed to Section 6.5.
-Theorem 4.3.
-Suppose that Assumption 3.2 holds. Then for a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ and $C>0$, the following are equivalent:
-  1. (i)
-    $\mathbf{S}$ has a matrix-valued representing measure supported on $\mathcal{X}\cap[-C,C]^{n}$;
-  2. (ii)
-    $M(\mathbf{S})\succeq 0$, $M^{\nu}(G\mathbf{S})\succeq 0$ and there is a constant $C_{0}>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-In case that UQs are not present, we obtain the following corollary.
+In analogy with Corollary 4.1, we can obtain the following inhomogeneous counterpart of Theorem 4.2. We omit the similar proof for the sake of clarity.
 Corollary 4.3.
-For a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ and $C>0$, the following are equivalent:
-  1. (i)
-    $\mathbf{S}$ admits a matrix-valued representing measure supported on
-    $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\}\cap[-C,C]^{n},$
-    where $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, $q\in\mathbb{N}$;
-  2. (ii)
-    $M(\mathbf{S})\succeq 0$, $M(H\mathbf{S})\succeq 0$ and there is a constant $C_{0}>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-4.3.2. Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs
-Inspired by the result of Lasserre and Netzer [30], we next derive a Positivstellensatz to characterize a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ that is PSD on $\mathcal{X}\cap[-1,1]^{n}$ using high-degree perturbations.
-For $d\in\mathbb{N}$, let $\Theta_{d}\coloneqq 1+\sum_{i=1}^{n}x_{i}^{2d}$.
-Let $d(G)\coloneqq\lceil\deg_{{\boldsymbol{x}}}G/2\rceil$. Given $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, let us consider the following optimization problem:
-$r_{d,k}^{\star}\coloneqq\left\{\begin{aligned} \inf_{r,\Sigma_{0},\Sigma}&\ r\\ \text{s.t.}\,\,&\ F({\boldsymbol{x}})+r\Theta_{d}I_{p}=\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\ \text{are SOS},\\ &\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G),\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2k,\end{aligned}\right.$ (6)
-and it dual problem reads as
-$\sup_{\mathbf{S}}&\ -\mathscr{L}_{\mathbf{S}}(F)\\ \text{s.t.}&\ \mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}}\subset{\mathbb{S}}^{p},\ \mathscr{L}_{\mathbf{S}}(\Theta_{d}I_{p})\leq 1,\\ &\ M_{d}(\mathbf{S})\succeq 0,\ M_{d-d(G),k}^{\nu}(G\mathbf{S})\succeq 0.\end{aligned}\right.$ (7)
-As the sequence of zero matrices is feasible to (7), it holds $r^{\star}_{d,k}\geq 0$ by the weak duality.
-Assumption 4.1.
-There are positive number $\lambda>0$, and open and bounded subsets $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ of $\mathcal{X}$ and $\mathcal{Y}$, respectively, such that $G({\boldsymbol{x}},{\boldsymbol{y}})\succeq\lambda I_{q}$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$.
-The following proposition shows that if Assumption 4.1 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$, then the optimal value $r^{\star}_{d,k}\to 0$ as $d,k\to\infty$.
-Its proof is postponed to Section 6.6.
-Proposition 4.2.
-Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$.
-Then, for any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that $r_{d,k}^{\star}\leq\varepsilon$ for all $d\geq\ell_{1},k\geq\ell_{2}$.
-We can now give the Lasserre-Netzer-type Positivstellensatz for polynomial matrices with UQs.
-Theorem 4.4.
-Suppose that Assumptions 3.2 and 4.1 hold, and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$.
-For any $\varepsilon>0$, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1},k\geq\ell_{2}$, it holds
-$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=\Sigma_{0}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
-for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$ and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$.
-Proof.
-By Proposition 4.2, there exist $\ell_{1},\ell_{2}\in\mathbb{N}$ such that for all $d\geq\ell_{1}$ and $k\geq\ell_{2}$, there is a feasible point $(r,\Sigma_{0}^{\prime},\Sigma)$ of (6) with $r\leq\varepsilon$.
-We have
-$F({\boldsymbol{x}})+r\Theta_{d}I_{p}=\Sigma_{0}^{\prime}+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
-with $\deg(\Sigma_{0}^{\prime})\leq 2d$, $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d-2d(G)$, and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$. Then, for all $d\geq\ell_{1}$ and $k\geq\ell_{2}$, it holds
-$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=F({\boldsymbol{x}})+r\Theta_{d}I_{p}+(\varepsilon-r)\Theta_{d}I_{p}=(\Sigma_{0}^{\prime}+(\varepsilon-r)\Theta_{d}I_{p})+\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-The conclusion then follows. ∎
-In case that UQs are not present, we obtain the following corollary.
-Corollary 4.4.
-Let $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$ and $\mathcal{W}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\}.$ Suppose that there are positive number $\lambda>0$, and open and bounded subset $\mathcal{O}$ of $\mathcal{X}$ such that $H({\boldsymbol{x}})\succeq\lambda I_{q}$ on $\mathcal{O}$.
-If $F({\boldsymbol{x}})$ is PSD on $\mathcal{W}\cap[-1,1]^{n}$, then for any $\varepsilon>0$, there exists some $\ell\in\mathbb{N}$ such that for all $d\geq\ell$, it holds
-$F({\boldsymbol{x}})+\varepsilon\Theta_{d}I_{p}=\Sigma_{0}+\left\langle\Sigma,G\right\rangle_{p}$
-for some SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d}^{p}$ and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}}]_{2d-2d(G)}^{pq}$.
+Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$, $F({\boldsymbol{x}})$ is PSD on ${\mathbb{R}}^{n}_{+}\cap\mathcal{X}$, and $\mathcal{Y}$ is compact.
+Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$ and polynomial matrices
+$P_{0}=\sum_{|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}-2d_{G}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
+satisfying $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}-2d_{G}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, and
+$(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{\varepsilon}}\left(F({\boldsymbol{x}})+\varepsilon(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2d_{F}}I_{p}\right)=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
 5. Applications to robust PMI constrained optimization
 Verifying PMIs over a prescribed set has a wide range of applications in many fields.
-For instance, many control problems for systems of ordinary differential equations can be formulated as convex optimization problems with PMI constraints that must be satisfied over a specified portion of the state space [15, 16, 19, 40, 51].
+For example, many control problems for systems of ordinary differential equations can be formulated as convex optimization problems with PMI constraints that must be satisfied over a specified portion of the state space [13, 15, 17, 37, 47].
 These problems are typically formulated as follows:
 $\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t.
-}\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})\coloneqq P_{0}({\boldsymbol{x}})-\sum_{i=1}^{r}P_{i}({\boldsymbol{x}})\gamma_{i}\succeq 0,\ \forall{\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\},$
-where ${\boldsymbol{c}}\in{\mathbb{R}}^{r}$, $P_{0},\ldots,P_{r}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
-However, due to estimation errors or lack of information, the data of real-world problems often involve uncertainty.
+}\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\coloneqq\Lambda_{0}({\boldsymbol{x}})-\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{x}})\gamma_{i}\succeq 0,\ \forall{\boldsymbol{x}}\in\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid H({\boldsymbol{x}})\succeq 0\},$
+where ${\boldsymbol{c}}\in{\mathbb{R}}^{r}$, $\Lambda_{0},\ldots,\Lambda_{r}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
+However, due to estimation errors or lack of information, data from real-world problems often involve uncertainty.
 Therefore, ensuring the robustness of PMIs over a prescribed set under uncertainty is a critical issue, which could be formulated as the following robust optimization problem:
 $\tau^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t.
-}\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\},$ (8)
+}\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\coloneqq\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G({\boldsymbol{x}},{\boldsymbol{y}})\succeq 0,\ \forall{\boldsymbol{y}}\in\mathcal{Y}\},$ (8)
 where $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is closed.
 Consequently, Positivstellensätze for polynomial matrices with UQs developed in this paper serve as a powerful mathematical tool for addressing this issue.
-Indeed, by leveraging the SOS-structured certificates provided by Positivstellensätze for positive definiteness of $P({\boldsymbol{x}},{\boldsymbol{\gamma}})$ over $\mathcal{X}$, we are able to establish converging hierarchies of SDP relaxations for the problem (8).
-Concretely, if $\mathcal{X}$ is compact, then one could apply the Scherer-Hol-type Positivestellensatz (Theorem 3.1) to construct a hierarchy of SDP relaxations for (8):
-$\tau_{k}\coloneqq\inf_{{\boldsymbol{\gamma}},\Sigma_{0},\Sigma}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ P({\boldsymbol{x}},{\boldsymbol{\gamma}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg\Sigma_{0},\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2k,\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2k-\deg_{{\boldsymbol{x}}}G.\end{aligned}\right.$
-It is clear that the sequence $(\tau_{k})_{k\in\mathbb{N}}$ is non-increasing upper bounds of $\tau^{\star}$.
-Assuming that the Slater condition holds for (8), i.e., there exists a point $\bar{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$ such that $P({\boldsymbol{x}},\bar{{\boldsymbol{\gamma}}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$, the convergence of $\tau_{k}\to\tau^{\star}$ as $k\to\infty$ is guaranteed under Assumptions 2.1 and 3.2 by Theorem 3.1.
-Other Positivstellensätze presented in this paper can also be employed to derive the corresponding hierarchies of SDP relaxations for (8).
-The study of convergence of these hierarchies provides an intriguing avenue for future research.
-For instance, if $\mathcal{X}$ is non-compact, then one could apply the Putinar–Vasilescu-type Positivstellensatz for the inhomogeneous case (Corollary 4.1) with a fixed $\varepsilon>0$ to construct the following hierarchy of SDP relaxations for (8) ($d_{P}\coloneqq\max_{i,j\in[p]}\lfloor\deg_{{\boldsymbol{x}}}(P_{ij})/2\rfloor+1$):
-$\tau_{k}(\varepsilon)\coloneqq\inf_{{\boldsymbol{\gamma}},\Sigma_{0},\Sigma}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \theta^{k}(P({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon\theta^{d_{P}}I_{p})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg\Sigma_{0},\ \deg_{{\boldsymbol{y}}}\Sigma\leq 2(k+d_{P}),\ \deg_{{\boldsymbol{x}}}\Sigma\leq 2(k+d_{P}-d_{G}).\end{aligned}\right.$
-Assuming that $\tau^{\star}$ is attainable and Assumption 3.2 holds, by Corollary 4.1, there exists $\bar{k}\in\mathbb{N}$ such that $\tau_{k}(\varepsilon)\leq\tau^{\star}$ for all $k\geq\bar{k}$.
-How to bound the number $\bar{k}$ from above, and whether $\tau_{k}(\varepsilon)\to\tau^{\star}$ as $\varepsilon\to 0$ and $k\to\infty$ under certain conditions, could be interesting topics for further study.
+Indeed, by leveraging the SOS-structured certificates provided by Positivstellensätze for the positive definiteness of $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})$ over $\mathcal{X}$, we are able to establish hierarchies of SDP relaxations for the problem (8).
+5.1. The compact case
+If $\mathcal{X}$ is compact, then one could apply Theorem 3.1 to construct a hierarchy of SDP relaxations for (8):
+$\left\{\begin{aligned} \tau(k)\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg S_{0}\leq 2k,\ \deg_{{\boldsymbol{x}}}S\leq 2(k-\lceil\deg_{{\boldsymbol{x}}}G/2\rceil),\ \deg_{{\boldsymbol{y}}}S\leq 2(k-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil).\end{aligned}\right$.
+(9)
+It is clear that $\{\tau(k)\}_{k\in\mathbb{N}}$ is a sequence of non-increasing upper bounds on $\tau^{\star}$.
+Theorem 5.1.
+Suppose that Assumptions 3.1 and 3.2 hold and there exists a point $\bar{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$ such that $\Lambda({\boldsymbol{x}},\bar{{\boldsymbol{\gamma}}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$.
+Then we have $\lim_{k\to\infty}\tau(k)=\tau^{\star}$.
+Proof.
+Note that for any feasible point $\boldsymbol{\gamma}$$^{\prime}$ of (8) and any $\varepsilon>0$, the point $(1-\varepsilon){\boldsymbol{\gamma}}^{\prime}+\varepsilon\bar{{\boldsymbol{\gamma}}}$ satisfies $\Lambda({\boldsymbol{x}},(1-\varepsilon){\boldsymbol{\gamma}}^{\prime}+\varepsilon\bar{{\boldsymbol{\gamma}}})\succ 0$ on $\mathcal{X}$.
+Then, the conclusion follows by Theorem 3.1.
+∎
+Given that the SDP problem derived from (9) scales rapidly and becomes computationally prohibitive as $k$ increases, we revise it by imposing separate degree constraints on $\boldsymbol{x}$} and $\boldsymbol{y}$}, leading to the following problem
+$\left\{\begin{aligned} \tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg S_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G/2\rceil),\ \deg_{{\boldsymbol{y}}}S\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil).\end{aligned}\right$.
+(10)
+If $\mathcal{Y}$ is low-dimensional, we can maintain a small $k_{{\boldsymbol{x}}}$ while increasing $k_{{\boldsymbol{y}}}$ separately; this allows us to solve (10) in reasonable time and obtain a nontrivial upper bound $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ of $\tau^{\star}$.
+In practice, the SDP relaxation (10) requires an efficient evaluation of the integral $\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\beta}}}\mathrm{d}\nu({\boldsymbol{y}})$.
+If $\nu$ is chosen to be the Lebesgue measure on $\mathcal{Y}$ (as in the following illustrating examples), this integral is available in closed form or by numerical methods whenever $\mathcal{Y}$ is a simple domain; examples include hypercubes, spheres [9], balls [9], and polytopes [2].
+All numerical experiments in the sequel were carried out on a PC with an 8-Core Intel i7 3.8GHz CPU and 8G RAM, using the SDP solver MOSEK [1] and a customized version of the MATLAB optimization toolbox YALMIP [28].
+Example 5.1.
+Consider the feasible set of the problem (8):
+$\mathcal{F}=\{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}\mid\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\forall{\boldsymbol{x}}\in\mathcal{X}\}.$
+In this example, we illustrate how Theorem 3.1 enables one to approximate the convex set of vectors $\boldsymbol{\gamma}$} for which $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})$ is PSD on $\mathcal{X}$ – that is, to obtain approximate semidefinite representations of $\mathcal{F}$.
+For any $k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}\in\mathbb{N}$ with $k_{{\boldsymbol{x}}}\geq\lceil\deg_{{\boldsymbol{x}}}G/2\rceil$ and $k_{{\boldsymbol{y}}}\geq\lceil\deg_{{\boldsymbol{y}}}G/2\rceil$, let $\nu$ in (10) be the Lebesgue measure on $\mathcal{Y}$ and consider
+$\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})=\left\{\begin{array}[]{c|c}{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}&\begin{aligned} &\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}{\boldsymbol{y}},\\ &S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\deg S_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G/2\rceil),\ \deg_{{\boldsymbol{y}}}S\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil).\end{aligned}\end{array}\right\}.$
+Clearly, $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ is an inner approximation of $\mathcal{F}$.
+On the other hand, Theorem 3.1 guarantees that any ${\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}$ for which $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})$ is PD on $\mathcal{X}$ belongs to $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ for sufficiently large $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$.
+Therefore, the sets $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ can approximate $\mathcal{F}$ from inside arbitrarily accurately in the sense that any compact subset of the interior of $\mathcal{F}$ is included in $\mathcal{F}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ for some sufficiently large integers $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$.
+Let us construct an illustrating example. Let $n=r=2$, $m=1$, $\mathcal{Y}=[0,1]$,
+$G({\boldsymbol{x}},y)=\left[\begin{array}[]{cc}(1/2-y)^{2}&y(x_{1}^{2}+x_{2}^{2}-1)\\ y(x_{1}^{2}+x_{2}^{2}-1)&x_{1}+x_{2}-1\end{array}\right],$
+and
+$\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=\left[\begin{array}[]{cc}1-(x_{1}\gamma_{1}-x_{2}\gamma_{2}+1/2)&2(x_{1}\gamma_{2}+x_{2}\gamma_{1})\\ 2(x_{1}\gamma_{2}+x_{2}\gamma_{1})&1+(x_{1}\gamma_{1}-x_{2}\gamma_{2}+1/2)\end{array}\right].$
+It is clear that
+$\mathcal{X}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{2}\mid x_{1}^{2}+x_{2}^{2}=1,\ x_{1}x_{2}\geq 0\}.$
+Geometrically, the set $\mathcal{F}$ is constructed as the intersection of all rotated copies of the shape defined by $(\gamma_{1}+1/2)^{2}+4\gamma_{2}^{2}\leq 1$ in the $\boldsymbol{\gamma}$}-plane as it is continuously rotated clockwise by $90^{\circ}$ about the origin.
+The resulting set $\mathcal{F}$ (light blue region with blue outline) is shown in Figure 1.
+Three inner approximations – $\mathcal{F}(1,1)$, $\mathcal{F}(1,2)$, and $\mathcal{F}(1,6)$ – each depicted with a red outline, appear in Figures 1.
+These red outlines were obtained by solving and extracting optimal solutions of (8) with objective functions $\gamma_{1}\cos\theta+\gamma_{2}\sin\theta$ over $80$ equispaced values of $\theta$ in the interval $0,2\pi]$.
+Figure 1. The region $\mathcal{F}$ and three inner approximations $\mathcal{F}(1,1)$, $\mathcal{F}(1,2)$, $\mathcal{F}(1,6)$ in Example 5.1.
+5.2. Exploiting sparsity
+We illustrate the computational advantage of our sparsity-exploiting Positivstellensatz when the correlative sparsity pattern in Assumption 3.3 is satisfied. Consider the problem
+$f^{\star}\coloneqq\min\ f({\boldsymbol{x}})\ \text{ s.t. }{\boldsymbol{x}}\in\widehat{\mathcal{X}}$,
+where $f({\boldsymbol{x}})\in{\mathbb{R}}[{\boldsymbol{x}}]$ and the set $\widehat{\mathcal{X}}$ is defined in (4). By Theorem 3.1, dense SDP relaxations for such a problem can be formulated as
+$\left\{\begin{aligned} \tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\sup_{\gamma,\sigma_{0},S_{j}}&\ \gamma\\ \text{s.t.}&\ f({\boldsymbol{x}})-\gamma=\sigma_{0}({\boldsymbol{x}})+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S_{j}({\boldsymbol{x}},{\boldsymbol{y}}),G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ \sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}],S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}},\,j\in[s],\,\ \ \text{are SOS},\\ &\ \deg\sigma_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S_{j}\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G_{j}/2\rceil),\ \deg_{{\boldsymbol{y}}}S_{j}\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G_{j}/2\rceil).\end{aligned}\right$.
+(11)
+For any $k_{{\boldsymbol{x}}}\geq\lceil\deg_{{\boldsymbol{x}}}G_{j}/2\rceil$ and $k_{{\boldsymbol{y}}}\geq\lceil\deg_{{\boldsymbol{y}}}G_{j}/2\rceil$, the number $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ provides a lower bound of $f^{\star}$.
+The sequence $\{\tau(k,k)\}_{k\in\mathbb{N}}$, the limit $\tau(\infty,\infty)\coloneqq\lim_{k\to\infty}\tau(k,k)$ exists.
+By Theorem 5.1, we have $\tau(\infty,\infty)=\tau^{\star}$ if Assumptions 3.1 and 3.2 hold.
+Provided the presence of the correlative sparsity pattern stated in Assumption 3.3, sparse SDP relaxations for such a problem can be constructed as follows:
+$\left\{\begin{aligned} \tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\sup_{\gamma,\sigma_{0},S_{j}}&\ \gamma\\ \text{s.t.}&\ f({\boldsymbol{x}})-\gamma=\sum_{\ell=1}^{t}\left(\sigma_{\ell,0}({\boldsymbol{x}}(\mathcal{I}_{\ell}))+\sum_{j\in\mathcal{J}_{\ell}}\int_{\mathcal{Y}}\left\langle S_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}),G_{j}({\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}})\right\rangle\mathrm{d}\nu({\boldsymbol{y}})\right),\\ &\ \sigma_{\ell,0}\in{\mathbb{R}}[{\boldsymbol{x}}(\mathcal{I}_{\ell})],S_{j}\in{\mathbb{S}}[{\boldsymbol{x}}(\mathcal{I}_{\ell}),{\boldsymbol{y}}]^{q_{j}},\ j\in\mathcal{J}_{\ell},\ \ell\in[t],\,\text{are SOS},\\ &\ \deg\sigma_{\ell,0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S_{j}\leq 2(k_{{\boldsymbol{x}}}-\lceil\deg_{{\boldsymbol{x}}}G_{j}/2\rceil),\ \deg_{{\boldsymbol{y}}}S_{j}\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G_{j}/2\rceil).\end{aligned}\right$.
+(12)
+For the sequence $\{\tau^{\mbox{\tiny cs}}(k,k)\}_{k\in\mathbb{N}}$, if the assumptions in Theorem 3.2 hold, we have the convergence
+$\tau^{\mbox{\tiny cs}}(\infty,\infty)\coloneqq\lim_{k\to\infty}\tau^{\mbox{\tiny cs}}(k,k)=\tau^{\star}.$
+When the number $\max\{|\mathcal{I}_{\ell}|,\ell\in[t]\}$ is small, dramatic computational savings can be expected.
+Example 5.2.
+Consider the problem $f^{\star}\coloneqq\min_{{\boldsymbol{x}}\in\widehat{X}}f({\boldsymbol{x}})$ where $f({\boldsymbol{x}})=\sum_{i=1}^{n}(x_{i}+1)^{4}$,
+$\widehat{\mathcal{X}}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid G_{j}({\boldsymbol{x}},y)\succeq 0,\ \forall j\in[n-1],\ y\in\mathcal{Y}\},\quad G_{j}({\boldsymbol{x}},y)=\left[\begin{array}[]{cc}1-yx_{j}&x_{j+1}\\ x_{j+1}&1+yx_{j}\end{array}\right],\ j\in[n-1],$
+and $\mathcal{Y}=[0,1]$. Clearly, Assumption 3.3 is satisfied with t=n1t=n-1, $\mathcal{I}_{\ell}=\{\ell,\ell+1\}$ and $\mathcal{J}_{\ell}=\{\ell\}$ for $\ell\in[n-1]$. It is easy to see that
+$\widehat{\mathcal{X}}=\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid x_{i}^{2}+x_{i+1}^{2}\leq 1,\ i\in[n-1]\}.$
+Since the problem $\min_{{\boldsymbol{x}}\in\widehat{X}}f({\boldsymbol{x}})$ is convex, numerical values for $f^{\star}$ can be obtained by solving the associated KKT system.
+These results can then be used to evaluate the accuracy of the SDP lower bounds $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
+We choose $\nu$ in the SDP relaxations (11) and (12) to be the Lebesgue measure on $\mathcal{Y}$ and present the numerical results in Table 3, where the symbol ‘–/–’ indicates that MOSEK runs out of memory.
+From the table, we could see that dramatic computational advantages are gained by exploiting correlative sparsity.
+∎
+Table 3. Computational results for Example 5.2.
+$n$ $f^{\star}$ $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 10) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 15)
+$\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau^{\mbox{\tiny cs}}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time
+33 0.0206 0.0184/13.1s 0.0184/4.45s 0.0191/83.3s 0.0191/19.9s
+44 0.0294 0.0262/65.3s 0.0262/9.23s –/– 0.0273/36.5s
+55 0.0359 –/– 0.0320/13.9s –/– 0.0333/45.1s
+66 0.0441 –/– 0.0394/19.2s –/– 0.0409/62.5s
+5.3. The non-compact case
+Let us consider the problem (8) with a non‑compact set $\mathcal{X}$.
+For such a problem, after fixing a constant $\varepsilon>0$, we can apply Corollary 4.1 to construct the following hierarchy of SDP relaxations:
+$\left\{\begin{aligned} \tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ \theta^{k_{{\boldsymbol{x}}}-d_{\Lambda}}(\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon\theta^{d_{\Lambda}}I_{p})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\,\,\text{are SOS matrices},\\ &\ \deg S_{0}\leq 2k_{{\boldsymbol{x}}},\ \deg_{{\boldsymbol{x}}}S\leq 2(k_{{\boldsymbol{x}}}-d_{G}),\ \deg_{{\boldsymbol{y}}}S\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil),\end{aligned}\right$.
+(13)
+where $d_{\Lambda}\coloneqq\max_{i,j\in[p]}\lfloor\deg_{{\boldsymbol{x}}}(\Lambda_{ij})/2\rfloor+1$.
+Let $\mathcal{F}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ denote the feasible set of (13).
+For the sequence $\{\tau_{\varepsilon}(\ell,\ell)\}_{\varepsilon>0,\ell\in\mathbb{N}}$, let $\tau_{0}(\infty,\infty)\coloneqq\lim_{\varepsilon\to 0}\lim_{\ell\to\infty}\tau_{\varepsilon}(\ell,\ell)$.
+Theorem 5.2.
+Let Assumption 3.2 hold.
+Then the limit $\tau_{0}(\infty,\infty)$ exists, and $\tau_{0}(\infty,\infty)\leq\tau^{\star}$.
+Moreover, if the optimal solution set $\mathcal{S}$ of (8) is nonempty and bounded, we have $\tau_{0}(\infty,\infty)=\tau^{\star}$.
+Proof.
+Since Assumption 3.2 holds, Corollary 4.1 implies that, for any $\varepsilon>0$ and sufficiently large $\ell$, $\mathcal{F}_{\varepsilon}(\ell,\ell)$ contains the feasible set $\mathcal{F}$ of (8) and hence $\tau_{\varepsilon}(\ell,\ell)\leq\tau^{\star}$.
+It is clear that $\tau_{\varepsilon}(\ell,\ell)$ is non-increasing as $\ell\to\infty$ for fixed $\varepsilon>0$ and is non-decreasing as $\varepsilon\to 0$ for fixed $\ell$.
+Hence, the limit $\tau_{0}(\infty,\infty)$ exists and $\tau_{0}(\infty,\infty)\in[-\infty,\tau^{\star}]$.
+Assuming that the optimal solution set $\mathcal{S}$ of (8) is nonempty and bounded, we show that $\tau_{0}(\infty,\infty)=\tau^{\star}$.
+Suppose on the contrary that $\tau_{0}(\infty,\infty)<\tau^{\star}$.
+Then, there exists a sequence $\{\varepsilon_{j}\}_{j\in\mathbb{N}}$ with $\lim_{j\to\infty}\varepsilon_{j}=0$ and a sequence $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ of feasible points of (13) with $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})=(\varepsilon_{j},\ell_{j},\ell_{j})$ such that for each $j\in\mathbb{N}$, ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}\leq\tau^{\star}-\delta$ for some $\delta>0$.
+We claim that $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ is bounded.
+Suppose that $\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|\to\infty$ as $j\to\infty$ and without loss of generality that $\lim_{j\to\infty}{\boldsymbol{\gamma}}^{(\ell_{j})}/\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|={\boldsymbol{v}}$ for some ${\boldsymbol{v}}=(v_{1},\ldots,v_{r})\in{\mathbb{R}}^{r}$ with $\|{\boldsymbol{v}}\|=1$.
+Since ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}$ is bounded by $\tau^{\star}-\delta$ for all $j\in\mathbb{N}$, we have ${\boldsymbol{c}}^{\intercal}{\boldsymbol{v}}=0$.
+Moreover, for arbitrarily fixed ${\boldsymbol{u}}\in\mathcal{X}$, we have
+$\Lambda({\boldsymbol{u}},{\boldsymbol{\gamma}}^{(\ell_{j})})+\varepsilon_{j}\theta({\boldsymbol{u}})^{d_{\Lambda}}I_{p}=\Lambda_{0}({\boldsymbol{u}})-\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{u}})\gamma^{(\ell_{j})}_{i}+\varepsilon_{j}\theta({\boldsymbol{u}})^{d_{\Lambda}}I_{p}\succeq 0,\ \forall j\in\mathbb{N}.$ (14)
+Dividing the left-hand side of (14) by $\|{\boldsymbol{\gamma}}^{(\ell_{j})}\|$ and letting $j\to\infty$, we get $\sum_{i=1}^{r}\Lambda_{i}({\boldsymbol{u}})v_{i}\succeq 0$.
+Since $\boldsymbol{u}$} is an arbitrary point in $\mathcal{X}$ and ${\boldsymbol{c}}^{\intercal}{\boldsymbol{v}}=0$, for any point ${\boldsymbol{\gamma}}^{\star}\in\mathcal{S}$, we have ${\boldsymbol{\gamma}}^{\star}+t{\boldsymbol{v}}\in\mathcal{S}$ for all $t>0$, which contradicts the boundedness of $\mathcal{S}$.
+As we have proved that $\{{\boldsymbol{\gamma}}^{(\ell_{j})}\}_{j\in\mathbb{N}}$ is bounded, we can assume without loss of generality that $\lim_{j\to\infty}{\boldsymbol{\gamma}}^{(\ell_{j})}=\hat{{\boldsymbol{\gamma}}}$ for some $\hat{{\boldsymbol{\gamma}}}\in{\mathbb{R}}^{r}$.
+Then letting $j\to\infty$ in (14) yields the result that $\hat{{\boldsymbol{\gamma}}}$ is feasible to (8).
+Thus, we have
+$\tau^{\star}\leq{\boldsymbol{c}}^{\intercal}\hat{{\boldsymbol{\gamma}}}=\lim_{j\to\infty}{\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}^{(\ell_{j})}\leq\tau^{\star}-\delta,$
+a contradiction. ∎
+Now, let us consider the problem
+$\tilde{\tau}^{\star}\coloneqq\inf_{{\boldsymbol{\gamma}}\in{\mathbb{R}}^{r}}\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\ \text{ s.t.
+}\ \Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0,\ \forall{\boldsymbol{x}}\in\mathcal{X}\cap{\mathbb{R}}_{+}^{n},$ (15)
+where $\mathcal{X}$ is non-compact and defined in (8), and $\mathcal{Y}$ is defined in (7).
+After fixing a constant $\varepsilon>0$, we apply Corollary 4.3 to construct the following hierarchy of SDP relaxations:
+$\left\{\begin{aligned} \tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})\coloneqq\inf_{{\boldsymbol{\gamma}},S_{0},S}&\ {\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}\\ \text{s.t.}&\ (1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2k_{{\boldsymbol{x}}}-2d_{\Lambda}}(\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})+\varepsilon(1+{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{2d_{\Lambda}}I_{p})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),\\ &\ P_{0}=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2k_{{\boldsymbol{x}}}}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}\text{ with each }P_{0,{\boldsymbol{\alpha}}}\succeq 0,\\ &\ P=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2(k_{{\boldsymbol{x}}}-d_{G})}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq},\ \deg_{{\boldsymbol{y}}}P\leq 2(k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil),\\ &\ \text{ with each }P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}^{pq}_{k_{{\boldsymbol{y}}}-\lceil\deg_{{\boldsymbol{y}}}G/2\rceil}(\mathbf{H}).\end{aligned}\right$.
+(16)
+For the sequence $\{\tilde{\tau}_{\varepsilon}(\ell,\ell)\}_{\varepsilon>0,\ell\in\mathbb{N}}$, let $\tilde{\tau}_{0}(\infty,\infty)\coloneqq\lim_{\varepsilon\to 0}\lim_{\ell\to\infty}\tilde{\tau}_{\varepsilon}(\ell,\ell)$.
+Recall the set $\mathbf{H}$ in Corollary 4.2.
+Similarly to Theorem 5.2, we have the following convergence result.
+Theorem 5.3.
+Let the Archimedean condition hold for the quadratic module $\mathcal{Q}^{pq}(\mathbf{H})$.
+Then the limit $\tilde{\tau}_{0}(\infty,\infty)$ exists and $\tilde{\tau}_{0}(\infty,\infty)\leq\tau^{\star}$.
+Moreover, if the optimal solution set $\mathcal{S}$ of (15) is nonempty and bounded, we have $\tilde{\tau}_{0}(\infty,\infty)=\tau^{\star}$.
+Note that we set the same degree bounds on $\boldsymbol{x}$} and $\boldsymbol{y}$} in the representations of $\Lambda({\boldsymbol{x}},{\boldsymbol{y}})$ in (13) and (16), so that we can compare the behaviors of $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ when applying both (13) and (16) to (8) if $\mathcal{X}=\mathcal{X}\cap{\mathbb{R}}^{n}_{+}$.
+Example 5.3.
+Consider the problem (8) where we let $m=1$, $n=r=2$, $\mathcal{Y}=[0,1]$,
+$G({\boldsymbol{x}},y)=Q_{1}^{\intercal}\left[\begin{array}[]{cc}x_{1}+x_{2}&0\\ 0&-2y(1-y)+(1-y)^{2}x_{1}+y^{2}x_{2}\end{array}\right]Q_{1},$
+and
+$\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})=Q_{2}^{\intercal}\left[\begin{array}[]{cc}-2+\left(-\frac{\gamma_{1}}{2}+\frac{3\gamma_{2}}{2}+3\right)x_{2}+\left(\frac{3\gamma_{1}}{2}-\frac{\gamma_{2}}{2}+3\right)x_{1}&0\\ 0&-2-\gamma_{1}x_{2}-\gamma_{2}x_{1}\end{array}\right]Q_{2}.$
+Here, $Q_{1}$ and $Q_{2}$ are two fixed arbitrary $2\times 2$ orthogonal matrices.
+For any fixed $(x_{1},x_{2})$, consider the function
+$f(y)=-2y(1-y)+(1-y)^{2}x_{1}+y^{2}x_{2}=(x_{1}+x_{2}+2)y^{2}-2(x_{1}+1)y+x_{1}.$
+The evaluations $f(0)=x_{1}$ and $f(1)=x_{2}$ imply that $x_{1}\geq 0$ and $x_{2}\geq 0$. Assuming $x_{1},x_{2}\geq 0$, the minimum of $f(y)$ is
+$f\left(\frac{x_{1}+1}{x_{1}+x_{2}+2}\right)=\frac{x_{1}x_{2}-1}{x_{1}+x_{2}+2}\ \text{ where }\ \frac{x_{1}+1}{x_{1}+x_{2}+2}\in\mathcal{Y}.$
+Thus we have
+$\mathcal{X}=\{(x_{1},x_{2})\in{\mathbb{R}}^{2}\mid x_{1}\geq 0,\ x_{2}\geq 0,\ x_{1}x_{2}\geq 1\}.$
+Then similarly, we can show that $\Lambda({\boldsymbol{x}},{\boldsymbol{\gamma}})\succeq 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\cap{\mathbb{R}}_{+}^{n}$ if and only if
+$2+\left(-\frac{\gamma_{1}}{2}+\frac{3\gamma_{2}}{2}+3\right)x_{2}+\left(\frac{3\gamma_{1}}{2}-\frac{\gamma_{2}}{2}+3\right)x_{1}\geq 0\ \text{ and }\ -2-\gamma_{1}x_{2}-\gamma_{2}x_{1}\geq 0$
+on $\mathcal{X}$ if and only if
+$\frac{\gamma_{1}}{2}-\frac{3\gamma_{2}}{2}-3\leq 0,\ -\frac{3\gamma_{1}}{2}+\frac{\gamma_{2}}{2}-3\leq 0,\ \phi_{1}({\boldsymbol{\gamma}})\coloneqq\left(\frac{\gamma_{1}}{2}-\frac{3\gamma_{2}}{2}-3\right)\left(-\frac{3\gamma_{1}}{2}+\frac{\gamma_{2}}{2}-3\right)-1\geq 0,$
+and
+$\gamma_{1}\leq 0,\ \gamma_{2}\leq 0,\ \phi_{2}({\boldsymbol{\gamma}})\coloneqq\gamma_{1}\gamma_{2}-1\geq 0.$
+The feasible set $\mathcal{F}$ is shown in light blue in Figure 2.
+Since $\mathcal{X}=\mathcal{X}\cap{\mathbb{R}}^{2}_{+}$, we can apply both SDP relaxations (13) and (16) to this problem and compare the behaviors of $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
+Consider the objective function ${\boldsymbol{c}}^{\intercal}{\boldsymbol{\gamma}}=\gamma_{1}-\gamma_{2}$.
+One can check that the optimal solution is one of the solutions of $\phi_{1}({\boldsymbol{\gamma}})=\phi_{2}({\boldsymbol{\gamma}})=0$:
+${\boldsymbol{\gamma}}^{\star}=\left(1-\sqrt{5}-\sqrt{5-2\sqrt{5}},1-\sqrt{5}+\sqrt{5-2\sqrt{5}}\right)\approx(-1.9626,-0.5095),$
+and the optimal value $\tau^{\star}\approx-1.4531$.
+The computational results for $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ with different $\varepsilon$ and $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ are presented in Table 4.
+As expected, for a given $\varepsilon>0$, both $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ provide lower bounds for the optimal value $\tau^{\star}$ when $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ are sufficiently large (except for $\tau_{0.01}(2,15)$ and $\tilde{\tau}_{0.01}(2,15)$ which requires larger $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ as $\varepsilon=0.01$ is small).
+Moreover, for a fixed $\varepsilon>0$, both quantities are non-increasing as $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$ increase, and for fixed $k_{{\boldsymbol{x}}}$ and $k_{{\boldsymbol{y}}}$, they are non-decreasing as $\varepsilon$ decreases.
+Compared with $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, the values $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ with the same $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ are smaller and can be computed in much less time.
+To illustrate the behavior of optimal solutions of (13) and (16) under varying $\varepsilon$ and $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, like in Example 5.1, we solve (13) and (16) with the objective function $\gamma_{1}\cos\theta+\gamma_{2}\sin\theta$ for $40$ evenly spaced values of $\theta$ over $0,2\pi]$, and extract the corresponding optimal solutions.
+The feasible sets of (13) and (16), which serve as approximations of $\mathcal{F}$, are then outlined by connecting these optimal solutions and shown in Figure 2.
+As one can see, the behavior of these optimal points aligns with the quantities $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ and $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, confirming the theoretical expectations discussed above.
+In particular, it is evident from the figure that the feasible set of (16) contains that of (13) with the same $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$, which corresponds to the fact that $\tilde{\tau}_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$ is smaller than $\tau_{\varepsilon}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
+∎
+Table 4. Computational results for Example 5.3.
+$\varepsilon$ $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 5) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 10) $(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}}$)=(2, 15)
+$\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tau(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time $\tilde{\tau}(k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$/time
+$0.1$ -1.3716/2.20s -1.4922/1.69s -1.4844/32.3s -1.6587/7.32s -1.5650/547s -1.6837/27.2s
+$0.05$ -1.2664/2.34s -1.3163/1.93s -1.3698/31.1s -1.4899/7.21s -1.4552/507s -1.5230/29.9s
+$0.01$ -1.1873/2.57s -1.1806/2.17s -1.2949/33.3s -1.3625/7.86s -1.3721/451s -1.3900/35.7s
+Figure 2.
+The feasible set (in light blue) of the problem in Example 5.3 and the feasible sets of its SDP relaxations (13) (outlined with blue line) and (16) (outlined with red line) with the same values $(\varepsilon,k_{{\boldsymbol{x}}},k_{{\boldsymbol{y}}})$.
 6. Proofs
-6.1. Proof of Theorem 3.1
-Proof of Theorem 3.1.
-As mentioned earlier, the proof proceeds in four steps: Step 1: Let $\mathscr{L}$ be the linear functional in Corollary 3.1.
-The linear functional $\mathscr{L}$ induces a symmetric bilinear form on the space ${\mathbb{R}}[{\boldsymbol{x}}]^{p}$ by
-$(\psi,\phi)\mapsto\langle\psi,\phi\rangle_{\mathscr{L}}\coloneqq\frac{1}{2}\mathscr{L}(\psi\phi^{\intercal}+\phi\psi^{\intercal}),\quad\forall\psi,\phi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}.$
-This form is PSD since $\psi\psi^{\intercal}\in\mathcal{Q}^{p}(G,\nu)$ for any $\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}$. Let
-$\mathcal{N}\coloneqq\{\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}\mid\langle\psi,\psi\rangle_{\mathscr{L}}=0\}$
-be the $\langle\cdot,\cdot\rangle_{\mathscr{L}}$-null vector space corresponding to $\mathscr{L}$.
-By the Cauchy-Schwarz inequality, $\mathcal{N}$ is a vector subspace of ${\mathbb{R}}[{\boldsymbol{x}}]^{p}$ and $\mathscr{L}$ induces an inner product on the quotient space ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$.
-Denote by $\mathcal{H}$ the Hilbert space completion of ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$ and $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ its inner product.
-Recall that by the proof of Lemma 3.1, there exists some $C>0$ such that $C-x_{i}^{2}\in\mathcal{Q}^{1}(G,\nu)$ for all $i\in[n]$.
-It implies that $(C-x_{i}^{2})\psi\psi^{\intercal}\in\mathcal{Q}^{p}(G,\nu)$ for any $\psi\in{\mathbb{R}}[{\boldsymbol{x}}]^{p}$.
-Hence,
-$0\leq\langle x_{i}\psi+\mathcal{N},x_{i}\psi+\mathcal{N}\rangle_{\mathcal{H}}=\mathscr{L}(x_{i}^{2}\psi\psi^{\intercal})\leq C\mathscr{L}(\psi\psi^{\intercal}).$
-Therefore, the multiplication $\hat{x}_{i}:\psi\mapsto x_{i}\psi$ defines a bounded operator on ${\mathbb{R}}[{\boldsymbol{x}}]^{p}/\mathcal{N}$ and can be extended to a bounded operator on $\mathcal{H}$.
-Fix an orthonormal basis of $\mathcal{H}$ and let $\widehat{A}_{i}$ be the corresponding representative of the multiplication by $x_{i}$ in $\mathcal{B}(\mathcal{H})$ with respect to this basis.
-Let $\widehat{\mathbf{A}}\coloneqq(\widehat{A}_{i})_{i\in[n]}$.
-Denote by $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ the $p\times p$ matrix algebra over the $C^{\star}$-algebra $\mathcal{B}(\mathcal{H})$.
-That is, $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ consists of $p\times p$ matrices whose entries are in $\mathcal{B}(\mathcal{H})$.
-Then, $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ is a $C^{\star}$-algebra with involution $[B_{ij}]^{\star}=[B_{ji}^{\star}]$ where $B_{ij}\in\mathcal{B}(\mathcal{H})$.
-For $k\in[p]$, let ${\boldsymbol{v}}_{k}\in\mathcal{H}$ be the image of the constant polynomial vector $e_{k}\in{\mathbb{R}}^{p}$, where $\{e_{1},\ldots,e_{p}\}$ is the standard basis for ${\mathbb{R}}^{p}$.
-Define a linear functional $\Pi$ on $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$ by
-$\Pi(\mathbf{B})=\sum_{1\leq i,j\leq p}\langle B_{ij}{\boldsymbol{v}}_{j},{\boldsymbol{v}}_{i}\rangle_{\mathcal{H}},\ \ \forall\mathbf{B}=[B_{ij}]\in\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})).$
-Then, $\Pi$ is a state on $\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$.
-Note that for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, it holds that $\Pi(H(\widehat{\mathbf{A}}))=\mathscr{L}(H)$.
-Step 2: By applying the GNS construction to $(\mathcal{M}_{p}(\mathcal{B}(\mathcal{H})),\Pi)$, we obtain a Hilbert space $\mathcal{K}$ with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{K}}$, a representation $\pi:\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))\to\mathcal{B}(\mathcal{K})$ and a vector $\xi\in\mathcal{K}$ such that $\Pi(\mathbf{B})=\langle\pi(\mathbf{B})\xi,\xi\rangle_{\mathcal{K}}$ for all $\mathbf{B}\in\mathcal{M}_{p}(\mathcal{B}(\mathcal{H}))$.
-For each $i,j\in[p]$, denote by $E_{ij}$ the $p\times p$ matrix whose $(i,j)$-th entry is the identity operator and all other entries are the zero operator in $\mathcal{B}(\mathcal{H})$.
-For each $i\in[p]$, let $\mathcal{K}_{i}=\pi(E_{ii})(\mathcal{K})$.
-Then, the subspaces $\mathcal{K}_{i}$’s are all mutually orthogonal and isomorphic.
-In fact, $\pi(E_{ij})$ is a linear isomorphism between $\mathcal{K}_{j}$ and $\mathcal{K}_{i}$ with the inverse $\pi(E_{ji})$.
-Moreover,
-$\mathcal{K}=1_{\mathcal{B}(\mathcal{K})}\mathcal{K}=\pi\left(\sum_{i=1}^{p}E_{ii}\right)\mathcal{K}=\bigoplus_{i=1}^{p}\mathcal{K}_{i}.$
-Hence, each $\mathcal{K}_{i}$ is a Hilbert subspace of $\mathcal{K}$. Let us decompose $\xi$ as $\xi=\xi_{1}+\cdots+\xi_{p}$ where each $\xi_{i}\in\mathcal{K}_{i}$, i.e., $\xi_{i}=\pi(E_{ii})\xi$.
-For $H({\boldsymbol{x}})=[H_{ij}({\boldsymbol{x}})]\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, it holds that
-$\mathscr{L}(H)$ $=\Pi(H(\widehat{\mathbf{A}}))=\left\langle\pi\left(\sum_{i,j}H_{ij}(\widehat{\mathbf{A}})E_{ij}\right)\sum_{\ell=1}^{p}\xi_{\ell},\sum_{\ell=1}^{p}\xi_{\ell}\right\rangle_{\mathcal{K}}$ (9)
-$=\left\langle\sum_{i,j}\left(\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{ij})\right)\sum_{\ell=1}^{p}\xi_{\ell},\sum_{\ell=1}^{p}\xi_{\ell}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{ij})\xi_{j},\xi_{i}\right\rangle_{\mathcal{K}}.$
-For $i,j\in[p]$, let $\eta_{j}=\pi(E_{1i})\pi(E_{ij})\xi_{j}$ and $\eta_{i}=\pi(E_{1i})\xi_{i}$. Then, $\eta_{j},\eta_{i}\in\mathcal{K}_{1}$ and
-$\mathscr{L}(H)$ $=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\pi(E_{i1})\eta_{j},\pi(E_{i1})\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\pi(E_{1i})\pi(E_{i1})\eta_{i}\right\rangle_{\mathcal{K}}$
-$=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\pi(E_{11})\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}}.$
-For each $i\in[p]$, let $\mathbf{A}=(A_{1},\ldots,A_{n})$ with $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ being the restriction of $\pi(\widehat{A}_{i}I_{p})$ to $\mathcal{K}_{1}$ and denote by $\langle\cdot,\cdot\rangle_{\mathcal{K}_{1}}$ the inner product on $\mathcal{K}_{1}$ induced from $\langle\cdot,\cdot\rangle_{\mathcal{K}}$.
-Then, we have
-$\mathscr{L}(H)=\sum_{i,j}\left\langle\pi(H_{ij}(\widehat{\mathbf{A}})I_{p})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}}=\sum_{i,j}\left\langle H_{ij}(\mathbf{A})\eta_{j},\eta_{i}\right\rangle_{\mathcal{K}_{1}}.$ (10)
-Step 3: For each $i\in[n]$, $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ is self-adjoint. Indeed, for every $\omega,\gamma\in\mathcal{K}_{1}$,
-$\langle A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}=\langle\pi(\widehat{A}_{i}I_{p})\omega,\gamma\rangle_{\mathcal{K}}=\langle\omega,(\pi(\widehat{A}_{i}I_{p}))^{\star}\gamma\rangle_{\mathcal{K}}=\langle\omega,\pi((\widehat{A}_{i}I_{p})^{\star})\gamma\rangle_{\mathcal{K}}=\langle\omega,\pi(\widehat{A}_{i}I_{p})\gamma\rangle_{\mathcal{K}}=\langle\omega,A_{i}\gamma\rangle_{\mathcal{K}_{1}}.$
-Moreover, $(A_{i})_{i\in[n]}$ are pairwise commute. In fact, for every $\omega,\gamma\in\mathcal{K}_{1}$,
-$\langle A_{i}A_{j}\omega,\gamma\rangle_{\mathcal{K}_{1}}$ $=\langle\pi(\widehat{A}_{i}I_{p})\pi(\widehat{A}_{j}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle\pi(\widehat{A}_{i}\widehat{A}_{j}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}$
-$=\langle\pi(\widehat{A}_{j}\widehat{A}_{i}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle\pi(\widehat{A}_{j}I_{p})\pi(\widehat{A}_{i}I_{p})\omega,\gamma)\rangle_{\mathcal{K}}=\langle A_{j}A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}.$
-For each $i\in[n]$, as $A_{i}\in\mathcal{B}(\mathcal{K}_{1})$ is self-adjoint, by the spectral theorem for bounded self-adjoint operators on a Hilbert space (see, e.g.
-[48, Theorem 5.1]), there exists a unique projection-valued measure $\mathcal{E}_{i}$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}})$ such that the support of $\mathcal{E}_{i}$ is contained in a compact interval, $A_{i}=\int_{{\mathbb{R}}}t\mathrm{d}\mathcal{E}_{i}(t)$ and
-$\langle A_{i}\omega,\gamma\rangle_{\mathcal{K}_{1}}=\int_{{\mathbb{R}}}t\mathrm{d}\langle\mathcal{E}_{i}(t)\omega,\gamma\rangle_{\mathcal{K}_{1}},\quad\forall\omega,\gamma\in\mathcal{K}_{1}.$
-Moreover, as $(A_{i})_{i\in[n]}$ are pairwise commute, by [48, Theorem 4.10], we can define on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ the product measure $\mathcal{E}$ of $\mathcal{E}_{i},i\in[n]$, such that $\mathcal{E}\left(\prod_{i\in[n]}\mathcal{N}_{i}\right)=\prod_{i\in[n]}\mathcal{E}_{i}(\mathcal{N}_{i})$ for all $\mathcal{N}_{i}\in\mathfrak{B}({\mathbb{R}}),\ i\in[n].$ By the multidimensional spectral theorem (see, [48, Theorem 5.23] or [4, Sec.
-6.5, Theorem 1]), $A_{i}=\int_{{\mathbb{R}}^{n}}x_{i}\mathrm{d}\mathcal{E}({\boldsymbol{x}})$, $i\in[n]$, and
-$\left\langle\left(\prod_{i\in[n]}A_{i}^{\alpha_{i}}\right)\omega,\gamma\right\rangle_{\mathcal{K}_{1}}=\int_{{\mathbb{R}}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\langle\mathcal{E}({\boldsymbol{x}})\omega,\gamma\rangle_{\mathcal{K}_{1}},\quad\forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n},\ \omega,\gamma\in\mathcal{K}_{1}.$
-Note that the support of the product measure $\mathcal{E}$ is compact since each $\mathcal{E}_{i}$ has compact support.
-For $i,j\in[p]$, recall that the vectors $\eta_{i},\eta_{j}\in\mathcal{K}_{1}$.
-Define the real-valued measure $\phi_{ij}(\cdot)\coloneqq\langle\mathcal{E}(\cdot)\eta_{j},\eta_{i}\rangle_{\mathcal{K}_{1}}$ on the Borel $\sigma$-algebra $\mathfrak{B}({\mathbb{R}}^{n})$ and the symmetric matrix-valued measure $\Phi\coloneqq[\phi_{ij}]$.
-Then, each $\phi_{ij}$ has a compact support and for any $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, we have $\mathscr{L}(H)=\int_{{\mathbb{R}}^{n}}H({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})$ by (10).
-Moreover, $\Phi$ is a PSD matrix-valued measure.
-In fact, for any $\mathcal{N}\in\mathfrak{B}({\mathbb{R}}^{n})$ and ${\boldsymbol{v}}=(v_{i})\in{\mathbb{R}}^{p}$, as $\mathcal{E}(\mathcal{N})$ is a projection operator on $\mathcal{K}_{1}$, we have
-${\boldsymbol{v}}^{\intercal}\Phi(\mathcal{N}){\boldsymbol{v}}=\sum_{i,j}\phi_{ij}(\mathcal{N})v_{i}v_{j}=\sum_{i,j}\langle\mathcal{E}(\mathcal{N})v_{j}\eta_{j},v_{i}\eta_{i}\rangle_{\mathcal{K}_{1}}=\left\langle\mathcal{E}(\mathcal{N})\sum_{i}v_{i}\eta_{i},\mathcal{E}(\mathcal{N})\sum_{j}v_{j}\eta_{j}\right\rangle_{\mathcal{K}_{1}}\geq 0.$
-Step 4: Now we show that for $i,j\in[p]$, the support of $\phi_{ij}$ is contained in $\mathcal{X}$.
-Suppose that ${\boldsymbol{u}}\in{\mathbb{R}}^{n}$ and ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$.
-Then, there is a ball $\mathcal{O}_{1}\subset{\mathbb{R}}^{n}$ (resp., $\mathcal{O}_{2}\subset{\mathbb{R}}^{m}$) with a radius $\rho>0$ around ${\boldsymbol{u}}$ (resp., ${\boldsymbol{w}}$) such that $G({\boldsymbol{x}},{\boldsymbol{y}})\not\succeq 0$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$.
-Without loss of genericity, we may assume that there exists ${\boldsymbol{v}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}\leq-\delta$ on $2\mathcal{O}_{1}\times 2\mathcal{O}_{2}$.
-Define a continuous function $h^{(1)}({\boldsymbol{x}})$ (resp.
-$h^{(2)}({\boldsymbol{y}})$) on ${\mathbb{R}}^{n}$ (resp., ${\mathbb{R}}^{m}$) by $h^{(1)}({\boldsymbol{x}})=2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|$ (resp., $h^{(2)}({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$)for ${\boldsymbol{x}}\in 2\mathcal{O}_{1}$ (resp., ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$) and $h^{(1)}({\boldsymbol{x}})=0$ (resp., $h^{(2)}({\boldsymbol{y}})=0$) otherwise.
-Take a compact set $\mathcal{C}\subset{\mathbb{R}}^{n}$ containing $2\mathcal{O}_{1}$ and the support of $\phi_{ii}$, $i\in[p]$.
-By the Weierstrass theorem, there is a sequence of square polynomials $\{h^{2}_{\ell}\}_{\ell\in\mathbb{N}}\subset{\mathbb{R}}[{\boldsymbol{x}}]$ which converges to $h^{(1)}$ uniformly on $\mathcal{C}$.
-By Proposition 3.1, there exists a sequence of SOS polynomilas $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h^{(2)}$ in the $L^{2}$-norm.
-Hence,
-$\lim_{\ell\to\infty}\lim_{k\to\infty}\sum_{i=1}^{p}\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$ (11)
-$=$ $\lim_{\ell\to\infty}\lim_{k\to\infty}\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
-$=$ $\lim_{\ell\to\infty}\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})h^{(2)}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
-$=$ $\sum_{i=1}^{p}\int_{\mathcal{C}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h^{(1)}({\boldsymbol{x}})h^{(2)}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
-$=$ $\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}\left(\int_{\mathcal{Y}\cap 2\mathcal{O}_{2}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
-$\leq$ $\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}\left(\int_{\mathcal{Y}\cap\mathcal{O}_{2}}-\delta\rho(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
-$=$ $-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\int_{2\mathcal{O}_{1}}(2\rho-\|{\boldsymbol{x}}-{\boldsymbol{u}}\|)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$
-$\leq$ $-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\int_{\mathcal{O}_{1}}\rho\mathrm{d}\phi_{ii}({\boldsymbol{x}})=-\delta\rho^{2}\nu(\mathcal{Y}\cap\mathcal{O}_{2})\sum_{i=1}^{p}\phi_{ii}(\mathcal{O}_{1}).$
-As $\mathcal{C}$ is compact, the second and third equalities in (11) can be verified by Dominated Convergence Theorem and the convergence of $\{\sigma_{k}\}_{k}$ to $h^{(2)}$ in the $L^{2}$-norm.
-Note that for all $\ell,k\in\mathbb{N}$,
-$\sum_{i=1}^{p}\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}{\boldsymbol{v}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{v}}h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi_{ii}({\boldsymbol{x}})$ (12)
-$=$ $\int_{{\mathbb{R}}^{n}}\left(\int_{\mathcal{Y}}\left\langle h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\Phi({\boldsymbol{x}})$
-$=$ $\mathscr{L}\left(\int_{\mathcal{Y}}\left\langle h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
-where the last inequality is due to the fact that $h_{\ell}^{2}({\boldsymbol{x}})\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}$ is an SOS and $\mathscr{L}(\mathcal{Q}^{p}(G,\nu))\subseteq{\mathbb{R}}_{+}$.
-Since $\mathrm{supp}(\nu)=\mathcal{Y}$, we have $\nu(\mathcal{Y}\cap\mathcal{O}_{2})>0$.
-Therefore, by combining (11) and (12), we have $\phi_{ii}(\mathcal{O}_{1})=0$ for all $i\in[p]$.
-As $\Phi$ is PSD, it implies that $\phi_{ij}(\mathcal{O}_{1})=0$ for all $i,j\in[p]$.
-Hence, the support of each $\phi_{ij}$ is contained in $\mathcal{X}$.
-Then, as $F$ is positive on $\mathcal{X}$ by assumption and $\Phi$ is PSD, we have
-$0\geq\mathscr{L}(F)=\int_{\mathcal{X}}F({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})>0,$
-a contradiction. ∎
 6.1. Proof of Proposition 3.4
@@
 For each $j\in[s]$, let
-$M\coloneqq\max\left\{|\lambda_{\max}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|\colon{\boldsymbol{x}}\in[-C,C]^{n},\ {\boldsymbol{y}}\in\mathcal{Y},\ j\in[s]\right\}.$
-As $\mathcal{Y}$ is compact, the quantity $M$ is well-defined.
+$M\coloneqq\max\left\{|\lambda_{\max}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|,|\lambda_{\min}(G_{j}({\boldsymbol{x}},{\boldsymbol{y}}))|\colon{\boldsymbol{x}}\in[-C,C]^{n},\ {\boldsymbol{y}}\in\mathcal{Y},\ j\in[s]\right\}.$
+As $\mathcal{Y}$ is compact, the quantity $M$ is well-defined and positive.
 For this $M$, we show that there exists some $\bar{k}\in\mathbb{N}$ such that (5) holds for all $k\geq\bar{k}$.
@@
 $F({\boldsymbol{x}}^{(k)})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}}^{(k)},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}}^{(k)},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\not\succ 0.$ (17)
-As $[-C,C]^{n}$ is compact, without loss of generality, we may assume that $\lim_{k\to\infty}{\boldsymbol{x}}^{(k)}={\boldsymbol{x}}^{\star}$ for some ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}$.
-Next, we prove that there exists $k^{\prime}\in\mathbb{N}$ and a neighborhood $\mathcal{O}_{1}$ of ${\boldsymbol{x}}^{\star}$ such that (4) holds for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$, which yields a contradiction.
-We first consider the case that ${\boldsymbol{x}}^{\star}\in\widehat{\mathcal{X}}$.
+As $C,C]^{n}$ is compact, without loss of generality, we may assume that $\lim_{k\to\infty}{\boldsymbol{x}}^{(k)}={\boldsymbol{x}}^{\star}$ for some ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}$.
+Next, we claim that there exists $k^{\prime}\in\mathbb{N}$ and a neighborhood $\mathcal{O}_{1}$ of $\boldsymbol{x}$$^{\star}$ such that (5) holds for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$, which yields a contradiction.
+We first consider the case that $^{\boldsymbol{x}}^{\star}\in\widehat{\mathcal{X}}$.
 As $F({\boldsymbol{x}})\succ 0$ on $\widehat{\mathcal{X}}$, there exists $\varepsilon>0$ and a neighborhood $\mathcal{O}_{1}$ of $\boldsymbol{x}$$^{\star}$ such that $F({\boldsymbol{x}})\succeq\varepsilon I_{p}$ for all ${\boldsymbol{x}}\in\mathcal{O}_{1}$.
@@
 Then,
-[…]
+$\begin{aligned} \left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{u}},{\boldsymbol{w}})/M\right)^{2k},G_{j}({\boldsymbol{u}},{\boldsymbol{w}})\right\rangle=&\left\langle\left(I_{q_{j}}-Q_{j}(D_{j}/M)Q_{j}^{\intercal}\right)^{2k},Q_{j}D_{j}Q_{j}^{\intercal}\right\rangle\\ =&\left\langle Q_{j}\left(I_{q_{j}}-D_{j}/M\right)^{2k}Q_{j}^{\intercal},Q_{j}D_{j}Q_{j}^{\intercal}\right\rangle=M\sum_{i=1}^{q_{j}}\frac{\lambda_{i}^{(j)}}{M}\left(1-\frac{\lambda_{i}^{(j)}}{M}\right)^{2k}\end{aligned}.$ (18)
 Let $q^{\prime}\coloneqq\max_{j\in[s]}q_{j}$. Note that there exists $k^{\prime}\in\mathbb{N}$ such that for all $k\geq k^{\prime}$,
@@
 Hence, for any pair of ${\boldsymbol{x}}\in[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, from (18) we obtain that
-$\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\leq M\sum_{\lambda_{i}^{(j)}>0}\frac{\lambda_{i}^{(j)}}{M}\left(1-\frac{\lambda_{i}^{(j)}}{M}\right)^{2k}\leq Mq_{j}\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}\leq\frac{\varepsilon}{2s\nu(\mathcal{Y})}$ (15)
+$\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle\leq Mq_{j}\frac{\varepsilon}{2q^{\prime}sM\nu(\mathcal{Y})}\leq\frac{\varepsilon}{2s\nu(\mathcal{Y})}$ (19)
 for all $k\geq k^{\prime}$, where $k^{\prime}$ does not depends on the choice of ${\boldsymbol{x}}\in[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{Y}$.
@@
 $F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succeq\varepsilon I_{p}-\frac{\varepsilon}{2}I_{p}=\frac{\varepsilon}{2}I_{p}\succ 0.$
-Next we consider the case that ${\boldsymbol{x}}^{\star}\in[-C,C]^{n}\setminus\widehat{\mathcal{X}}$.
+Next, we consider the case that $^{\boldsymbol{x}}^{\star}\in[-C,C]^{n}\setminus\widehat{\mathcal{X}}$.
 There exists $j_{0}\in[s]$ and ${\boldsymbol{y}}^{(0)}\in\mathcal{Y}$ such that $\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}}^{\star},{\boldsymbol{y}}^{(0)})<0$.
@@
 For any pair of ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and ${\boldsymbol{y}}\in\mathcal{O}_{2}\cap\mathcal{Y}$, letting $\lambda_{\min}^{(j_{0})}\coloneqq\lambda_{\min}(G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}}))$, from (18) we obtain that
-$\left\langle\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle$ $\leq\sum_{\lambda_{i}^{(j_{0})}>0}\lambda_{i}^{(j_{0})}+\sum_{\lambda_{i}^{(j_{0})}<0}\lambda_{i}^{(j_{0})}\left(1-\frac{\lambda_{i}^{(j_{0})}}{M}\right)^{2k}$
+$\left\langle\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle$ $\leq\sum_{\lambda_{i}^{(j_{0})}>0}\lambda_{i}^{(j_{0})}+\sum_{\lambda_{i}^{(j_{0})}<0}\lambda_{i}^{(j_{0})}\left(1-\frac{\lambda_{i}^{(j_{0})}}{M}\right)^{2k}$ (20)
 $\leq Mq^{\prime}+\lambda_{\min}^{(j_{0})}\left(1-\frac{\lambda_{\min}^{(j_{0})}}{M}\right)^{2k}\leq Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}.$
@@
 $\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}=\left\langle\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle I_{p}$
-could be uniformly bounded from above on $[-C,C]^{n}\times\mathcal{Y}$ for all $k\in\mathbb{N}$. Therefore, as $[-C,C]^{n}$ and $\mathcal{Y}$ are compact, there exists $R\in{\mathbb{R}}$ such that
-$F({\boldsymbol{x}})-\sum_{j\neq j_{0}}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
+could be uniformly bounded from above on $C,C]^{n}\times\mathcal{Y}$ for all $k\in\mathbb{N}$. Therefore, as $C,C]^{n}$ and $\mathcal{Y}$ are compact, there exists $R\in{\mathbb{R}}$ such that
+$F({\boldsymbol{x}})-\sum_{j\neq j_{0}}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$ (21)
 $-\int_{\mathcal{Y}\setminus\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})-(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1)I_{p}\succeq RI_{p}$
-for all ${\boldsymbol{x}}\in[-C,C]^{n}$.
-Since $11-\bar{\lambda}/M>$ 1, there exists $k^{\prime}\in\mathbb{N}$ such that $\bar{\lambda}\left(1-\bar{\lambda}/{M}\right)^{2k}<R/\nu(\mathcal{Y}\cap\mathcal{O}_{2})$ holds for all $k\geq k^{\prime}$.
-Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
+for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Since $1-\bar{\lambda}/M>1$, there exists $k^{\prime}\in\mathbb{N}$ such that
+$\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}<\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}$ (22)
+holds for all $k\geq k^{\prime}$. Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
 $F({\boldsymbol{x}})-\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
-$\succeq$ $-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R)I_{p}$
-$\succeq$ $\left(-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left(Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}\right)\mathrm{d}\nu({\boldsymbol{y}})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}$
-$\succeq$ $\left(-\left(Mq^{\prime}+\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}\right)\nu(\mathcal{Y}\cap\mathcal{O}_{2})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}=I_{p},$
-which completes the proof. ∎
+$\succeq$ $-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R)I_{p}$ (by (21))
+$\succeq$ $\left(-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left(Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}\right)\mathrm{d}\nu({\boldsymbol{y}})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}$ (by (20))
+$\succeq$ $\left(-\left(Mq^{\prime}+\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}\right)\nu(\mathcal{Y}\cap\mathcal{O}_{2})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}=I_{p},$ (by (22))
+which validates the claim and completes the proof. ∎
 Using Proposition 3.4, we can recover Theorem 3.1 in the case that $\mathcal{Y}$ is compact.
@@
 Proof.
-Write $h=\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\langle\Sigma_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})$ where all $\sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$ are SOS.
-Then,
-$Sh=S\sigma_{0}+\sum_{j=1}^{s}S\int_{\mathcal{Y}}\langle\Sigma_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})=S\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S\otimes\Sigma_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-As $S\sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $S\otimes\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq_{j}}$ are all SOS matrices, we have $Sh\in\mathcal{Q}^{p}(\mathbf{G},\nu)$.
+Write $h=\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\langle S_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})$ where all $\sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$ are SOS.
+Then
+$Sh=S\sigma_{0}+\sum_{j=1}^{s}S\int_{\mathcal{Y}}\langle S_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})=S\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S\otimes S_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+As $S\sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $S\otimes S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq_{j}}$ are all SOS matrices, we have $Sh\in\mathcal{Q}^{p}(\mathbf{G},\nu)$.
 ∎
@@
 Then $F({\boldsymbol{u}})\succ 0$ for all ${\boldsymbol{u}}\in\widetilde{\mathcal{X}}$. By Corollary 3.1, it holds that
-$F({\boldsymbol{x}})=\Sigma^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+H({\boldsymbol{x}})(1-\|{\boldsymbol{x}}\|^{2}),$
-[…]
-$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{-\deg{F}}=\Sigma^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)+\int_{\mathcal{Y}}\left\langle\Sigma^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right),G\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
+$F({\boldsymbol{x}})=S^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+H({\boldsymbol{x}})(1-\|{\boldsymbol{x}}\|^{2}),$
+where $^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ are SOS matrices and $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$.
+Replacing $\boldsymbol{x}$} by ${\boldsymbol{x}}/\|{\boldsymbol{x}}\|$ in the above equality yields
+$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{-\deg{F}}=S^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)+\int_{\mathcal{Y}}\left\langle S^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right),G\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
 Let
-$k^{\prime}\coloneqq\max\left\{\deg{F},\ \deg(\Sigma^{\prime}_{0}),\ \deg_{{\boldsymbol{x}}}{G}+\deg_{{\boldsymbol{x}}}\Sigma^{\prime}\right\}.$
+$k^{\prime}\coloneqq\max\left\{\deg{F},\ \deg S^{\prime}_{0},\ \deg_{{\boldsymbol{x}}}{G}+\deg_{{\boldsymbol{x}}}S^{\prime}\right\}.$
 By assumption, $k^{\prime}$ is even. Multiplying the two sides of the last equality with $\|{\boldsymbol{x}}\|^{k^{\prime}}$ gives
-$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=\overline{\Sigma}_{0}\left({\boldsymbol{x}}\right)+\int_{\mathcal{Y}}\left\langle\overline{\Sigma}\left({\boldsymbol{x}},{\boldsymbol{y}}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (16)
+$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=\overline{S}_{0}\left({\boldsymbol{x}}\right)+\int_{\mathcal{Y}}\left\langle\overline{S}\left({\boldsymbol{x}},{\boldsymbol{y}}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (23)
 where
-$\overline{\Sigma}_{0}\coloneqq\Sigma^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)\|{\boldsymbol{x}}\|^{k^{\prime}}\ \ \text{and}\ \ \overline{\Sigma}\coloneqq\Sigma^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\|{\boldsymbol{x}}\|^{k^{\prime}-\deg_{{\boldsymbol{x}}}{G}}.$
-Since $\Sigma^{\prime}$ is an SOS matrix and $k^{\prime}-\deg_{{\boldsymbol{x}}}{G}\geq\deg_{{\boldsymbol{x}}}\Sigma^{\prime}$, we have $\overline{\Sigma}=H^{\intercal}H$ with $H=H_{1}+H_{2}\|{\boldsymbol{x}}\|$ where $H_{1},H_{2}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times m}$ for some $\ell\in\mathbb{N}$, are homogeneous in ${\boldsymbol{x}}$ of degree $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2$ and $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2-1$, respectively.
+$\overline{S}_{0}\coloneqq S^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)\|{\boldsymbol{x}}\|^{k^{\prime}}\ \ \text{and}\ \ \overline{S}\coloneqq S^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\|{\boldsymbol{x}}\|^{k^{\prime}-\deg_{{\boldsymbol{x}}}{G}}.$
+Since $S^{\prime}$ is an SOS matrix and $k^{\prime}-\deg_{{\boldsymbol{x}}}{G}\geq\deg_{{\boldsymbol{x}}}S^{\prime}$, we have $\overline{S}=H^{\intercal}H$ with $H=H_{1}+H_{2}\|{\boldsymbol{x}}\|$ where $H_{1},H_{2}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell\in\mathbb{N}$, are homogeneous in $\boldsymbol{x}$} of degree $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2$ and $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2-1$, respectively.
 Thus,
-$\overline{\Sigma}=H^{\intercal}H=(H_{1}+H_{2}\|{\boldsymbol{x}}\|)^{\intercal}(H_{1}+H_{2}\|{\boldsymbol{x}}\|)=\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right)\|{\boldsymbol{x}}\|.$
+$\overline{S}=H^{\intercal}H=(H_{1}+H_{2}\|{\boldsymbol{x}}\|)^{\intercal}(H_{1}+H_{2}\|{\boldsymbol{x}}\|)=\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right)\|{\boldsymbol{x}}\|.$
 Similarly, there exist homogeneous polynomial matrices $H_{0,1}\in{\mathbb{R}}[x]^{\ell_{0}\times p}_{k^{\prime}/2}$ and $H_{0,2}\in{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}_{k^{\prime}/2-1}$ for some $\ell_{0}\in\mathbb{N}$ such that
-$\overline{\Sigma}_{0}=\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)\|{\boldsymbol{x}}\|.$
+$\overline{S}_{0}=\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)\|{\boldsymbol{x}}\|.$
 Then, by (23), it holds that
@@
 $\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)+\int_{\mathcal{Y}}\left\langle\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})=0.$
-Then, letting $N\coloneqq(k^{\prime}-\deg{F})/2\in\mathbb{N}$, $\Sigma_{0}\coloneqq H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}$ and $\Sigma\coloneqq H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}$, we obtain the desired conclusion.
+Then, letting $N\coloneqq(k^{\prime}-\deg{F})/2\in\mathbb{N}$, $S_{0}\coloneqq H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}$ and $S\coloneqq H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}$, we obtain the desired conclusion.
 ∎
@@
 Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$.
-Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ ($\deg(S_{0})=N+\deg F$ and $\deg_{{\boldsymbol{x}}}S=N+\deg F-\deg_{{\boldsymbol{x}}}G$), and satisfy that $S_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$ (resp., $S({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$) are SOS matrices in ${\boldsymbol{z}}$ (resp., ${\boldsymbol{z}}$ and ${\boldsymbol{y}}$), such that
+Then, there exists $N\in\mathbb{N}$ and polynomial matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} ($\deg S_{0}=N+\deg F$ and $\deg_{{\boldsymbol{x}}}S=N+\deg F-\deg_{{\boldsymbol{x}}}G$), and satisfy that $S_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$ (resp., $S({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$) are SOS matrices in $\boldsymbol{z}$} (resp., $\boldsymbol{z}$} and $\boldsymbol{y}$}), such that
 $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
 Consider the polynomial matrices $F({\boldsymbol{z}}\circ{\boldsymbol{z}})\in{\mathbb{S}}[{\boldsymbol{z}}]^{p}$ and $G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{q}$.
-Then, by assumption and Theorem 4.1, there exists $N$ and polynomial matrices $H_{0}\in{\mathbb{R}}[{\boldsymbol{z}}]^{\ell_{0}\times p}$, $H\in{\mathbb{R}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell_{0},\ell\in\mathbb{N}$ which are homogeneous in ${\boldsymbol{z}}$, such that
+Then, by assumption and Theorem 4.1, there exists $N$ and polynomial matrices $H_{0}\in{\mathbb{R}}[{\boldsymbol{z}}]^{\ell_{0}\times p}$, $H\in{\mathbb{R}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell_{0},\ell\in\mathbb{N}$ that are homogeneous in $\boldsymbol{z}$}, such that
 $\|{\boldsymbol{z}}\|^{2N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})=(H_{0}^{\intercal}H_{0})({\boldsymbol{z}})+\int_{\mathcal{Y}}\left\langle(H^{\intercal}H)({\boldsymbol{z}},{\boldsymbol{y}}),G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
-Moreover, by Theorem 4.1, $\deg(H_{0})=N+\deg F$ and $\deg_{{\boldsymbol{x}}}(H)=N+\deg F-\deg_{{\boldsymbol{x}}}G$. Note that there are sets of polynomial matrices
-$\{H_{0,{\boldsymbol{\alpha}}}:{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}\quad\text{and}\quad\{H_{{\boldsymbol{\alpha}}}:{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times p}$
+Moreover, by Theorem 4.1, $\deg H_{0}=N+\deg F$, and $\deg_{{\boldsymbol{x}}}H=N+\deg F-\deg_{{\boldsymbol{x}}}G$. Note that there are sets of polynomial matrices
+$\{H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}}):{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}\quad\text{and}\quad\{H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}):{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times p}$
 which are homogeneous in $\boldsymbol{x}$}, such that
 $H_{0}({\boldsymbol{z}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{z}}\circ{\boldsymbol{z}}),\quad H({\boldsymbol{z}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{z}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}}).$
-Then,
+Then
 $({\boldsymbol{e}}^{\intercal}({\boldsymbol{z}}\circ{\boldsymbol{z}}))^{N}F({\boldsymbol{z}}\circ{\boldsymbol{z}})$
@@
 It follows that the equality
-$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
-holds for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}$.
-As +${\mathbb{R}}^{n}_{+}$ has interior points, the polynomial matrices on the left and right side of the above equality are identical.
-Letting $S_{0}\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})$ and $S\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})$, the conclusion follows.
-∎
+$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}})H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}},{\boldsymbol{y}})H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
+holds for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}$. As ${\mathbb{R}}^{n}_{+}$ has interior points, the polynomial matrices on the left and right side of the above equality are identical. Letting
+$S_{0}({\boldsymbol{x}})\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}})H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}})\text{ and }S({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}},{\boldsymbol{y}})H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}),$
+the conclusion follows. ∎
 As an extension of Pólya’s result [35], Scherer and Hol [42] provided the following certificate for homogeneous polynomial matrices being positive on the nonnegative orthant.
 Theorem 6.2.
-[45, Theorem 3] Suppose that the polynomial matrix $P({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is homogeneous and $P({\boldsymbol{x}})\succeq\lambda I_{p}$ for some $\lambda>0$ on $\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$.
-Write $P({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(P)}P_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ with each $P_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ and let
-$L(P)\coloneqq\max_{{\boldsymbol{\alpha}}\in\mathrm{supp}(P)}\frac{{\boldsymbol{\alpha}}!}{\deg(P)!}\|P_{{\boldsymbol{\alpha}}}\|,$
-[…]
-$N\geq\frac{\deg(P)(\deg(P)-1)L(P)}{2\lambda}-\deg(P),$
-all coefficients of $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}P({\boldsymbol{x}})$ are PD.
-We will use this theorem as another intermediary result. Let $D\coloneqq\max\left\{\deg(F),\deg_{{\boldsymbol{x}}}(G)\right\}$.
+[42, Theorem 3] Suppose that the polynomial matrix $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is homogeneous and $H({\boldsymbol{x}})\succeq\lambda I_{p}$ for some $\lambda>0$ on $\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$.
+Write $H({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}H_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ with each $H_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ and let
+$L(H)\coloneqq\max_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}\frac{{\boldsymbol{\alpha}}!}{\deg H!}\|H_{{\boldsymbol{\alpha}}}\|,$
+where $\|\cdot\|$ denotes the spectral norm. Then, for all
+$N\geq\frac{\deg H(\deg(H)-1)L(H)}{2\lambda}-\deg H,$
+the coefficient of ${\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ in $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}H({\boldsymbol{x}})$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{N+\deg H}$ is PD.
+We will use this theorem as another intermediary result. Let $D\coloneqq\max\left\{\deg F,\deg_{{\boldsymbol{x}}}G\right\}$.
 Lemma 6.2.
@@
 Then, there exists $\varepsilon>0$ such that the homogeneous polynomial matrix
-$F_{\varepsilon}({\boldsymbol{x}})\coloneqq({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg(F)}F({\boldsymbol{x}})-\varepsilon\left(({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D}I_{p}+\int_{\mathcal{Y}}\left\langle({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G}I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$
+$F_{\varepsilon}({\boldsymbol{x}})\coloneqq({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg F}F({\boldsymbol{x}})-\varepsilon\left(({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D}I_{p}+\int_{\mathcal{Y}}\left\langle({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G}I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$
 for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$.
@@
 By homogeneity, we only need to prove $F_{\varepsilon}\succ 0$ on $\overline{\mathcal{X}}\coloneqq{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}.$ That is, there exists $\varepsilon>0$ such that $F({\boldsymbol{x}})-\varepsilon\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$ for all ${\boldsymbol{x}}\in\overline{\mathcal{X}}$.
-Since $\overline{\mathcal{X}}$ is compact and $F({\boldsymbol{x}})\succ 0$ on +${\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, letting
+Since $\overline{\mathcal{X}}$ is compact and $F({\boldsymbol{x}})\succ 0$ on ${\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, letting
 $\varepsilon\coloneqq\frac{\min_{{\boldsymbol{x}}\in\overline{\mathcal{X}}}\lambda_{\min}(F({\boldsymbol{x}}))}{2\max_{{\boldsymbol{x}}\in\overline{\mathcal{X}}}|\lambda_{\max}\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)|},$ (24)
-the conclusion follows. ∎
-We now prove the following Pólya-type Positivstellensatz for polynomial matrices with UQs.
+we obtain the desired conclusion. ∎
+We now prove Theorem 4.2.
 Proof of Theorem 4.2.
 Let $F_{\varepsilon}({\boldsymbol{x}})$ be the polynomial matrix in Lemma 6.2 where $\varepsilon>0$ is defined in (24).
-Then, applying Theorem 6.1 to $F_{\varepsilon}({\boldsymbol{x}})$ yields $N_{1}\in\mathbb{N}$ and polynomial matrices $S^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$’s are SOS matrices, such that
+Then, applying Theorem 6.1 to $F_{\varepsilon}({\boldsymbol{x}})$ yields $N_{1}\in\mathbb{N}$ and polynomial matrices $^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} and satisfy that $S^{\prime}_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$, $S^{\prime}({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$’s are SOS matrices, such that
 $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{1}}F_{\varepsilon}({\boldsymbol{x}})=S^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
 Write $\Gamma({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ and define
-$L(\Gamma)\coloneqq\max_{{\boldsymbol{y}}\in\mathcal{Y}}\max_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\frac{{\boldsymbol{\alpha}}!}{\deg_{{\boldsymbol{x}}}(\Gamma)!}\|\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\|.$
+$L(\Gamma)\coloneqq\max_{{\boldsymbol{y}}\in\mathcal{Y}}\max_{{\boldsymbol{\alpha}}\in\mbox{supp}_{{\boldsymbol{x}}}(\Gamma)}\frac{{\boldsymbol{\alpha}}!}{\deg_{{\boldsymbol{x}}}\Gamma!}\|\Gamma_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\|.$
 As $\mathcal{Y}$ is compact, $L(\Gamma)$ is well-defined. Let
-$N_{2}\coloneqq\max\left\{\frac{\deg(\Gamma_{0})(\deg(\Gamma_{0})-1)L(\Gamma_{0})}{2\varepsilon}-\deg(\Gamma_{0}),\ \frac{\deg_{{\boldsymbol{x}}}(\Gamma)(\deg_{{\boldsymbol{x}}}(\Gamma)-1)L(\Gamma)}{2\varepsilon}-\deg_{{\boldsymbol{x}}}(\Gamma)\right\}.$
-Let $S_{0}({\boldsymbol{x}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma_{0}$ and $S({\boldsymbol{x}},{\boldsymbol{y}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma$.
+$N_{2}\coloneqq\max\left\{\frac{\deg\Gamma_{0}(\deg\Gamma_{0}-1)L(\Gamma_{0})}{2\varepsilon}-\deg\Gamma_{0},\ \frac{\deg_{{\boldsymbol{x}}}\Gamma(\deg_{{\boldsymbol{x}}}\Gamma-1)L(\Gamma)}{2\varepsilon}-\deg_{{\boldsymbol{x}}}\Gamma\right\}.$
+Let $P_{0}({\boldsymbol{x}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma_{0}$ and $P({\boldsymbol{x}},{\boldsymbol{y}})=({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N_{2}}\Gamma$.
 Write
-$S_{0}({\boldsymbol{x}})=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\quad\text{ and }\quad S({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}.$
-By Theorem 6.2, $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$ and ${\boldsymbol{y}}\in\mathcal{Y}$.
+$P_{0}({\boldsymbol{x}})=\sum_{|{\boldsymbol{\alpha}}|=\deg P_{0}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\quad\text{ and }\quad P({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}P}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}.$
+By Theorem 6.2, $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg P_{0}$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}P$ and ${\boldsymbol{y}}\in\mathcal{Y}$.
 Letting $N=D-\deg F+N_{1}+N_{2}$, the conclusion follows from (25).
 ∎
-6.5. Proof of Theorem 4.3
-To proof Theorem 4.3, we need some intermediary lemmas.
-Lemma 6.3.
-Assume that $M(\mathbf{S})\succeq 0$ and there are constants $C_{0},C>0$ such that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-Then, $\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{\mathbf{0}}\|C^{|{\boldsymbol{\alpha}}|}\leq\mbox{tr}\left(S_{\mathbf{0}}\right)C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-Proof.
-As $S_{\mathbf{0}}\succeq 0$, we only need to prove the first inequality.
-If $\|S_{\mathbf{0}}\|=0$, then $S_{\mathbf{0}}=0$ and hence $S_{{\boldsymbol{\alpha}}}=0$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$ since the null space of $S_{\mathbf{0}}$ contains that of $S_{{\boldsymbol{\alpha}}}$ by [31, Lemma 1.2 (i)].
-As $M(\mathbf{S})\succeq 0$, by [17, Theorem 7.7.11], there is a contraction $N_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$ such that $S_{{\boldsymbol{\alpha}}}=S_{\mathbf{0}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\alpha}}}^{1/2}$.
-Then, it holds
-$\|S_{{\boldsymbol{\alpha}}}\|=\|S_{\mathbf{0}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\alpha}}}^{1/2}\|\leq\|S_{\mathbf{0}}^{1/2}\|\|N_{{\boldsymbol{\alpha}}}\|\|S_{2{\boldsymbol{\alpha}}}^{1/2}\|\leq\|S_{\mathbf{0}}\|^{1/2}\|S_{2{\boldsymbol{\alpha}}}\|^{1/2}.$
-By induction, we have
-$\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{\mathbf{0}}\|^{\sum_{i=1}^{k}1/2^{i}}\|S_{2^{k}{\boldsymbol{\alpha}}}\|^{1/2^{k}}\leq\|S_{\mathbf{0}}\|^{1-1/2^{k}}(C_{0}C^{2^{k}|{\boldsymbol{\alpha}}|})^{1/2^{k}}=\|S_{\mathbf{0}}\|^{1-1/2^{k}}C_{0}^{1/2^{k}}C^{|{\boldsymbol{\alpha}}|}.$
-for any $k\geq 1$. We obtain $\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{\mathbf{0}}\|C^{|{\boldsymbol{\alpha}}|}$ by letting $k\to\infty$. ∎
-Lemma 6.4.
-Suppose that Assumption 3.2 holds. For $C>0$, the set
-$\mathscr{M}\coloneqq\{\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}\mid\mbox{tr}\left(S_{\mathbf{0}}\right)=1,\ M(\mathbf{S})\succeq 0,\ M^{\nu}(G\mathbf{S})\succeq 0,\ \|S_{{\boldsymbol{\alpha}}}\|\leq C^{|{\boldsymbol{\alpha}}|},\ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}\}$
-is a convex set whose extreme points are ${\boldsymbol{v}}{\boldsymbol{v}}^{\intercal}\zeta_{{\boldsymbol{u}}}$ where ${\boldsymbol{v}}\in{\mathbb{R}}^{p}$ with $\|{\boldsymbol{v}}\|=1$ and $\zeta_{{\boldsymbol{u}}}=({\boldsymbol{u}}^{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ is the Zeta vector at some ${\boldsymbol{u}}\in\mathcal{X}\cap[-C,C]^{n}$.
-Proof.
-It is clear that $\mathscr{M}$ is a convex set.
-Let $\mathbf{S}$ be an extreme point of $\mathscr{M}$.
-We next show that there exists a sequence $(\xi_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{R}}^{\mathbb{N}^{n}}$ such that $S_{{\boldsymbol{\alpha}}}=\xi_{{\boldsymbol{\alpha}}}S_{\mathbf{0}}$ and $\xi_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}\xi_{{\boldsymbol{\beta}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$.
-Now we fix an arbitrary ${\boldsymbol{\alpha}}_{0}\in\mathbb{N}^{n}$ and prove that $S_{{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{\mathbf{0}}$ for some $\xi_{{\boldsymbol{\alpha}}_{0}}\in{\mathbb{R}}$.
-For $\varepsilon\in\{\pm 1\}$, define two sequences $\mathbf{S}^{(\varepsilon)}\coloneqq(S^{(\varepsilon)}_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{S}}^{p}$ by letting $S^{(\varepsilon)}_{{\boldsymbol{\alpha}}}\coloneqq C^{|{\boldsymbol{\alpha}}_{0}|}S_{{\boldsymbol{\alpha}}}+\varepsilon S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-We claim that $M(\mathbf{S}^{(\varepsilon)})\succeq 0$ and $M^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$, that is, $M_{d}(\mathbf{S}^{(\varepsilon)})\succeq 0$ and $M_{d,k}^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$ for each $d,k\in\mathbb{N}$.
-We only prove that $M^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$ and that $M(\mathbf{S}^{(\varepsilon)})\succeq 0$ could be proven in a similar way.
-Fix an arbitrary vector ${\boldsymbol{w}}=(w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}}})_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}}\in{\mathbb{R}}^{pq|\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}|}$ with each $w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\in{\mathbb{R}}^{pq}$, we need to prove that
-${\boldsymbol{w}}^{\intercal}M^{\nu}_{d,k}(G\mathbf{S}^{(\varepsilon)}){\boldsymbol{w}}$
-$=$ $\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S^{(\varepsilon)}_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
-$=$ $C^{|{\boldsymbol{\alpha}}_{0}|}\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
-$+\varepsilon\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\alpha}}_{0}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}\geq 0.$
-Define a new sequence ${\boldsymbol{z}}\coloneqq(z_{{\boldsymbol{\kappa}}})_{\kappa\in\mathbb{N}^{n}}\in{\mathbb{R}}^{\mathbb{N}^{n}}$ by letting
-$z_{{\boldsymbol{\kappa}}}\coloneqq\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\kappa}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
-for each ${\boldsymbol{\kappa}}\in\mathbb{N}^{n}$.
-Then to prove ${\boldsymbol{w}}^{\intercal}M^{\nu}_{d,k}(G\mathbf{S}^{(\varepsilon)}){\boldsymbol{w}}\geq 0$, it suffices to show that $-z_{0}C^{|{\boldsymbol{\kappa}}|}\leq z_{{\boldsymbol{\kappa}}}\leq z_{0}C^{|{\boldsymbol{\kappa}}|}$ for all ${\boldsymbol{\kappa}}\in\mathbb{N}^{n}$ and then let ${\boldsymbol{\kappa}}={\boldsymbol{\alpha}}_{0}$.
-To this end, we next show that $M({\boldsymbol{z}})\succeq 0$ and ${\boldsymbol{z}}$ is exponentially bounded, and then apply Lemma 6.3.
-For $\ell\in\mathbb{N}$ and ${\boldsymbol{v}}\coloneqq(v_{{\boldsymbol{\kappa}}})_{{\boldsymbol{\kappa}}\in\mathbb{N}^{n}_{\ell}}\in{\mathbb{R}}^{|\mathbb{N}^{n}_{\ell}|}$, we have
-${\boldsymbol{v}}^{\intercal}M_{k}({\boldsymbol{z}}){\boldsymbol{v}}$
-$=$ $\sum_{{\boldsymbol{\kappa}},{\boldsymbol{\delta}}\in\mathbb{N}^{n}_{\ell}}\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})v_{{\boldsymbol{\kappa}}}w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}^{\intercal}\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\kappa}}+{\boldsymbol{\delta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)v_{{\boldsymbol{\delta}}}w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}$
-$=$ $\,({\boldsymbol{v}}\otimes{\boldsymbol{w}})^{\intercal}M_{d+\ell,k}^{\nu}(G\mathbf{S})({\boldsymbol{v}}\otimes{\boldsymbol{w}})\geq 0.$
-Moreover,
-$|z_{{\boldsymbol{\kappa}}}|$ $\leq\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\left|\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\right|\|w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\|\|w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}\|\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}\|S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}+{\boldsymbol{\kappa}}}\|\|G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\|$
-$\leq\left(\sum_{\begin{subarray}{c}({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\\ ({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}\end{subarray}}\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\left|\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\right|\|w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\|\|w_{({\boldsymbol{\beta}},{\boldsymbol{\xi}})}\|\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}\|G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\|C^{|{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}|}\right)C^{|{\boldsymbol{\kappa}}|}.$
-Then, applying Lemma 6.3 to ${\boldsymbol{z}}$ yielding $-z_{0}C^{|{\boldsymbol{\kappa}}|}\leq z_{{\boldsymbol{\kappa}}}\leq z_{0}C^{|{\boldsymbol{\kappa}}|}$.
-Letting ${\boldsymbol{\gamma}}={\boldsymbol{\alpha}}_{0}$, we get ${\boldsymbol{w}}^{\intercal}M_{d,k}^{\nu}(G\mathbf{S}^{(\varepsilon)}){\boldsymbol{w}}\geq 0$, i.e., $M^{\nu}(G\mathbf{S}^{(\varepsilon)})\succeq 0$.
-Similarly, we can prove that $M(\mathbf{S}^{(\varepsilon)})\succeq 0$.
-As $\|S^{(\varepsilon)}_{{\boldsymbol{\alpha}}}\|\leq(2C^{|{\boldsymbol{\alpha}}_{0}|})C^{|{\boldsymbol{\alpha}}|}$, applying Lemma 6.3 again yields that
-$\|S^{(\varepsilon)}_{{\boldsymbol{\alpha}}}\|\leq\mbox{tr}\left(S^{(\varepsilon)}_{0}\right)C^{|{\boldsymbol{\alpha}}|}=(C^{|{\boldsymbol{\alpha}}_{0}|}+\varepsilon\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right))C^{|{\boldsymbol{\alpha}}|}$
-for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-Additionally, as $M(\mathbf{S}^{(\varepsilon)})\succeq 0$, we have $\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)\geq 0$, i.e., $|\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right)|\leq C^{|{\boldsymbol{\alpha}}_{0}|}$.
-Case 1: $\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)=0$ for some $\varepsilon\in\{\pm 1\}$.
-Without loss of generality, we may assume that $\mbox{tr}\left(S^{(-1)}_{\mathbf{0}}\right)=0$ which implies $S^{(-1)}_{{\boldsymbol{\alpha}}}=0$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$ since $M(\mathbf{S}^{(-1)})\succeq 0$.
-Letting ${\boldsymbol{\alpha}}=\mathbf{0}$, we get $S_{{\boldsymbol{\alpha}}_{0}}=C^{|{\boldsymbol{\alpha}}_{0}|}S_{\mathbf{0}}$.
-So, we can set $\xi_{{\boldsymbol{\alpha}}_{0}}=C^{|{\boldsymbol{\alpha}}_{0}|}$ in this case and it holds that $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{{\boldsymbol{\alpha}}}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-Case 2: Now we assume $\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)\neq 0$ for both $\varepsilon\in\{\pm 1\}$ which implies $\mathbf{S}^{(\varepsilon)}/\mbox{tr}\left(S^{(\varepsilon)}_{\mathbf{0}}\right)\in\mathscr{M}$.
-Then, we have a convex combination of two points in $\mathscr{M}$:
-[…]
-which implies $\mathbf{S}=\mathbf{S}^{(1)}/\mbox{tr}\left(S^{(1)}_{\mathbf{0}}\right)$ or $\mathbf{S}^{(-1)}/\mbox{tr}\left(S^{(-1)}_{\mathbf{0}}\right)$.
-Without loss of generality, we may assume that $\mathbf{S}=\mathbf{S}^{(1)}/\mbox{tr}\left(S^{(1)}_{\mathbf{0}}\right)$.
-Then, for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$,
-$C^{|{\boldsymbol{\alpha}}_{0}|}S_{{\boldsymbol{\alpha}}}+S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}=S^{(1)}_{{\boldsymbol{\alpha}}}=\mbox{tr}\left(S^{(1)}_{\mathbf{0}}\right)S_{{\boldsymbol{\alpha}}}=C^{|{\boldsymbol{\alpha}}_{0}|}S_{{\boldsymbol{\alpha}}}+\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right)S_{{\boldsymbol{\alpha}}}.$
-Letting ${\boldsymbol{\alpha}}={\mathbf{0}}$ and $\xi_{{\boldsymbol{\alpha}}_{0}}=\mbox{tr}\left(S_{{\boldsymbol{\alpha}}_{0}}\right)$, we get $S_{{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{\mathbf{0}}$.
-Also, in this case, it holds that $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\alpha}}_{0}}=\xi_{{\boldsymbol{\alpha}}_{0}}S_{{\boldsymbol{\alpha}}}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-Now we have shown that there exists a sequence $(\xi_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subset{\mathbb{R}}^{\mathbb{N}^{n}}$ such that $S_{{\boldsymbol{\alpha}}}=\xi_{{\boldsymbol{\alpha}}}S_{\mathbf{0}}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-Moreover, it is clear from the above arguments that $\xi_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}S_{\mathbf{0}}=S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}S_{{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}\xi_{{\boldsymbol{\beta}}}S_{\mathbf{0}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$ which implies $\xi_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}=\xi_{{\boldsymbol{\alpha}}}\xi_{{\boldsymbol{\beta}}}$.
-Since $S_{\mathbf{0}}\succeq 0$, decompose $S_{\mathbf{0}}$ as $S_{\mathbf{0}}=\sum_{i=1}^{t}\lambda_{i}{\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}$ for some $\lambda_{i}>0$ and ${\boldsymbol{v}}^{(i)}\in{\mathbb{R}}^{p}$, $i\in[t]$, with $\sum_{i=1}^{t}\lambda_{i}=1$ and $\|{\boldsymbol{v}}^{(i)}\|=1$.
-Letting ${\boldsymbol{u}}=(u_{1},\ldots,u_{n})$ with $u_{i}=\xi_{{\boldsymbol{e}}_{i}}$, we get $\mathbf{S}=\sum_{i=1}^{t}\lambda_{i}{\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}\zeta_{{\boldsymbol{u}}}$.
-We have ${\boldsymbol{u}}\in[-C,C]^{n}$ because $|\xi_{{\boldsymbol{e}}_{i}}|=|\mbox{tr}\left(S_{{\boldsymbol{e}}_{i}}\right)|\leq C^{|{\boldsymbol{e}}_{i}|}=C$.
-It remains to show that ${\boldsymbol{u}}\in\mathcal{X}$ and $t=1$.
-First, suppose on the contrary that ${\boldsymbol{u}}\not\in\mathcal{X}$, i.e., there exists ${\boldsymbol{w}}\in\mathcal{Y}$ such that $G({\boldsymbol{u}},{\boldsymbol{w}})\not\succeq 0$.
-Then, there is a ball $\mathcal{O}_{2}\subset{\mathbb{R}}^{m}$ with a radius $\rho>0$ around ${\boldsymbol{w}}$ such that $G({\boldsymbol{u}},{\boldsymbol{y}})\not\succeq 0$ for all ${\boldsymbol{y}}\in\mathcal{O}_{2}$.
-Without loss of genericity, we may assume that there exists ${\boldsymbol{\xi}}\in{\mathbb{R}}^{q}$ and $\delta>0$ such that ${\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\leq-\delta$ for all ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$.
-Define a continuous function $h^{(2)}({\boldsymbol{y}})$ on ${\mathbb{R}}^{m}$ by $h^{(2)}({\boldsymbol{y}})=2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|$ for ${\boldsymbol{y}}\in 2\mathcal{O}_{2}$ and $h^{(2)}({\boldsymbol{y}})=0$ otherwise.
-By Proposition 3.1, there exists a sequence of SOS polynomilas $\{\sigma_{k}\}_{k}$ in ${\mathbb{R}}[{\boldsymbol{y}}]$ that converges to $h^{(2)}$ in the $L^{2}$-norm.
-Hence,
-$\lim_{k\to\infty}\int_{\mathcal{Y}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=\int_{\mathcal{Y}\cap 2\mathcal{O}_{2}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}}){\boldsymbol{\xi}}(2\rho-\|{\boldsymbol{y}}-{\boldsymbol{w}}\|)\mathrm{d}\nu({\boldsymbol{y}})$ (19)
-$\leq$ $\int_{\mathcal{Y}\cap\mathcal{O}_{2}}-\delta\rho\mathrm{d}\nu({\boldsymbol{y}})=-\delta\rho\nu(\mathcal{Y}\cap\mathcal{O}_{2})<0.$
-The last inequality is due to the fact that $\mathrm{supp}(\nu)=\mathcal{Y}$. Then for $k\in\mathbb{N}$ large enough, we have the following contradiction,
-$0>\sum_{i=1}^{t}\lambda_{i}\|{\boldsymbol{v}}^{(i)}\|^{2}\int_{\mathcal{Y}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})=$ $\sum_{i=1}^{t}\int_{\mathcal{Y}}{\boldsymbol{\xi}}^{\intercal}G({\boldsymbol{u}},{\boldsymbol{y}}){\boldsymbol{\xi}}\sigma_{k}({\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\langle I_{p},\lambda_{i}{\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}\rangle$ (20)
-$=$ $\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\left\langle\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{\xi}}{\boldsymbol{\xi}}^{\intercal},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
-where the last inequality is due to the fact that $\sigma_{k}({\boldsymbol{y}})I_{p}\otimes{\boldsymbol{\xi}}{\boldsymbol{\xi}}^{\intercal}$ is an SOS matrix and Proposition 4.1.
-As ${\boldsymbol{u}}\in\mathcal{X}\cap[-1,1]^{n}$, it is easy to see that each ${\boldsymbol{v}}^{(i)}({\boldsymbol{v}}^{(i)})^{\intercal}\zeta_{{\boldsymbol{u}}}\in\mathscr{M}$.
-Then since $\mathbf{S}$ is an extreme point of $\mathscr{M}$, we must have $t=1$.
-∎
-Proof of Theorem 4.3.
-(i) $\Longrightarrow$ (ii).
-Assume that $\mathbf{S}$ has a $p\times p$ matrix-valued representing measure $\Phi=[\phi_{ij}]_{i,j}$ supported on $\mathcal{X}\cap[-C,C]^{n}$.
-Then each measure $\phi_{ij}$ is supported on $\mathcal{X}\cap[-C,C]^{n}$.
-Fix $d\in\mathbb{N}$ and an arbitrary vector ${\boldsymbol{w}}\in{\mathbb{R}}^{p|\mathbb{N}^{n}_{d}|}$.
-Let $\Sigma_{0}({\boldsymbol{x}})\coloneqq([{\boldsymbol{x}}]_{d}\otimes I_{p})^{\intercal}{\boldsymbol{w}}{\boldsymbol{w}}^{\intercal}([{\boldsymbol{x}}]_{d}\otimes I_{p})$.
-We have
-${\boldsymbol{w}}^{\intercal}M_{d}(\mathbf{S}){\boldsymbol{w}}=\mathscr{L}_{\mathbf{S}}(\Sigma_{0})=\int_{\mathcal{X}}\mbox{tr}\left(\Sigma_{0}({\boldsymbol{x}})\mathrm{d}\Phi({\boldsymbol{x}})\right)\geq 0,$
-which implies $M_{d}(\mathbf{S})\succeq 0$. Fix $d,k\in\mathbb{N}$ and an arbitrary PSD matrix $Z\in\mathbb{S}_{+}^{pq|\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}|}$. Let
-$\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k})\otimes I_{pq})^{\intercal}Z(([{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k})\otimes I_{pq}).$
-We have $\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\succeq 0$ for any ${\boldsymbol{x}}\in\mathcal{X}\cap[-C,C]^{n}$. Hence,
-$\mathscr{L}_{\mathbf{S}}\left(\int_{\mathcal{Y}}\langle\Sigma,G\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\geq 0,$
-which implies that $M_{d,k}^{\nu}(G\mathbf{S})\succeq 0$ by Proposition 4.1. Moreover, for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$,
-$\|S_{{\boldsymbol{\alpha}}}\|\leq\|S_{{\boldsymbol{\alpha}}}\|_{F}=\left\|\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})\right\|_{F}=\left\|\left[\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi_{ij}\right]_{i,j}\right\|_{F}\leq\sqrt{\sum_{i,j=1}^{p}(\phi_{ij}(\mathcal{X}\cap[-C,C]^{n}))^{2}}C^{|{\boldsymbol{\alpha}}|}.$
-(ii) $\Longrightarrow$ (i).
-Assume that $\|S_{{\boldsymbol{\alpha}}}\|\leq C_{0}C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-By Lemma 6.3, $\|S_{{\boldsymbol{\alpha}}}\|\leq\mbox{tr}\left(S_{\mathbf{0}}\right)C^{|{\boldsymbol{\alpha}}|}$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}$.
-If $\mbox{tr}\left(S_{\mathbf{0}}\right)=0$, then $S_{\mathbf{0}}=0$ which implies that all $\mathbf{S}=0$ and we are done.
-Now rescale $\mathbf{S}$ if necessary and assume that $\mbox{tr}\left(S_{\mathbf{0}}\right)=1$.
-Then, by Lemma 6.3, $\mathbf{S}$ belongs the set $\mathscr{M}$ in Lemma 6.4.
-As the set $\mathscr{M}$ is compact by Tychonoff’s theorem, applying the Krein-Milman theorem [24] to the convex compact set $\mathscr{M}$ yields that $\mathbf{S}$ belongs to the closure of the convex hull of the extreme set of $\mathscr{M}$.
-Therefore, by Lemma 6.4, we have a coordinate-wise convergence $\mathbf{S}=\lim_{i\to\infty}\mathbf{S}^{(i)}$ where each $\mathbf{S}^{(i)}$ admits a finitely atomic matrix-valued measure supported by $\mathcal{X}\cap[-C,C]^{n}$.
-Then, by the coordinate-wise convergence and Haviland’s theorem (Theorem 2.2), we conclude that $\mathbf{S}$ has a matrix-valued representing measure supported by $\mathcal{X}\cap[-C,C]^{n}$.
-∎
-6.6. Proof of Proposition 4.2
-We first show that (7) is solvable, and strong duality holds between (6) and (7) if Assumption 4.1 holds.
-Lemma 6.5.
-[…]
-Proof.
-Define a linear form $L:{\mathbb{R}}[{\boldsymbol{x}}]_{2d}\to{\mathbb{R}}$ by
-$L(f)=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(f)}\mbox{tr}\left(S_{{\boldsymbol{\alpha}}}\right)f_{{\boldsymbol{\alpha}}},\ \ \forall f=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(f)}f_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{R}}[{\boldsymbol{x}}]_{2d}.$
-Then it is clear that $L(f^{2})=\mathscr{L}_{\mathbf{S}}(f^{2}I_{p})$ for all $f\in{\mathbb{R}}[{\boldsymbol{x}}]_{d}$.
-As $M_{d}(\mathbf{S})\succeq 0$, we have $L(f^{2})\geq 0$ for all $f\in{\mathbb{R}}[{\boldsymbol{x}}]_{d}$.
-Moreover, since $\mathscr{L}_{\mathbf{S}}(\Theta_{d}I_{p})\leq 1$, it holds that $L(1)\leq 1$ and $L(x_{i}^{2d})\leq 1$ for all $i\in[n]$.
-Then, by [30, Lemmas 4.1 and 4.3], $|L({\boldsymbol{x}}^{{\boldsymbol{\alpha}}})|\leq 1$ and hence $|\mbox{tr}\left(S_{{\boldsymbol{\alpha}}}\right)|\leq 1$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}$.
-In particular, for any ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{d}$, as $S_{2{\boldsymbol{\alpha}}}\succeq 0$, we have $\|S_{2{\boldsymbol{\alpha}}}\|\leq\mbox{tr}\left(S_{2{\boldsymbol{\alpha}}}\right)\leq 1$.
-Fix an ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}$.
-We may write ${\boldsymbol{\alpha}}={\boldsymbol{\beta}}+{\boldsymbol{\gamma}}$ for some ${\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in\mathbb{N}^{n}_{d}$.
-As $M_{d}(\mathbf{S})\succeq 0$, by [17, Theorem 7.7.11], there is a contraction $N_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ such that $S_{{\boldsymbol{\alpha}}}=S_{2{\boldsymbol{\beta}}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\gamma}}}^{1/2}$.
-Therefore,
-$\|S_{{\boldsymbol{\alpha}}}\|=\|S_{2{\boldsymbol{\beta}}}^{1/2}N_{{\boldsymbol{\alpha}}}S_{2{\boldsymbol{\gamma}}}^{1/2}\|\leq\|S_{2{\boldsymbol{\beta}}}^{1/2}\|\|N_{{\boldsymbol{\alpha}}}\|\|S_{2{\boldsymbol{\gamma}}}^{1/2}\|\leq\|S_{2{\boldsymbol{\beta}}}^{1/2}\|\|S_{2{\boldsymbol{\gamma}}}^{1/2}\|\leq\mbox{tr}\left(S_{2{\boldsymbol{\beta}}}\right)\mbox{tr}\left(S_{2{\boldsymbol{\gamma}}}\right)\leq 1.$
-holds true. ∎
-Proposition 6.1.
-The problem (7) is solvable. If Assumption 4.1 holds, then there is no dual gap between (6) and (7).
-Proof.
-As the sequence of zero matrices is feasible to (7), by Lemma 6.5, the feasible set of (7) is compact and hence (7) is solvable.
-Now let $\lambda$, $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ be as in Assumption 4.1.
-To prove strong duality, it suffices to show that (7) is strictly feasible.
-Let $\Phi\in\mathfrak{M}^{p}_{+}(\mathcal{X})$ be such that $\Phi=\mbox{diag}(\phi,\ldots,\phi)$ where $\phi$ is the probability measure with uniform distribution on $\mathcal{O}_{1}$, and ${\mathbf{S}}^{\circ}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}}$ where each
-$S_{{\boldsymbol{\alpha}}}=\frac{1}{1+p\int_{\mathcal{X}}\Theta_{d}\mathrm{d}\phi({\boldsymbol{x}})}\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\Phi({\boldsymbol{x}})=\left(\frac{1}{1+p\int_{\mathcal{X}}\Theta_{d}\mathrm{d}\phi({\boldsymbol{x}})}\int_{\mathcal{X}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\mathrm{d}\phi({\boldsymbol{x}})\right)I_{p},\ \ \forall{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d}.$
-Clearly, $\mathscr{L}_{\mathbf{S}^{\circ}}(\Theta_{d}I_{p})<1$.
-We next prove that $M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ})\succ 0$.
-Since $\mathbf{S}^{\circ}$ has a PSD matrix-valued representing measure supported on $\mathcal{X}$, by Proposition 4.1, we obtain $M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ})\succeq 0$.
-Suppose on the contrary that $M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ})\not\succ 0$.
-Now fix a nonzero vector ${\boldsymbol{w}}=(w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}}})_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})\in\mathbb{N}^{n}_{d-d(G)}\times\mathbb{N}^{m}_{k}}\in{\mathbb{R}}^{pq|\mathbb{N}^{n}_{d-d(G)}\times\mathbb{N}^{m}_{k}|}$ with each $w_{({\boldsymbol{\alpha}},{\boldsymbol{\eta}})}\in{\mathbb{R}}^{pq}$ such that ${{\boldsymbol{w}}^{\intercal}}M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ}){\boldsymbol{w}}=0$.
-Let
-$\Sigma({\boldsymbol{y}})=([{\boldsymbol{x}}]_{d-d(G)}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq})^{\intercal}{\boldsymbol{w}}{\boldsymbol{w}}^{\intercal}([{\boldsymbol{x}}]_{d-d(G)}\otimes[{\boldsymbol{y}}]_{k}\otimes I_{pq}).$
-Then it holds
-$\mathscr{L}_{{\mathbf{S}}^{\circ}}\left(\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)=\langle M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ}),{\boldsymbol{w}}{\boldsymbol{w}}^{\intercal}\rangle={{\boldsymbol{w}}^{\intercal}}M_{d-d(G),k}^{\nu}(G{\mathbf{S}}^{\circ}){\boldsymbol{w}}=0.$
-Let
-$T({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{d-d(G)}}\sum_{{\boldsymbol{\eta}}\in\mathbb{N}^{m}_{k}}w_{{\boldsymbol{\alpha}},{\boldsymbol{\eta}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}{\boldsymbol{y}}^{{\boldsymbol{\eta}}}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}\quad\text{and}\quad\Sigma({\boldsymbol{x}},{\boldsymbol{y}})=T({\boldsymbol{x}},{\boldsymbol{y}})T({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}.$
-For each $j\in[p]$, let
-$H_{j}({\boldsymbol{x}},{\boldsymbol{y}})=[T_{(j-1)q+1}({\boldsymbol{x}},{\boldsymbol{y}}),\ldots,T_{jq}({\boldsymbol{x}},{\boldsymbol{y}})]\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}.$
-Then,
-$\left\langle\Sigma,G\right\rangle_{p}=[H_{i}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})]_{i,j\in[p]},$
-and
-$0=\mathscr{L}_{{\mathbf{S}}^{\circ}}\left(\int_{\mathcal{Y}}\left\langle\Sigma,G\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)=\frac{1}{1+p\int_{\mathcal{X}}\Theta_{d}\mathrm{d}\phi({\boldsymbol{x}})}\int_{\mathcal{X}}\sum_{j=1}^{p}\left(\int_{\mathcal{Y}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\right)\mathrm{d}\phi({\boldsymbol{x}}).$
-As $H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\geq 0$ for all ${\boldsymbol{x}}\in\mathcal{X}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, we have
-$\int_{\mathcal{X}}\int_{\mathcal{Y}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})=0,\quad\forall j\in[p].$
-As $G({\boldsymbol{x}},{\boldsymbol{y}})\succeq\lambda I_{p}$ on $\mathcal{O}_{1}\times\mathcal{O}_{2}$, for each $j\in[p]$, we have
-00 0 $=\int_{\mathcal{X}}\int_{\mathcal{Y}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})\geq\int_{\mathcal{O}_{1}}\int_{\mathcal{O}_{2}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}G({\boldsymbol{x}},{\boldsymbol{y}})H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})$
-$\geq\lambda\int_{\mathcal{O}_{1}}\int_{\mathcal{O}_{2}}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})^{\intercal}H_{j}({\boldsymbol{x}},{\boldsymbol{y}})\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}})=\lambda\int_{\mathcal{O}_{1}}\int_{\mathcal{O}_{2}}\sum_{i=1}^{q}T_{(j-1)q+i}({\boldsymbol{x}},{\boldsymbol{y}})^{2}\mathrm{d}\nu({\boldsymbol{y}})\mathrm{d}\phi({\boldsymbol{x}}).$
-Since $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are open, we have $T_{i}({\boldsymbol{x}},{\boldsymbol{y}})\equiv 0$ for each $i=1,\ldots,pq$.
-We then conclude ${\boldsymbol{w}}=0$, a contradiction.
-Similar arguments can be used to prove $M_{d}({\mathbf{S}}^{\circ})\succ 0$.
-Thus, $\mathbf{S}^{\circ}$ is strictly feasible to (7).
-∎
-Proof of Proposition 4.2.
-Suppose that the conclusion is false.
-Then, we can find $d_{i},k_{i}\in\mathbb{N}$, $i\in\mathbb{N}$, such that $d_{i},k_{i}\to\infty$ as $i\to\infty$, and $r_{d_{i},k_{i}}^{\star}>\varepsilon$ for all $i\in\mathbb{N}$.
-By Proposition 6.1, let $\mathbf{S}^{(i)}=(S^{(i)}_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d_{i}}}$ be an optimal solution of (7) with $d=d_{i}$ and $k=k_{i}$.
-Hence, by Proposition 6.1, $\mathscr{L}_{\mathbf{S}^{(i)}}(F)=-r_{d_{i},k_{i}}^{\star}$.
-By Lemma 6.5, $\|S^{(i)}_{{\boldsymbol{\alpha}}}\|\leq 1$ for all ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{2d_{i}}$.
-Complete the sequence $\mathbf{S}^{(i)}$ with zero matrices to obtain $\widetilde{\mathbf{S}}^{(i)}=(S^{(i)}_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$.
-By Tychonoff’s Theorem, we could find a subsequence of $\{\mathbf{S}^{(i)}\}_{i\in\mathbb{N}}$ which pointwisely converges to some $\mathbf{S}^{\star}=(S^{\star}_{\boldsymbol{\alpha}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}$ in the product topology with each $\|S^{\star}_{\boldsymbol{\alpha}}\|\leq 1$.
-For the sake of simplicity and without loss of generality, we may assume that the whole sequence $\{\mathbf{S}^{(i)}\}_{i\in\mathbb{N}}$ converges to $\mathbf{S}^{\star}$.
-It holds that $M(\mathbf{S}^{\star})\succeq 0$ and $M^{\nu}(G\mathbf{S}^{\star})\succeq 0$ from the pointwise convergence.
-Then, by Theorem 4.3, $\mathbf{S}^{\star}$ admits a matrix-valued representing measure $\Phi^{\star}$ supported on $\mathcal{X}\cap[-1,1]^{n}$.
-Therefore, as $F({\boldsymbol{x}})$ is PSD on $\mathcal{X}\cap[-1,1]^{n}$, we have
-$-\varepsilon>-r_{d_{i},k_{i}}^{\star}=\mathscr{L}_{\mathbf{S}^{(i)}}(F)\to\mathscr{L}_{\mathbf{S}^{\star}}(F)=\int_{\mathcal{X}\cap[-1,1]^{n}}F\mathrm{d}\Phi^{\star}\geq 0,$
-a contradiction. ∎
 7. Conclusions
 We extend various classical Positivstellensätze to provide SOS-structured characterizations for polynomial matrices that are positive (semi)definite over a semialgebraic set defined by a PMI with UQs.
-Under the Archimedean condition, we first present a Scherer-Hol-type Positivstellensatz, along with a sparse version for scalar polynomials being positive over the semialgebraic set in the presence of correlative sparsity.
-We also derive Putinar–Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze without assuming the Archimedean condition.
-These results significantly extend existing work on Positivstellensätze, and some of them remain novel and intriguing even in the absence of UQs.
-Our results are valuable for ensuring robustness of PMIs over a prescribed set with uncertainty, allowing potential applications in areas such as optimal control, systems theory, and certifying stability.
+Under the Archimedean condition, we first present a matrix-valued Positivstellensatz incorporating universal quantifiers, along with a sparse version for scalar-valued polynomial objectives, leveraging the correlative sparsity patterns.
+We also derive two generalized Positivstellensätze without assuming the Archimedean condition.
+These results significantly extend the existing work on Positivstellensätze, and some of them remain novel and intriguing even in the absence of UQs.
+Our results are valuable in ensuring the robustness of PMIs over a prescribed set with uncertainty, allowing potential applications in areas such as optimal control, systems theory, and certifying stability.
 Acknowledgements
+The authors appreciate Professor Konrad Schmüdgen for providing the result in [44, Theorem 10.25], which significantly simplified the proof of Theorem 3.1.
 Feng Guo was supported by the Chinese National Natural Science Foundation under grant 12471478.